### v1

# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

Abstract.
We concentrate our attention on the qualitative study of the phase portraits of the three-parameter FitzHugh-Nagumo family and its compactification. We divide the study into three scenarios based on the parameters. One of the scenarios is characterised by the existence of a Double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry (singularity of codimension two). In this case, we explicitly exhibit the Pitchfork, Hopf, Belyakov, Double Homoclinic bifurcation/transition curves unfolding the singularity of codimension 2 and plot the bifurcation diagrams. We bridge this analysis with the theory on the associated slow-fast family and the existence of canards. We complete our study with the global compactification of the phase portraits for the family under consideration. This study complements the work summarised in Georgescu, Rocşoreanu and Giurgiţeanu, Global Bifurcations in FitzHugh-Nagumo Model, Trends in Mathematics: Bifurcations, Symmetry and Patterns (2003).
Key words and phrases:
Keywords: FitzHugh-Nagumo equation, Bifurcations, Global phase portrait, Poincaré compactification, Singular system.
[…]
1 Lisbon School of Economics and Management, Lisbon University, Centro de Matemática Aplicada à Previsão e Decisão Económica,
Rua do Quelhas 6, 1200-781 Lisboa, Portugal,
2 School of Mathematics, Statistics and Computer Science, College of Science, University of Tehran, Tehran, Iran and School of Mathematics, Institute for Research in Fundamental Sciences (IPM), P. O. Box 19395-5746, Tehran, Iran
2020 Mathematics Subject Classification:
Primary: 34C60. Secondary: 34E15, 34E17, 37G15, 37G35
1. Introduction
Information in nerve fibers is encoded through action potentials (electrical membrane changes). Alan Hodgkin and Andrew Huxley, in 1939, managed to make the first intracellular recording of an action potential by inserting microelectrodes into the giant axons of squid. In their experiments, they demonstrated that the action potentials are the result of two effects:
  • a rapid inward current carried by sodium (Na+) ions and
  • a slow activating outward current carried by potassium (K+) ions.
Hodgkin and Huxley discovered that the permeability of the membrane for $\text{Na}^{+}$ and $\text{K}^{+}$ was regulated independently, with the conductance depending on both time and the membrane potential. They provided a computational model for the action potential in a single cell, known as the Hodgkin-Huxley (HH) model [11]. Details and an overview may be found in the review [3].
1.1. From the Hodgkin-Huxley model to the FitzHugh-Nagumo equations
Hodgkin and Huxley modelled the measured changes in current by introducing precise probabilistic terms to capture ion channels, which may be either open or closed. The HH model elucidates the electrical behaviour of ion channels within the cell membrane, by addressing the passage of $\text{Na}^{+}$ and $\text{K}^{+}$ ions. The HH model consists of four ordinary differential equations, each corresponding to one of the state variables [11, Equation 1]. This system is nonlinear, making it analytically difficult to solve explicitly. Nonetheless, numerical simulations enable the exploration of certain properties and general behaviours, such as the existence of oscillations and excitability.
The fact that the HH model correctly captures these excitable dynamics was a major realisation as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (cf. [11, Equation 1]). While the HH model effectively replicates many neuronal physiological phenomena, it s very complex. In [7], Richard FitzHugh introduced a simplified model to capture the dynamics of neuronal excitability, which was refined by Jinichi Nagumo two years later [19]. This model is known as the FitzHugh-Nagumo (FHN) model. FitzHugh focused on preserving dynamical features of the HH model, namely the presence of excitability and oscillations.
1.2. The FitzHugh-Nagumo equations
FitzHugh started from the oscillator equation (introduced in 1920) by Balthasar Van der Pol [25], that admits oscillations which are relaxation-like. This means there are periods of “low” and “high” states characterised by rapid transitions between them. Van der Pol’s equation is built from the simple differential equation for the damped harmonic oscillator. He replaced the damping constant by a damping function that depends quadratically on $x,$ thus introducing a nonlinearity ($\dot{x}$ represents the first derivative of the physical position $x$ with respect to $t$ and $c\in\mathbb{R}^{+}_{0}$):
$\ddot{x}+c(x^{2}-1)\dot{x}+x=0.$
There is only “effective” damping for $|x|<1$ while for $x^{2}>1$ the nonlinear term describes amplification. To interpret the dynamics of the van der Pol equation, one can use the Liénard map [16]
$y=\frac{\dot{x}}{c}+\left(\frac{x^{3}}{3}-x\right),$
giving rise to a system of two differential equations (on the plane):
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}x,\end{array}\right.$ (1.4)
from which one may see the separation in time scales of both equations. For $c\gg 1$, while the first one evolves fast (order of $\mathcal{O}(c)$), the second one is much slower of the order of $\mathcal{O}(1/c),$ where $\mathcal{O}$ stands for the usual Landau notation Given a real map $f:\mathbb{R}\to\mathbb{R}$ and $c\in\mathbb{R}$, we say that $f$ is of order $\mathcal{O}(c)$ if there exist $M\in\mathbb{R}^{+}$ and $x_{0}\in\mathbb{R}$ such that $|f(x)|<Mc$, for $x>x_{0}$.. Building on the van der Pol oscillator, FitzHugh considered the equation:
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)+z\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x-a+by)\end{array}\right.$ (1.5)
where (see [3] and Remark 1.1):
  1. (1)
    the parameter $z$ mimics the membrane current density;
  2. (2)
    the variable $x$ is related to the membrane voltage and the $\text{Na}^{+}$ activation and
  3. (3)
    the variable $y$ is related to the $\text{Na}^{+}$ inactivation and the $\text{K}^{+}$ activation.
Equations (1.5) work as an activator-inhibitor model. Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with an electrical circuit [19]. Although the FHN model is primarily used to describe neuronal and cardiac systems, it can also be applied in a range of other biological contexts. The FHN equations provide a simplified framework for representing interconnected positive and negative feedback loops, enabling the generation of a wide range of responses such as switches, pulses, and stable oscillations [24]. Among the numerous models used to investigate cardiac cells, the FHN model is notable for being one of the simplest and most extensively researched, effectively representing the general dynamic characteristics of cardiac cells [1, 4].
Remark 1.1.
System (1.5) mimics an excitable system and there are limitations regarding the interpretation of the original variables of the HH model. Therefore, in the rest of the paper, we will “play safe” and stick to a mathematical motivation in terms of bifurcations and dynamical systems.
1.3. Novelty
Results in [21, pp.180] synthesised the global bifurcation diagram for the FHN model. It has been obtained by putting together as in a “huge puzzle”, all local bifurcation diagrams obtained in Chapters 2–4. See also Section 2.4 of [20]
The main novelty of this paper is the dissection of the global bifurcations diagrams of [21, pp. 180] and provide a more complete bifurcation analysis of model (1.5), namely the location of Hopf, Pitchfork and Double-zero bifurcations in some cases that will be specified in Subsection 1.4.
Besides the case-study $c=\mathcal{O}(1)$, we also study the asymptotic case $c\to\pm\infty$ ($\Leftrightarrow\varepsilon=1/c^{2}\to 0$), where canards are observed [12, 14], and we finish the analysis with the compactification of the phase portrait of (1.5) on the Poincaré disc. To the best of our knowledge, the application of this procedure to equation (1.5) is new, providing additional information about the trajectories which tend to or come from infinity. For some cases, we give the complete description of the phase portraits for (1.5) in the Poincaré disc (i.e. in the compactification of $\mathbb{R}^{2}$ adding the circle $\mathbb{S}^{1}$ of the infinity) modulo topological equivalence.
1.4. Structure
This paper is structured as follows: Section 2 introduces the definitions and preliminary concepts of blow-up and Poincaré compactification. In Section 3, we discuss the finite equilibria of (3.1) for the following three cases:
Case A:
$a=0$
Case B:
$b=0$
Case C:
[…]
and investigate possible bifurcations, including Pitchfork, Double Homoclinic and Hopf bifurcations. The general bifurcation analysis of (1.5) is difficult to tackle; however we have been able to make some progress in Cases A, B and C. In Section 4, we analyse the asymptotic dynamics of (1.5) when $c\to\pm\infty$ and we relate the existence of canards with the periodic solution emerging from the Hopf bifurcation.
Section 5 is dedicated to the study of the global phase portrait of (1.5) in the Poincaré disc. The phase portraits for each connected component in Cases A, B and C are completely illustrated. This completes the work started in [8]. Section 6 finishes the paper.
Throughout this paper, we have endeavoured to make a self-contained exposition bringing together all topics related to the proofs. We have drawn illustrative figures to make the paper easily readable and all results are illustrated with numerical simulations using the software Matlab R2015b.
2. Preparatory section
In this section, we introduce some terminology for polynomial vector fields on $\mathbb{R}^{2},$ that will be used in the remaining sections. Let $f_{(a,b,c)}$ be a smooth vector field on $\mathbb{R}^{2}$ with flow given by the unique solution $x(t)=\varphi(t,x)\in\mathbb{R}^{2}$ of the three-parameter family
$\dot{x}=f_{(a,b,c)}(x),\qquad x(0)=x_{0}\in\mathbb{R}^{2},$ (2.1)
where $\0(a,b,c)\in\mathbb{R}\times\mathbb{R}\times\mathbb{R}\backslash\{0\}$.
2.1. Useful terminology
The center manifold of a non-hyperbolic equilibrium $p\in\mathbb{R}^{2}$ of (2.1) is the set of solutions whose behaviour around $p$ is controlled neither by the exponential attraction of the stable manifold nor by the exponential repulsion of the unstable manifold.
If $Df_{(a,b,c)}$ evaluated at a given equilibrium $p$, $Df_{(a,b,c)}(p)$, has an eigenvalue with zero real part, the center manifold plays an important role and this is the set where bifurcations might occur.
For a fixed triple $\0(a,b,c)\in\mathbb{R}\times\mathbb{R}\times\mathbb{R}\backslash\{0\}$, an equilibrium point $p$ of (2.1) is called nilpotent singularity when all eigenvalues of Jacobian matrix $Df_{(a,b,c)}(p)$ are zero but $Df_{(a,b,c)}(p)\neq 0$, and is called linearly zero when $Df_{(a,b,c)}(p)\equiv 0.$ Further, the singular point $p$ is called semi-hyperbolic if exactly one eigenvalue of $Df_{(a,b,c)}(p)$ is equal to $0$
Throughout this paper, we study the Double-zero (DZ) singularity with symmetry $\mathbb{Z}_{2}$ for a family of differential equations corresponding to the case described in [27, pp. 400]. The unfolding of this singularity of codimension two involves lines of Belyakov transitions, Pitchfork, Hopf, Double Homoclinic and saddle-node of limit cycles. For the notion of super/subcritical Hopf bifurcation, we use the terminology of Golubitsky and Schaeffer [9, Chapter IV, Section 2]. We suggest the reading of [15, 18, 27] for a complete understanding of these bifurcations, as well as the sufficient conditions that prompt their existence.
Remark 2.1.
In this paper, we refer to an equilibrium point as undergoing a Belyakov transition if at least one pair of eigenvalues of the Jacobian matrix (evaluated at the equilibrium) changes from real to complex conjugate or vice versa, while the sign of their real part remains unchanged. Although such a transition is typically considered in higher-dimensional systems, we adopt this terminology here for the sake of brevity and clarity.
2.2. Blow-up Method
The blow-up method is one of the important methods to study the topological behaviour of solutions of a dynamical system in the neighbourhood of nilpotent and linearly zero singularities. In this method, an equilibrium point is “expanded” into a line or a circle, and new equilibria on the line or circle are examined. There are various methods for doing a blow-up. In this paper, we concentrate our attention on blow-ups in a given direction, a method which will be used in the Section 5 of the present article. In what follows, following [2], we review some concepts related to the topic.
Consider a planar polynomial differential equation associated with the vector field $\mathbf{X}=(P,Q),$ in $\mathbb{R}^{2}$ of the form:
$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{2})=P_{m}(x_{1},x_{2})+\cdots\\ \\ \dot{x}_{2}=Q(x_{1},x_{2})=Q_{m}(x_{1},x_{2})+\cdots\end{array}\right.$ (2.2)
where
  • $P,Q$ are coprime polynomials;
  • $P_{m},Q_{m}$ are homogeneous polynomials of degree $m\in\mathbb{N}$ and
  • the dots $\cdots$ stand for high-order terms in the variables $x_{1},x_{2}.$
Since $m>0,$ the origin is a singularity. By using the change of coordinates
$(x_{1},x_{2})\mapsto(r\cos(\theta),r\sin(\theta)),$
system (2.2) may be written in polar coordinates (if $r\neq 0$) as:
$\left\{\begin{array}[]{l}\dot{r}=\mathcal{R}(\theta)r+\cdots\\ \\ \dot{\theta}=\mathcal{T}(\theta)+\cdots\end{array}\right.$ (2.3)
where:
  • $\mathcal{R}$ and $\mathcal{T}$ are polynomials in $\cos{(\theta)}$ and $\sin{(\theta)}$ and
  • the dots $\cdots$ stand for high order terms in $r.$
If $\mathcal{T}\neq 0,$ all the solution curves tending (as $t\to\pm\infty$) to the origin are tangent to the solutions $\theta^{\star}\in[0,2\pi)$ of the equation $\mathcal{T}(\theta)=0.$ The map $\mathcal{T}$ is often called the characteristic polynomial and $\theta^{\star}$ is called characteristic direction associated with (2.2). The polynomial $\mathcal{T}$ in the cartesian coordinates $(x_{1},x_{2})$ may be written as:
$\mathcal{T}(x_{1},x_{2})=x_{1}Q_{m}(x_{1},x_{2})-x_{2}P_{m}(x_{1},x_{2}).$
Definition 2.2.
Blow-up in $x_{1}$ direction is a “non-bijective” change of coordinates defined as
$(x_{1},z)\mapsto(x_{1},zx_{1})=(x_{1},x_{2}),$
where $z\geq 0$ is a new variable.
This change of coordinates projects the origin of (2.2) on the line $x_{1}=0.$ The expression of system (2.2) after the blow-up in the $x_{1}$ direction may be written as:
$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{1}z),\\ \\ \dot{z}=\dfrac{Q(x_{1},x_{1}z)-zP(x_{1},x_{1}z)}{x_{1}}.\end{array}\right.$ (2.4)
After the blow-up, we may cancel the common factor $x_{1}^{m-1}.$ The $x_{1}$ directional blow-up is equivalent on $x_{1}\neq 0$ and $\theta\neq\pi/2,3\pi/2,$ since there exists an analytic change of coordinates bringing $(r,\theta)$ and $(x_{1},x_{2}).$ To investigate the behavior of the solutions of system (2.2) around equilibria, it is necessary to examine the behaviour of the equilibrium points of (2.4) on the line $x_{1}=0.$ If some of these points are singular, then we may continue the blow-up process.
2.3. Poincaré Compactification
In this section, we present a concise overview of the Poincaré compactification following Chapter 5 of [5], a technique used in Section 5 of the present paper. Consider the polynomial vector field (2.2) in $\mathbb{R}^{2}$ where $d_{1}$ and $d_{2}$ are the algebraic degrees of $P$ and $Q$ respectively and $d={\max}\{d_{1},d_{2}\}.$
2.3.1. Preliminaries for the construction
First identify $\mathbb{R}^{2}$ with the plane $\Pi$ in $\mathbb{R}^{3}$ defined by $(y_{1},y_{2},y_{3})=(x_{1},x_{2},1).$ The sphere $\mathbb{S}^{2}=\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}:y_{1}^{2}+y_{2}^{2}+y_{3}^{2}=1\}$ is called the Poincaré sphere and is tangent to the plane $\Pi$ at $(0,0,1).$ The sphere may be written as
$H^{+}\cup H^{-}\cup\mathbb{S}^{1}$
where
$H^{+}=\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{3}>0\},\quad H^{-}=\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{3}<0\}$
and the equator $\mathbb{S}^{1}$ given by:
$\mathbb{S}^{1}=\{(y_{1},y_{2},y_{3})\in\mathbb{S}^{2}:y_{3}=0\}.$
If $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ define $\Delta(x)$ as $\sqrt{x_{1}^{2}+x_{2}^{2}+1}\neq 0.$
Definition 2.3.
The projection of $\mathbf{X}$ from $\mathbb{R}^{2}$ to $\mathbb{S}^{2}$ is given by the central projections
$f^{+}:\mathbb{R}^{2}\to\mathbb{S}^{2}\quad\text{and}\quad f^{-}:\mathbb{R}^{2}\to\mathbb{S}^{2}$
where $f^{+}(x_{1},x_{2})$ is the intersection point of the line passing through the origin and $(x_{1},x_{2})\in\mathbb{R}^{2}$, with $H^{+}$. More precisely:
$f^{+}(x_{1},x_{2})=\left(\frac{x_{1}}{\Delta(x)},\frac{x_{2}}{\Delta(x)},\frac{1}{\Delta(x)}\right).$
Analogously, we define $f^{-}$ (substituting $H^{+}$ by $H^{-}$) as:
$f^{-}(x_{1},x_{2})=\left(-\frac{x_{1}}{\Delta(x)},-\frac{x_{2}}{\Delta(x)},-\frac{1}{\Delta(x)}\right).$
We obtain vector fields in each hemisphere (analytically conjugate to the initial vector field $\mathbf{X}$. For $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the induced vector field on $H^{+}$ and $H^{-}$ is defined, respectively, by:
$\overline{\mathbf{X}}(y)$ $=Df^{+}(x)\mathbf{X}(x),\quad\text{where}\quad y=f^{+}(x),$
$\overline{\mathbf{X}}(y)$ $=Df^{-}(x)\mathbf{X}(x),\quad\text{where}\quad y=f^{-}(x).$
Remark 2.4.
The vector field $\overline{\mathbf{X}}$ on $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ is everywhere tangent to $\mathbb{S}^{2}.$
The points at infinity of $\mathbb{R}^{2}$ (each direction is associated with two points) are in bijective correspondence with the points of the equator of $\mathbb{S}^{2}.$ In the next subsection, we are going to extend the induced vector field $\overline{\mathbf{X}}$ from $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ to $\mathbb{S}^{2}$.
Definition 2.5.
The extended vector field on $\mathbb{S}^{2}$ is called the Poincaré compactification of the vector field $\mathbf{X}$ on $\mathbb{R}^{2},$ and it is denoted by $p(\mathbf{X}).$
2.3.2. About the construction
We use smooth charts to make calculations. For $y=(y_{1},y_{2},y_{3})\in\mathbb{S}^{2},$ we use the six local charts given by
$U_{k}=\{y\in\mathbb{S}^{2}:y_{k}>0\},\quad V_{k}=\{y\in\mathbb{S}^{2}:y_{k}<0\},\quad k=1,2,3.$
The corresponding local maps $\Phi_{k}:U_{k}\to\mathbb{R}^{2}$ and $\Psi_{k}:V_{k}\to\mathbb{R}^{2}$ are defined as
$\Phi_{k}(y)=-\Psi_{k}(y)=\left(\frac{y_{m}}{y_{k}},\frac{y_{n}}{y_{k}}\right),\quad\text{for}\quad m<n\quad\text{and}\quad m,n\neq k.$
We denote by $z=(u,v)$ the value of $\Phi_{k}(y)$ or $\Psi_{k}(y)$ for any $k\in\{1,2,3\}$, such that $(u,v)$ will play different roles depending on the local chart we are considering. Observe that points lying in $\mathbb{S}^{1}$, in any chart, have $v=0$.
We perform a detailed calculation of the expression of $p(\mathbf{X})$ in the local chart $U_{1}$. From (2.2), we have $\mathbf{X}(x)=(P(x_{1},x_{2}),Q(x_{1},x_{2}))$. Then $\overline{\mathbf{X}}(y)=Df^{+}(x)\mathbf{X}(x)$ with $y=f^{+}(x)$ and, using the Chain rule, one gets:
$D\Phi_{1}(y)\overline{\mathbf{X}}(y)=D\Phi_{1}(y)\circ Df^{+}(x)X(x)=D(\Phi_{1}\circ f^{+})(x)\mathbf{X}(x).$
Let $\overline{\mathbf{X}}|_{U_{1}}$ denote the system defined as $D\Phi_{1}(y)\overline{\mathbf{X}}(y)$. Then since
$(\Phi_{1}\circ f^{+})(x)=\left(\frac{x_{2}}{x_{1}},\frac{1}{x_{1}}\right)=(u,v),$
we have
$\overline{\mathbf{X}}|_{U_{1}}=v^{2}\left(-\frac{u}{v}P\left(\frac{1}{v},\frac{u}{v}\right)+\frac{1}{v}Q\left(\frac{1}{v},\frac{u}{v}\right),-P\left(\frac{1}{v},\frac{u}{v}\right)\right).$
On the other hand, one knows that $\rho(y)=y_{3}^{d-1}=v^{d-1}m(z)$ where
$m(z)=(1+u^{2}+v^{2})^{\frac{1-d}{2}}.$
Then we get:
$\rho\left(\overline{\mathbf{X}}|_{U_{1}}\right)(z)=v^{d+1}m(z)\left(-\frac{u}{v}P\left(\frac{1}{v},\frac{u}{v}\right)+\frac{1}{v}Q\left(\frac{1}{v},\frac{u}{v}\right),-P\left(\frac{1}{v},\frac{u}{v}\right)\right).$
Remark 2.6.
In order to prove that the extension of $\rho\overline{\mathbf{X}}$ to $p(\mathbf{X})$ is defined on the whole of $\mathbb{S}^{2}$ we notice that while $\overline{\mathbf{X}}|_{U_{1}}$ is not well defined when $v=0$, and $p(\mathbf{X})|_{U_{1}}=\rho\overline{\mathbf{X}}|_{U_{1}}$ is well defined along $v=0$, since the multiplying factor $v^{d+1}$ cancels any factor of $v$ which may appear in the denominator. The same line of arguments may be applied to the other local charts.
To simplify the extended vector field we also make a change in the time variable and remove the factor $m(z)$. We still keep a vector field on $\mathbb{S}^{2}$ which is $C^{\omega}$–equivalent to $\mathbf{X}$ on any of the hemispheres $H^{+}$ and $H^{-}$.
2.3.3. Explicit expressions
The expression for $p(\mathbf{X})$ in local chart $(U_{1},\Phi_{1})$ is given by:
$\dot{u}=v^{d}\left[-uP\left(\frac{1}{v},\frac{u}{v}\right)+Q\left(\frac{1}{v},\frac{u}{v}\right)\right],\hskip 14.22636pt\dot{v}=-v^{d+1}P\left(\frac{1}{v},\frac{u}{v}\right),$ (2.5)
the expression for $(U_{2},\Phi_{2})$ is:
$\dot{u}=v^{d}\left[P\left(\frac{u}{v},\frac{1}{v}\right)-uQ\left(\frac{u}{v},\frac{1}{v}\right)\right],\hskip 14.22636pt\dot{v}=-v^{d+1}Q\left(\frac{u}{v},\frac{1}{v}\right),$ (2.6)
and the expression for $(U_{3},\Phi_{3})$ is:
$\dot{u}=P(u,v),\hskip 14.22636pt\dot{v}=Q(u,v).$ (2.7)
Remark 2.7.
For $i=1,2,3$, the expressions of the vector field $p(\mathbf{X})$ in the local chart $(V_{i},\Psi_{i})$ are identical to those in the local chart $(U_{i},\Phi_{i}),$ up to a multiplication by $({-1})^{d-1}$.
To study $\mathbf{X}$ in the complete plane $\mathbb{R}^{2}$, including its behaviour near infinity, it suffices to work on $H^{+}\cup\mathbb{S}^{1}$, which we call the Poincaré disc. All calculations can be performed in the three charts $(U_{1},\Phi_{1})$, $(U_{2},\Phi_{2})$, and $(U_{3},\Phi_{3})$ in which case the expressions are given by the formulas (2.5), (2.6) and (2.7). To obtain (2.5) we start with (2.2) and introduce coordinates $(u,v)$ through the equality:
$(x_{1},x_{2})=\left(\frac{1}{v},\frac{u}{v}\right).$
In each local chart, the local representative of $p(\mathbf{X})$ is a polynomial vector field.
Definition 2.8.
We call finite (resp. infinite) singular points of $\mathbf{X}$ the singular points of $p(\mathbf{X})$ which lie in $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ (resp. $\mathbb{S}^{1}$).
2.3.4. Useful consequences
Following Chapter 5 of [5], we list some useful consequences of the theory described above.
  1. (1)
    If $y\in\mathbb{S}^{1}$ is an infinite singular point, then $-y$ is also a singular point.
  2. (2)
    Since the local behavior near $-y$ is the local behavior near $y$ multiplied by $(-1)^{d-1}$, then:
    1. (a)
      the orientation of the orbits changes when the degree $d$ is even.
    2. (b)
      if $d$ is even and $y\in\mathbb{S}^{1}$ is a stable node of $p(\mathbf{X})$, then $-y$ is an unstable node.
  3. (3)
    Infinite singular points appear in pairs of diametrically opposite points so that it is enough to study half of them. Using the degree of the vector field one can determine the other half.
  4. (4)
    The integral curves of $\mathbb{S}^{2}$ are symmetric with respect to the origin, such that it is sufficient to represent the flow of $p(\mathbf{X})$ only in the closed northern hemisphere, the so called Poincaré disc.
2.3.5. Infinite singular points
The theory described here will be useful in Section 5 of the present paper. We aim to study the local phase portrait at infinite singular points. For this we choose an infinite singular point $(u,0)$ and start by looking at the expression of the linear part of the vector field $p(\mathbf{X})$. Denote by $P_{i}$ and $Q_{i}$ the homogeneous polynomials of degree $i\in\mathbb{N}_{0}$ for $i=0,1,...,d$ such that
$\left\{\begin{array}[]{l}P=P_{0}+P_{1}+...+P_{d},\\ \\ Q=Q_{0}+Q_{1}+...+Q_{d}.\end{array}\right.$
Then $(u,0)\in\mathbb{S}^{1}\cap(U_{1}\cup V_{1})$ is an infinite singular point of $p(\mathbf{X})$ if and only if
$F(u)\equiv Q_{d}(1,u)-uP_{d}(1,u)=0.$
Similarly $(u,0)\in\mathbb{S}^{1}\cap(U_{2}\cup V_{2})$ is an infinite singular point of $p(\mathbf{X})$ if and only if
$G(u)\equiv P_{d}(u,1)-uQ_{d}(u,1)=0.$
The Jacobian of the vector field $p(\mathbf{X})$ at the point $(u,0)$ is
$\left(\begin{array}[]{cc}F^{\prime}(u)&Q_{d-1}(1,u)-uP_{d-1}(1,u)\\ \\ 0&-P_{d}(1,u)\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}G^{\prime}(u)&P_{d-1}(u,1)-uQ_{d-1}(u,1)\\ \\ 0&-Q_{d}(u,1)\end{array}\right)$
if $(u,0)$ belongs to $U_{1}\cup V_{1}$ or $U_{2}\cup V_{2}$, respectively.
Our discussion is concentrated on isolated singularities in the equator.
Following [5], among the hyperbolic singular points at infinity only nodes and saddles may appear. All the semi-hyperbolic singular points can appear at infinity. If one of these hyperbolic or semi-hyperbolic singularities at infinity is a (topological) saddle, then the straight line defined by $v=0$, representing the equator of $\mathbb{S}^{2}$, is a stable or unstable manifold, or a center manifold. The same property also holds for semi-hyperbolic singularities of saddle-node type. They can have their hyperbolic sectors split in two different ways depending on the Jacobian matrix of the system in the charts $U_{1}$ or $U_{2}$. These matrices may be either
$\left(\begin{array}[]{cc}a&\star\\ \\ 0&0\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}0&\star\\ \\ 0&a\end{array}\right)$
with $a\neq 0$ and $\star\in\mathbb{R}$. The sense of the orbits can also be the opposite. The nilpotent points, as well as the singularities with zero linear part, have a behavior at infinity that is quite a bit more complicated than the hyperbolic and elementary singular points. Blow-up is needed to study them.
3. Finite equilibria and bifurcation analysis
Our object of study is the analysis of the family of differential equations (1.5), which can be written as:
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x-a+by)\end{array}\right.$ (3.1)
where $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}.$ We can absorb the parameter $z$ of (1.5) into the variable through the change of variable $y\to y+z$ and redefinition of $a\to a-bz.$ The model (3.1) is invariant under the transformation $(t,c)\to(-t,-c)$ so we can restrict to the case $c>0$, which means that the stability of the equilibria for $c>0$ is the reverse of that of $c<0$.
Let us denote by $f_{(a,b,c)}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ the vector field associated with (3.1). For $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}$, the Jacobian matrix of $f_{(a,b,c)}$ at a general point $(x,y)\in\mathbb{R}^{2}$ is given by:
$Df_{(a,b,c)}(x,y)=\left(\begin{array}[]{cc}c-cx^{2}&c\\ \\ -1/c&-b/c\end{array}\right).$ (3.5)
It is immediate to deduce:
Lemma 3.1.
The divergence of (3.1) is given by $c-cx^{2}-b/c,$ which is strictly negative for $b>c^{2}.$
By the Bendixson criterion [5, Theorem 7.10], the negativity of the previous result ensures that the system (3.1) cannot have periodic orbits and, consequently, no limit cycles in the open set defined by the inequality $b>c^{2}.$ From now on, we divide the analysis into Cases A, B and C (see Subsection 1.4), depending on the parameters $a,b$ and $c$ of (3.1).
3.1. Case A ($a=0$)
Consider the invertible linear map in $\mathbb{R}^{2}$ defined by $\kappa(x,y)=(-x,-y)$, whose action on $\mathbb{R}^{2}$ is isomorphic to that of $\mathbb{Z}_{2}$ (rotation of $\pi$ around the origin). Since
$f_{(0,b,c)}\circ\kappa=\kappa\circ f_{(0,b,c)},$
we may say that:
Lemma 3.2.
The vector field $f_{(0,b,c)}$ is $\mathbb{Z}_{2}(\kappa)$–equivariant.
For all $a=0$, $b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\},$ we know that $E_{1}=(0,0)$ is an equilibrium of (3.1). If $b\in(-\infty,0)\cup[1,+\infty),$ then system (3.1) has two extra equilibria given explicitly by
$E_{2}=\kappa(E_{3})=\left(-\sqrt{\dfrac{3(b-1)}{b}},\dfrac{1}{b}\sqrt{\dfrac{3(b-1)}{b}}\right)$ (3.6)
and $E_{3}=\kappa(E_{2})=\left(\sqrt{\dfrac{3(b-1)}{b}},-\dfrac{1}{b}\sqrt{\dfrac{3(b-1)}{b}}\right).$
From (3.5), we deduce that the trace and determinant operators of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,3,$ are given and denoted, respectively, by:
$\mbox{Tr}(E_{1})$ $=c-\dfrac{b}{c},\qquad\qquad\qquad\qquad\mbox{Det}(E_{1})=1-b,$
$\mbox{Tr}(E_{2,3})$ $=-\dfrac{2c^{2}b-3c^{2}+b^{2}}{cb},\qquad\mbox{Det}(E_{2,3})=2(b-1).$
It is easy to see that $E_{1}$ is a hyperbolic saddle for $b>1$ (since $\mbox{Det}(E_{1})<0$). Using (3.5), the jacobian matrix of the vector field $f_{(0,1,\pm 1)}$ at $E_{1}\equiv E_{2}\equiv E_{3}$ is given by:
(±$\left(\begin{array}[]{cc}\pm 1&\pm 1\\ \\ \mp 1&\mp 1\end{array}\right).$ (3.10)
It is easy to verify that the matrix (3.10) is non-hyperbolic and has a double zero eigenvalue. Our main result relies on the existence of a Double-zero bifurcation (DZ) with $\mathbb{Z}_{2}(\kappa)$–symmetry of $f_{(0,1,\pm 1)}$ at the singularity $E_{1}$.
Remark 3.3.
Consider the following linear change of coordinates, $\Psi,$ and its inverse:
$\Psi(x,y)=(x+y,-x)\quad\text{and}\quad\Psi^{-1}(x,y)=(-y,x+y).$
One may easily check that:
$\Psi^{-1}\circ f_{(0,1,\pm 1)}\circ\Psi(x,y)$ $=$ $\Psi^{-1}\circ f_{(0,1,\pm 1)}(x+y,-x)$
$=$ $\Psi^{-1}\left(-\frac{(x\pm y)^{3}}{3}+y,-y\right)$
$=$ $\left(y,-\frac{1}{3}(x\pm y)^{3}\right),$
so that the equations in Jordan form are given by
$\left\{\begin{array}[]{l}\dot{x}=y\\ \\ \dot{y}=-\frac{1}{3}(x\pm y)^{3},\end{array}\right.$ (3.11)
which contains part of the truncated form of degree 3 of the versal deformation of a Double-zero (DZ) bifurcation with $\mathbb{Z}_{2}$–symmetry [27, pp. 400, Case $c=-1$]. The next result deals with the description of all bifurcation curves passing through the bifurcation point $(b,c)=(1,\pm 1)$ and that characterises a DZ bifurcation of $f_{(0,b,c)}$ with $\mathbb{Z}_{2}(\kappa)$–symmetry. We focus the proof on the analysis around the bifurcation parameter $(b,c)=\left(1,1\right)$; the analysis close to $(b,c)=\left(1,-1\right)$ has a similar treatment (reversing time).
Theorem 3.4.
With respect to the vector field $f_{(0,b,c)},$ the equilibrium $E_{1}\equiv E_{2}\equiv E_{3}$ undergoes a DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1).$ The global representations of the transition/bifurcation curves in the space of parameters $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$ are as follows (schematic lines for $a=0$ have been plotted in Figures 1 and 3):
  1. (1)
    The equilibrium $E_{1}$ undergoes a supercritical (with respect to $b$) Pitchfork bifurcation along the line:
    $T_{P}=\{(b,c):b=1\}$
    giving rise to the equilibria $E_{2}$ and $E_{3}$ defined in (3.6) for $b>1$.
  2. (2)
    The equilibrium $E_{1}$ undergoes a Belyakov transition along the union of parabola:
    $T^{1}_{F}=\{(b,c):b=-c^{2}\pm 2c\}.$
  3. (3)
    The equilibrium $E_{1}$ undergoes a subcritical (with respect to $b$) Hopf bifurcation along part of the parabola:
    $T_{H}^{1}=\left\{(b,c):b=c^{2},\,|c|<1\right\}.$
  4. (4)
    The equilibria $E_{2}$ and $E_{3}$ undergo a Belyakov transition along the algebraic curves:
    $T^{2,3}_{F}=\left\{(b,c):\left(b^{2}+2cb-2c^{2}b+3c^{2}\right)\left(b^{2}-2cb-2c^{2}b+3c^{2}\right)=0\right\}.$
  5. (5)
    The equilibria $E_{2}$ and $E_{3}$ undergo a supercritical (with respect to $b$) Hopf bifurcation along the algebraic curves:
    $T^{2,3}_{H}=\left\{(b,c):b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>1\right\}.$
  6. (6)
    There is a double homoclinic cycle to $E_{1}$ along the approximated curve:
    $DH:\quad 0=\frac{7b^{2}+10bc^{2}-17c^{2}}{15c^{3}}\sqrt{\frac{-7b^{2}+5bc^{2}+2c^{2}}{5b}},\quad b\in(0,c).$
Remark 3.5.
According to [21, Theorem 3.5.1] one knows that there is a Bautin bifurcation along the curve parametrized by $c$ defined by
$\left\{(a,b,c)\in\mathbb{R}^{2}\times\mathbb{R}^{+}:\left(\pm\frac{4}{3}\left(c\sqrt{c^{2}-1}-c^{2}+1\right)\sqrt[4]{1-\frac{1}{c^{2}}},c^{2}\pm c\sqrt{c^{2}-1},c\right),c\geq 1\right\}$ (3.12)
In particular, in the three-dimensional bifurcation space $(a,b,c)$, this gives rise to a saddle-node bifurcation surface associated with two non-hyperbolic cycles, generating the line SNL defined by $b=c^{2}\pm c\sqrt{c^{2}-1}$ and $a=0$ of Figure 6. The formula of [21] is explicitly written for $c\geq 1$, however due to the invariance under the reversible transformation $(t,c)\to(-t,-c)$, it also holds for $c\leq-1$.
The curve SNL of Figure 6 does not make part of a generic unfolding of a DZ bifurcation with symmetry; this is why it is represented as a dashed line.
When $a=0$, the points $\textbf{Q}_{17}$ and $\textbf{Q}_{18}$ of [21, pp. 180] coincide with the point $(b,c)=(1,1)$. This coincidence yields a codimension-three DZ–bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$, whose analysis is beyond the scope of the present paper. A complete understanding of this bifurcation remains an open problem and is briefly discussed in Section 6.
Figure 1. Illustration of Theorem 3.4. Left: The bifurcation curves associated to (3.1) for Case A. Linear stability of $E_{1}$ for $f_{(0,b,c)}$. Right: Linear stability of $E_{2}$ and $E_{3}$ for $f_{(0,b,c)}$. The notation $T_{P}$, $T_{H}^{1}$, $T_{F}^{1}$, $T_{P}$ $T_{H}^{2,3}$, $T_{F}^{2,3}$ follows from Theorem 3.4 and sd, uf, un, sf, and sn refer to saddle, unstable focus, unstable node, stable focus and stable node. The curve $b=0$ corresponds to a subcritical pitchfork bifurcation from where the equilibria $E_{2}$ and $E_{3}$ (3.6) collapse into $E_{1}$.
Remark 3.6.
Theorem 3.4 generalizes the analysis of [8] and [21] by locating a codimension 2 bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$ for the vector field $f_{(0,b,c)}.$ For the sake of completeness, we plot a computer-assisted global phase portraits on Table 1 of Appendix. Figure 8 sketches the main features of the phase portraits in the different regions – within each region of the figure, the dynamics is $C^{1}$–conjugated to the respective figure.
Proof of Theorem 3.4.
The first statement of the result comes from the normal form (3.11). The eigenvalues of the Jacobian matrix (3.5) at $E_{1},$ $E_{2}$ and $E_{3},$ when the equilibria exist, are:
$E_{1}:\quad$ $\lambda_{\pm}^{1}=\dfrac{c^{2}-b\pm\sqrt{(c^{2}+b)^{2}-4c^{2}}}{2c},$
$E_{2},E_{3}:\quad$ $\lambda_{\pm}^{2,3}=\frac{3{c}^{2}-{b}^{2}-2{c}^{2}b\pm\sqrt{4{c}^{4}{b}^{2}-12{c}^{4}b-4{c}^{2}{b}^{3}+9{c}^{4}+2{c}^{2}{b}^{2}+{b}^{4}}}{2cb}.$
Based on linear analysis, the type of equilibria for (3.1) ($a=0$) and for $b,c\in[-4,4]\times[-2,2]$ in different regions defined by (3.1) are provided in Figures 1 and 3.
  1. (1)
    We prove the existence of a Pitchfork bifurcation for $b=1$ using the normal form truncated at order 3 [27, Section 20.1E]. Let $b=1+\mu$, where $\mu\in\mathbb{R}$ is an additional bifurcation parameter. For $\mu=0$ the Jacobian matrix (3.5) associated with $E_{1}$ admits eigenvalues: 00 and $\frac{c^{2}-1}{c}$, $c\neq 0,1$. Define the diffeomorphism $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ as
    $\mathbf{x}\equiv(x,y)\mapsto\left(\frac{c^{2}y-x}{c^{2}-1},\frac{x-y}{c^{2}-1}\right).$
[…]
    $\bar{f}_{(0,b,c)}(\mathbf{x})\equiv f_{(0,b,c)}(\mathbf{x})-Df_{(0,b,c)}(\mathbf{0})\mathbf{x}.$
    Therefore, we have
    $\dot{\mathbf{x}}=$ $T^{-1}Df_{(0,b,c)}(\mathbf{0})T\mathbf{x}+T^{-1}\bar{f}_{(0,b,c)}T\mathbf{x}$ (3.13)
    $=$ $\begin{pmatrix}1&c^{2}\\ 1&1\end{pmatrix}\begin{pmatrix}c&c\\ -\frac{1}{c}&-\frac{1}{c}\end{pmatrix}\begin{pmatrix}\frac{1}{1-c^{2}}&\frac{c^{2}}{c^{2}-1}\\ \frac{1}{c^{2}-1}&\frac{1}{1-c^{2}}\end{pmatrix}+$
    $\begin{pmatrix}-cy\mu-\frac{1}{3}cx^{3}\\ \frac{c^{2}-1}{c}x+\frac{c^{2}-\mu-1}{c}y-\frac{c}{3}x^{3}\end{pmatrix}\Big{|}_{x\rightarrow\frac{c^{2}y-x}{c^{2}-1},y\rightarrow\frac{x-y}{c^{2}-1}}$
    from where we conclude that (in the new coordinates above):
    $\left\{\begin{array}[]{l}\dot{x}=-\frac{c}{3(c^{2}-1)^{3}}\Big{(}3c^{2}x^{2}y-x^{3}-3c^{4}xy^{2}+c^{6}y^{3}+3(c^{2}-1)^{2}\mu x-3(c^{2}-1)^{2}\mu y\Big{)},\\ \\ \dot{y}=-\frac{1}{c(c^{2}-1)^{3}}\Big{(}3c^{4}x^{2}y-c^{2}x^{3}-3c^{6}xy^{2}+c^{8}y^{3}+3(c^{2}-1)^{2}\mu x-3(c^{2}-1)^{2}\mu y-3(c^{2}-1)^{4}y\Big{)}.\end{array}\right.$
    We aim to find a center manifold in the form of $h(x,\mu):={\alpha_{0}}^{2}\mu+\alpha_{1}x^{2}+\alpha_{2}x\mu.$ Since the center manifold must satisfy the equation ([27, Equation 3.2.7])
    $\dot{x}\frac{\partial h}{\partial x}-\dot{y}\big{|}_{(x,h(x,\mu))}=0,$
    we obtain $\alpha_{0}=\alpha_{1}=0,$ and $\alpha_{2}=\frac{1}{\left(c^{2}-1\right)^{2}}.$ Then, the dynamics of (3.13) restricted to the center manifold $y=\frac{\mu x}{(c^{2}-1)^{2}}$ is given by:
    $\dot{x}=-\frac{c}{c^{2}-1}\mu x+\frac{c}{3\left(c^{2}-1\right)^{3}}x^{3}$ (3.14)
    meaning that there is a Pitchfork bifurcation at $\mu=0$. If $c>1$, then the bifurcation is supercritical with respect to $\mu$ and also with respect to $b$ in the original equation, as depicted in Figure 2. The analysis for the case $c>1$ is different from the case $10<c<$ 1, although the conclusion is the same.
    Figure 2. Illustration of the Pitchfork bifurcation of (3.14) for $c>1$ and close to $\mu=0(\Leftrightarrow b=1)$. (a) $\mu<0$, (b) $\mu=0$ and (c) $\mu>0$.
  2. (2)
    We should prove that the eigenvalues of the Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ change from real to complex (non-real). The proof of this item follows by observing that the eigenvalues associated with the Jacobian matrix (3.5) at $E_{1}$ are complex (non-real) for $b>-c^{2}+2c$ or $b<-c^{2}-2c,$ and real otherwise. The sign of their real part does not change.
  3. (3)
    Let $b=c^{2}$ and $|c|<1$. The Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ has a pair of purely imaginary eigenvalues $i\sqrt{1-c^{2}}$ with $-1<c<1.$ Let $b=c^{2}+\mu$ where $\mu\in\mathbb{R}$ is a bifurcation parameter. The eigenvalues associated with the Jacobian matrix at $E_{1}$ for $\mu\neq 0$
    $\begin{pmatrix}c&c\\ -\frac{1}{c}&-c-\frac{1}{c}\mu\end{pmatrix},$
    are given by
    $\lambda(\mu)=-\frac{1}{2c}\mu\pm\frac{1}{2c}\sqrt{(\mu+2c^{2}+2c)(\mu+2c^{2}-2c)}.$ (3.15)
    For $\mu$ close to zero and $|c|<1$, there exists a pair of complex conjugate eigenvalues. Furthermore, since
    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}=-\frac{1}{2c}\neq 0,$
    this means that there is a non-degenerate Hopf bifurcation at $\mu=0$ ([17, pp. 299–300]). By using the linear transformation
    $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}&\dfrac{1}{2}\dfrac{c}{\sqrt{1-c^{2}}}\\ 0&-\dfrac{1}{2}\dfrac{1}{c\sqrt{1-c^{2}}}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix},$
    the system (3.1) turns into its Jordan form
    $\left\{\begin{array}[]{l}f^{1}:\,\dot{u}=-\frac{1}{12}c{u}^{3}-\frac{1}{4}\frac{{c}^{2}}{\sqrt{1-{c}^{2}}}v{u}^{2}-\frac{1}{4}{\frac{{c}^{3}}{1-{c}^{2}}}{v}^{2}u-\frac{1}{12}\frac{{c}^{4}}{\left(1-{c}^{2}\right)^{3/2}}{v}^{3}-\sqrt{1-{c}^{2}}v,\\ \\ f^{2}:\,\dot{v}=\sqrt{1-{c}^{2}}u.\end{array}\right.$
    By calculating the first Lyapunov coefficient ([26, Equations (2) and (3)]), we have:
    $l:=\dfrac{1}{16\omega}\left(R_{1}+\omega R_{2}\right),$
    where
    $\omega:=$ $\sqrt{1-{c}^{2}}$
    $R_{1}:=$ $f^{1}_{uv}\left(f^{1}_{uu}+f^{1}_{vv}\right)-f^{2}_{uv}\left(f^{2}_{uu}+f^{2}_{vv}\right)-f^{1}_{uu}f^{2}_{uu}+f^{1}_{vv}f^{2}_{vv}\quad\text{and}$
    $R_{2}:=$ $f^{1}_{uuu}+f^{1}_{uvv}+f^{2}_{uuv}+f^{2}_{vvv}.$
    Therefore, using Maple, we get
    $l=\frac{c}{32(c^{2}-1)}<0$
    and we may conclude that the limit cycle is stable for $10<c<$ 1 and the bifurcation is subcritical – [27, Remarks 1 and 2, pp. 383]. In this case $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable for $\mu>0$ with an asymptotically stable periodic solution for $\mu<0$. The same conclusion holds for $b$.
  4. (4)
    The proof is similar as item (2) by taking into account that
    $4{c}^{4}{b}^{2}-12{c}^{4}b-4{c}^{2}{b}^{3}+9{c}^{4}+2{c}^{2}{b}^{2}+{b}^{4}$
    $=\left(b^{2}+2cb-2c^{2}b+3c^{2}\right)\left(b^{2}-2cb-2c^{2}b+3c^{2}\right).$
  5. (5)
    Using linear analysis, one knows that Hopf bifurcation exists only if $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0$ and $\mbox{Det}(E_{2}),\mbox{Det}(E_{3})>0.$ Indeed,
    $1\displaystyle 3{c}^{2}-{b}^{2}-2{c}^{2}b=0\Leftrightarrow b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>$ 1 .
    Figure 3. The bifurcation curves associated to (3.1) for Case A. Within each depicted region, phase portraits are $C^{1}$–equivalent. (a) The global phase portraits on the Poincaré disc corresponding to Regions a–p have been sketched in Figures 8. (b) The local portraits corresponding to Regions 1–28 have been plotted in Table 1 of Appendix.
    The inequality $b>1$ serves to ensure that the equilibria $E_{2}$ and $E_{3}$ exist.
    Let $b=c\left(-c\pm\sqrt{c^{2}+3}\right).$ The Jacobian matrix at $E_{2,3}$ has a pair of purely imaginary eigenvalues ±$\pm i\sqrt{2c\sqrt{c^{2}+3}-2-2c^{2}}$ with $c>1.$ Let $b=-c^{2}\pm c\sqrt{c^{2}+3}+\mu$ where $\mu\in\mathbb{R}$ denotes the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ associated with the Jacobian matrix at $E_{2,3}$
    $\begin{pmatrix}-\frac{c\left(2{c}^{2}+2c\sqrt{{c}^{2}+3}\pm 2\mu\mp 3\right)}{\mp{c}^{2}+c\sqrt{{c}^{2}-3}\pm\mu}&c\\ -\frac{1}{c}&{\frac{{c}^{2}\mp c\sqrt{{c}^{2}+3}-\mu}{c}}\end{pmatrix},$
    are given by (computations performed in Maple):
    $\lambda(\mu)=\dfrac{-2c\sqrt{c^{2}+3}\mu-\mu^{2}\pm\sqrt{\sigma}}{2c(\mu-c^{2}+c\sqrt{c^{2}+3})}$
    where :
    $\sigma=\mu^{4}+4c(\sqrt{c^{2}+3}-2c)\mu^{3}-4c^{2}(6c\sqrt{c^{2}+3}-7c^{2}-5)\mu^{2}$
    $+8c^{3}(2\sqrt{c^{2}+3}(3c^{2}+1)-6c^{3}-11c)\mu-8c^{4}(c\sqrt{c^{2}+3}\left(4c^{2}+5\right)-11c^{2}-4c^{4}-3).$
    There is a pair of complex conjugate eigenvalues when $\sigma<0$ and
    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}=-\frac{\sqrt{c^{2}+3}}{c(\sqrt{c^{2}+3}\mp c)}.$
    Since the first Lyapunov coefficient is
    $l={\frac{c\left(\left(4c^{2}+3\right)\left(\sqrt{c^{2}+3}-c\right)-3c\right)\left(\sqrt{c^{2}+3}c\left(21+44c^{2}+16c^{4}\right)-c^{2}\left(69+68c^{2}+16c^{4}\right)-9\right)}{32\left(\sqrt{c^{2}+3}c-c^{2}-1\right)^{2}\left(\sqrt{c^{2}+3}-c\right)^{7}}},$
    we may deduce, using Maple, that these limit cycles are unstable for $c>0$.
  6. (6)
    This item follows from [8, pp. 201] and [21, formula (3.2.29)].
Figure 4. The classification of the equilibrium point $E_{1}$ and the corresponding bifurcation curves for (3.1) in Case B. In this figure, un, sn, uf and sf refer to unstable (stable) node and unstable (stable) focus, respectively. (a) The local phase portraits associated with Regions 1–8 have been plotted in Figure 2. (b) The global phase portraits on the Poincaré disc associated with Regions I–IV have been plotted in Figure 9.
Remark 3.7.
Numerically, we may observe that the non-degenerate Hopf bifurcation of items (3) and (5) of Theorem 3.4 are, respectively, subcritical and supercritical. The bifurcation $T_{H}^{1}$ yields a stable periodic solution, say $C^{s}$, and $T_{H}^{2,3}$ generates two unstable and $\kappa$–symmetric periodic solutions – confirm on Table 1 of Appendix.
3.2. Case B ($b=0$)
Model (3.1) for $b=0$ has only one equilibrium given explicitly by:
$E_{1}=\,\left(a,-a+\frac{1}{3}a^{3}\right),$
and the linear part of $f_{(a,0,c)}$ at $E_{1}$ has eigenvalues
$\frac{1}{2}c(1-a^{2})\pm\frac{1}{2}\sqrt{(ca^{2}-c+2)(ca^{2}-c-2)}.$ (3.16)
Then we may conclude that:
Theorem 3.8.
With respect to the vector field $f_{(a,0,c)}$ of (3.1), the global representations of the transition/bifurcation curves in the space of parameters $\0(a,c)\in\mathbb{R}\times\mathbb{R}\backslash\{0\}$ are as follows:
  • I.
    The equilibrium $E_{1}$ undergoes a Belyakov transition along the hyperbola:
    $T^{b0}_{F}=\left\{(a,c):c=\pm\frac{2}{a^{2}-1}\right\}.$
  • II.
    The equilibrium $E_{1}$ undergoes a Hopf bifurcation along the lines:
    $T^{b0}_{H}=\left\{(a,c):a=\pm 1\right\}.$
    The Hopf bifurcation is supercritical at $a=-1$ and subcritical at $a=1$.
Schematic bifurcation lines for $b=0$ have been plotted in Figure 4. Illustrative pictures in the phase space have been depicted in Figure 9 in the case where the eigenvalues of (3.5) at $E_{1}$ are complex non-real.
Proof.
We are going to concentrate the proof in the case $b=0$ and $c>0$. Again, the analysis of the case $b=0$ and $c<0$ has a similar treatment (reversing time).
  • I.
    The first claim is a direct consequence of (3.16).
  • II.
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues ±$\pm i.$ Let $a=\pm 1+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix (3.5) at $E_{1}$ are (computations performed with Maple):
    $\lambda(\mu)=\mp c\mu-\frac{1}{2}c{\mu}^{2}\pm\frac{1}{2}\sqrt{4{c}^{2}{\mu}^{2}\pm 4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4}.$
    For $\mu$ close to zero, we have $04{c}^{2}{\mu}^{2}+4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4<$ 0, which means that $\lambda(\mu)$ is complex non-real. Furthermore $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}\big{|}_{\mu=0}=\mp c\neq 0.$ Following [26, Equations (2) and (3)], the first Lyapunov coeficient is $l=-\frac{1}{32}c$ which is negative for $c>0.$ Hence the bifurcation is subcritical at $a=1$ and supercritical for $a=-1$.
    For the case $a=-1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable focus for $\mu>0$ with an asymptotically stable periodic solution emerging for $\mu>0$. For the case $a=1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ and unstable focus for $\mu<0$ with an asymptotically stable periodic solution emerging $\mu>0$. This finishes the proof.
Remark 3.9.
We may observe in Figure 9 that, for $c>0$ (resp. $c<0$), the Hopf bifurcation (with respect to $a$) is supercritical for $a=-1$ giving rise to a stable (resp. unstable) periodic solution which disappears at a subcritical Hopf bifurcation at $a=1,$ see regions 5-8 of Table 2.
3.3. Case C ($a>0,$ $10<b<$ 1)
If $(x^{\star},y^{\star})$ is an equilibrium of (3.1) under the conditions $a>0,$ $0<b<1,$ then:
$x^{\star}-a+by^{\star}=0\Leftrightarrow y^{\star}=\frac{a-x^{\star}}{b}.$
By substituting $y^{\star}=\frac{a-x^{\star}}{b}$ into the first equation of (3.1), then $x^{\star}$ is a root of $H(x)=0,$ where
$H(x)=-\frac{1}{3}cx^{3}+c\left(1-\frac{1}{b}\right)x+\frac{ac}{b}.$
The discriminant associated with the polynomial $H(x)$ is given by
$\Delta=\frac{4}{3}c^{4}\left(1-\frac{1}{b}\right)^{3}-3\frac{c^{4}a^{2}}{b^{2}}.$
For $10<b<$ 1, the discriminant satisfies $\Delta<0$. According to Cardano’s formula, this implies that the polynomial $H(x)$ has exactly one real root, denoted by $x^{\star}.$ Then the equilibrium of (3.1) is explicitly given by
$\left(x^{\star},\frac{a-x^{\star}}{b}\right)=(x^{\star},y^{\star})=:E_{1}.$
The characteristic equation of the Jacobian matrix evaluated at $(x^{*},y^{*})$ is
$\lambda^{2}+\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}\lambda+b{x^{*}}^{2}-b+1=0.$
We may then conclude that:
Proposition 3.10.
The following assertions hold for the vector field $f_{(a,b,c)}$ of (3.1), under the conditions $a>0,$ and $10<b<$ 1:
  1. (1)
    The equilibrium point $E_{1}$ is stable if $\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}>0.$ Conversely, it is unstable if $\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}<0.$
  2. (2)
    A non-degenerate Hopf bifurcation occurs at $c^{2}=\frac{b}{1-{x^{*}}^{2}}.$ The resulting limit cycle is attracting when $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and repelling for $c<\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
Note that $x^{\star}\neq\pm 1$ (if $x^{\star}=\pm 1$, then $b=0$, which is a contradiction). Illustrative pictures in the phase space have been depicted in Figure 10. As before, we use the notation $\mbox{Tr}(E_{1})$ and $\mbox{Det}(E_{1})$ the trace and determinant operators of $Df_{(a,b,c)},$ evaluated at $E_{1}$.
Proof.
Under the condition $0<b<1,$ one knows that $\det(E_{1})>0.$
  1. (1)
    The result follows by noticing that $\mbox{Tr}(E_{1})=-\frac{c^{2}{x^{*}}^{2}-c^{2}+b}{c}.$ If $\mbox{Tr}(E_{1})<0$, then the equilibrium $E_{1}$ is stable. Conversely, it is unstable if $\mbox{Tr}(E_{1})>0$.
  2. (2)
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues ±$\pm i\sqrt{b{x^{*}}^{2}-b+1}.$ Let $b=c^{2}\left(1-{x^{*}}^{2}\right)+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix at $E_{1},$
    $\begin{pmatrix}c(1-x^{*})&c\\ -\frac{1}{c}&c{x^{*}}^{2}-c-\frac{\mu}{c}\end{pmatrix},$
    are
    $\lambda(\mu)=\frac{-\mu\pm\sqrt{\left(2c^{2}{x^{*}}^{2}-2c^{2}-\mu-2c\right)\left(2c^{2}{x^{*}}^{2}-2c^{2}-\mu+2c\right)}}{2c}.$
    If
    $\min_{c\geq 0}{c\left(\pm 1-c+c{x^{*}}^{2}\right)}<\mu<\max_{c\geq 0}{c\left(\pm 1-c+c{x^{*}}^{2}\right)},$
    then the expression under the square root is negative. Hence there exists a pair of complex non real conjugate eigenvalues. The result follows by observing that (computations performed with Maple):
    $\frac{d{\rm Re}\,{\lambda(\mu)}}{d\mu}=-\frac{1}{2c}\neq 0.$
Remark 3.11.
The statement of Proposition 3.10 says nothing about the criticality of the Hopf Bifurcation. Analytically, it is very difficult to compute the first Lyapunov exponent of (3.1) at an equilibrium whose explicit expression we do not know. However, numerics of Figure 10 suggest that the Hopf bifurcation is supercritical for for $c=\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and subcritical for $c=-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
4. Canards for the singular case $c\to+\infty$ in Cases A and B
The goal of this section is the study of the dynamics of (3.1) when $|c|\to+\infty$. Many results concerning the asymptotic behaviour of the solution of the FN model as $|c|\to+\infty$ are available in the literature – see [21, 22, 23]. Our concern is to match these results with the local bifurcation study performed in Section 3. We are particularly interested in Cases A and B whose results have been described in Theorems 3.4 and 3.8. By changing the time scale of (3.1), considering $\tau\to t/c$ and $\varepsilon\to 1/c^{2},$ we obtain an equivalent version of (3.1) as follows (written with respect to the slow time $\tau$):
$\left\{\begin{array}[]{l}\varepsilon\dot{x}=\displaystyle\left[y-\left(\frac{x^{3}}{3}-x\right)\right]=:f(x,y,\varepsilon)\\ \\ \dot{y}=\displaystyle-(x-a+by)=:g(x,y,\varepsilon)\end{array}\right.$ (4.1)
where $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}.$ For $0<\varepsilon\ll 1,$ or equivalently $|c|\gg 1,$ system (4.1) is what we call a fast-slow system (see [10, 14] and references therein).
A fast-slow system is a system of differential equations in which some variables have their derivatives with larger magnitude than others. This leads to a two time-scale system. The general approach to this type of system starts by grouping the variables in two disjoint sets: fast variable $x$ and slow variable $y$. By allowing the dynamics to be separated into two distinct phases, slow and fast, we can study each one independently and understand its underlying dynamics. This separation is introduced in system (4.1) via the parameter $\varepsilon.$
4.1. The critical manifold and bifurcations
We first analyse the singular case $\varepsilon=0.$ We follow the analysis performed in [10, 12, 13]. The critical manifold associated with (4.1) is the set
$\mathcal{C}_{0}=\{(x,y)\in\mathbb{R}^{2}:f(x,y,0)=0\}=\{(x,y)\in\mathbb{R}^{2}:y=x^{3}/3-x\}$ (4.2)
Definition 4.1.
With respect to (4.1) with critical manifold $\mathcal{C}_{0}$:
  1. (1)
    The subset $S\subset\mathcal{C}_{0}$ is said to be normally hyperbolic if for all $p=(x,y)\in{S}$, we have:
[…]
  2. (2)
    A normally hyperbolic set $S$ is said to be attracting (resp. repelling) if, for all $p\in{S}$, all eigenvalues of $\frac{\partial f}{\partial x}\left(p,0\right)$ have negative (resp. positive) real part, respectively.
Definition 4.2.
With respect to (4.1) with critical manifold $\mathcal{C}_{0}$, the point $p\in\mathcal{C}_{0}$ is said to be a fold point if:
$\frac{\partial f}{\partial x}\left(p,0\right)=0\,,\quad\frac{\partial^{2}f}{\partial x^{2}}\left(p,0\right)\neq 0\quad\text{ and }\quad\frac{\partial f}{\partial y}\left(p,0\right)\neq 0.$
If $g\left(p,0\right)\neq 0$, the fold point is called regular.
Coming back to system (4.1), we have $f\left(x,y,0\right)=y-x^{3}/3+x$ and, for $p=(x,y)\in\mathcal{C}_{0}$, one has:
$\frac{\partial f}{\partial x}\left(p,0\right)=x^{2}-1\neq 0\qquad\text{ if }\,\,x\neq\pm 1.$
Therefore, if $b=3/2$, the equilibria $E_{2}=(-1,2/3)$ and $E_{3}=(1,-2/3)$ are fold points (according to the Definition 4.2) since:
$\frac{\partial f}{\partial x}(E_{2,3},0)=0,\quad\frac{\partial^{2}f}{\partial^{2}x}(E_{2,3},0)\neq 0\quad\text{and}\quad\frac{\partial f}{\partial y}(E_{2,3},0)\neq 0$
and $S_{0}=\mathcal{C}_{0}\setminus\left\{E_{2},E_{3}\right\}$ is normally hyperbolic. We may split $S_{0}$ into three disjoint open subsets:
$\mathcal{C}_{0\text{L}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:x<-1\right\}},$
$\mathcal{C}_{0\text{M}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:-1<x<1\right\}}\quad\text{and}$
$\mathcal{C}_{0\text{R}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:x>1\right\}}.$
Therefore, we get
$S_{0}=\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{M}}\cup\mathcal{C}_{0\text{R}},$ (4.3)
where $\mathcal{C}_{0\text{L}}$ and $\mathcal{C}_{0\text{R}}$ are repelling subsets and $\mathcal{C}_{0\text{M}}$ is attracting. This means that in the fast flow, trajectories of (4.1) move backwards either to $\mathcal{C}_{0\text{L}}$ or $\mathcal{C}_{0\text{R}}$ and towards to $\mathcal{C}_{0\text{M}}$ – see Figure 5.
Figure 5. (a) Illustration of the critical manifold associated to system (4.1). (b) Illustration of the dynamics of system (4.1) for the case $a=0,$ when there is a unique equilibrium.
Yet in the singular case ($\varepsilon=0$), depending on the values of $a,b\in\mathbb{R}$, there exists at least one and at most three equilibria, as result of the intersection of the cubic curve, $f(x,y,0)=0,$ with the line defined by $x-a+by=0$ ($\Leftrightarrow\dot{y}=0).$ Moreover, when the system has three equilibria only one of the following scenarios may occur (cf. [10]):
Case 1:
either all three equilibria belong to $\mathcal{C}_{0\text{M}}$ or else
Case 2:
they are each one in a distinct region, $\mathcal{C}_{0\text{L}},$ $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$.
The independence of the time scales when $\varepsilon=0$ makes the system simpler to understand. Fenichel’s Theorem [6] says that, near the normally hyperbolic part of $\mathcal{C}_{0}$, when $0<\varepsilon\ll 1,$ the dynamics of the perturbed system is similar to that of the singular system, with a deviation of $\mathcal{O}\left(\varepsilon\right).$ The smaller the $\varepsilon,$ the more similar the system trajectories are to those described in the singular system ($\varepsilon=0$). Fenichel’s Theorem guarantees that for $0<\varepsilon\ll 1,$ there exists a set $S_{\varepsilon}$, $C^{1}$–close to $S_{0}$ (in the Haussdorf topology), that exhibits the same behaviour as $S_{0}$.
Definition 4.3.
The set $S_{\varepsilon}$ is referred to as the slow manifold of system (4.1).
When the system’s equilibrium is also a fold point, a supercritical (with respect to $b$) Hopf bifurcation occurs, giving rise to a periodic solution. In this section, the canard phenomenon described in [12, 13] bridges the local periodic solution from the Hopf bifurcation to the global limit cycle, as we proceed to explain.
Definition 4.4.
A trajectory segment of a fast-slow system (4.1) is a canard if it stays within $\mathcal{O}\left(\varepsilon\right)$ distance to a repelling branch of a slow manifold for a time that is $\mathcal{O}\left(1\right)$ on the slow time scale $\tau=t\varepsilon$.
Canards are related to equilibria of (4.1) that occur at fold points of the slow manifold. The next definitions deal with the general equation parametrised by $\lambda\in\mathbb{R}$:
$\left\{\begin{array}[]{l}\varepsilon\dot{x}=f(x,y,\lambda,\varepsilon)\\ \\ \,\dot{y}=g(x,y,\lambda,\varepsilon)\end{array}\right.$ (4.4)
We now introduce the definition of singular fold point, which differs from that of a fold point – it is required to coincide with an equilibrium point of (4.1):
Definition 4.5.
With respect to (4.4), let $p=\left(x,y\right)$ be a fold point. The point $p$ is called a singular fold if:
$f\left(p,\lambda,0\right)=0,$ $\frac{\partial f}{\partial x}\left(p,\lambda,0\right)=0,$ (4.5)
$\frac{\partial^{2}f}{\partial x^{2}}\left(p,\lambda,0\right)\neq 0,$ $\frac{\partial f}{\partial y}\left(p,\lambda,0\right)\neq 0\quad\text{and}\quad g\left(p,\lambda,0\right)=0.$
Definition 4.6.
A singular fold point $p$ is said to be regular if:
$\frac{\partial g}{\partial x}\left(p,\lambda,0\right)\neq 0\quad\text{ and }\quad\frac{\partial g}{\partial\lambda}\left(p,\lambda,0\right)\neq 0.$ (4.6)
The behaviour around a regular singular fold point is described in the next theorem. The existence of canards is coupled with the existence of a Hopf bifurcation, which in such systems is referred to as a singular Hopf bifurcation.
Theorem 4.7 ([12, 13], adapted).
Consider a fast-slow system of the form (4.4) where $\left(x,y\right)=\left(0,0\right)$ is a regular singular fold point for $\lambda=0$, and that may be written in the following form:
$\dot{x}=-y\,l_{1}\left(x,y,\lambda,\varepsilon\right)+x^{2}\,l_{2}\left(x,y,\lambda,\varepsilon\right)+\varepsilon\,l_{3}\left(x,y,\lambda,\varepsilon\right)$ (4.7)
$\dot{y}=\varepsilon\left(\pm x\,l_{4}\left(x,y,\lambda,\varepsilon\right)-\lambda\,l_{5}\left(x,y,\lambda,\varepsilon\right)+y\,l_{6}\left(x,y,\lambda,\varepsilon\right)\right)$
where
$l_{3}\left(x,y,\lambda,\varepsilon\right)=\mathcal{O}\left(x,y,\lambda,\varepsilon\right)\quad\text{and}\quad l_{j}\left(x,y,\lambda,\varepsilon\right)=1+\mathcal{O}\left(x,y,\lambda,\varepsilon\right),$
for $j\in\{1,2,4,5,6\}$.
Assume that, for $\varepsilon=0$, there is a slow trajectory connecting the repelling and attracting regions of the critical manifold $\mathcal{C}_{0}$. Then, there exist $\varepsilon_{0}>0$ and $\lambda_{0}>0$ such that for $0<\varepsilon<\varepsilon_{0}$ and $|\lambda|<\lambda_{0}$, the system has an equilibrium point $p\in\mathbb{R}^{2}$ near the origin where $p\to\left(0,0\right)$ as $\left(\lambda,\varepsilon\right)\to\left(0,0\right)$. Then, there exists a smooth function $\lambda_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $\lambda$ that gives rise to a family of canards, asymptotically defined by:
$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\left(B+A\right)\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
and there exists a continuous function $\lambda_{H}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $\lambda$ of Hopf bifurcations in the system, asymptotically defined by:
$\lambda_{H}\left(\sqrt{\varepsilon}\right)=-B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
where
$A=\dfrac{-\dfrac{\partial l_{1}}{\partial x}+3\dfrac{\partial l_{2}}{\partial x}-2\dfrac{\partial l_{4}}{\partial x}+2l_{6}}{8}\qquad B=\dfrac{\dfrac{\partial l_{3}}{\partial x}+l_{6}}{2},$
where the functions $l_{i},\,i=1,\ldots,6$ and their partial derivatives are evaluated at the point $\left(x,y,\lambda,\varepsilon\right)=\left(0,0,0,0\right)$. The Hopf bifurcation is nondegenerate when $A\neq 0$, it is supercritical if $A<0$ and subcritical if $A>0$.
4.2. Case A $(a=0)$
The curves $T_{F}^{2,3}$, $T_{H}^{2,3}$, $DH$ and SNL, introduced in Theorem 3.4 and Remark 3.5, can be parameterised by $c\in[1,+\infty)$. From now on, we restrict our attention to the portions of these curves lying within the region defined by $b,c\geq 1$ – see Figure 6. We are going to denote by $T_{F,+}^{2,3}\backslash\{(1,1)\}$ the connected component of $T_{F}^{2,3}\backslash\{(1,1)\}$ with highest values of $b$ for the same $c$, and $T_{F,-}^{2,3}\backslash\{(1,1)\}$ the other. It is straightforward to verify that, for sufficiently large values of $c$ in $[1,+\infty)$, the following holds:
$T_{F,-}^{2,3}(c)\leq T_{H}^{2,3}(c)\leq DH(c)\leq SNL(c)\leq T_{F,+}^{2,3}.$ (4.8)
All the curves admit $(b,c)=(1,1)$ as accumulation point. Denote by $A_{1}\leq A_{2}<A_{3}<A_{4}$ the limits when $c\to+\infty$ of the previous curves (when it exists), respectively. More specifically, we have:
Figure 6. Partial scheme of the codimension 1 bifurcation curves emerging from the point $(b,c)=(1,1)$ associated with $f_{(0,b,c)}$ of (3.1), for $b,c>0$. The symbol $c^{\infty}$ represents the limit of the associated curves when $c\to+\infty$ (implicit compactification in $c$). The notation of the curves correspond to those of Theorem 3.4, for Case A. The line SNL is dashed because it does not make part of the DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry. Near the curve $b=T_{H}^{2,3}$, the emerging periodic solutions give rise to canards as a consequence of Proposition 4.9.
Lemma 4.8.
As illustrated in Figure 6, the following equalities hold:
  1. (1)
    $A_{1}=\displaystyle\lim_{c\to+\infty}T_{F,-}^{2,3}(c)=3/2$
  2. (2)
    $A_{2}=\displaystyle\lim_{c\to+\infty}T_{H}^{2,3}(c)=3/2$
  3. (3)
    $A_{3}=\displaystyle\lim_{c\to+\infty}DH(c)={17/10}$
  4. (4)
    $A_{4}=\displaystyle\lim_{c\to+\infty}SNL(c)=+\infty$
In particular, we have $1<A_{1}\leq A_{2}<A_{3}<A_{4}.$
We omit the proof since it follows from straightforward computations using the expressions of Theorem 3.4 and Remark 3.5. A partial representation of these curves has been plotted in Figure 6.
Proposition 4.9.
With respect to (4.4), there exists a smooth function $b_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $b$ that gives rise to a family of canards in the system, asymptotically defined by:
$b_{c}\left(\sqrt{\varepsilon}\right)=\frac{3}{2}+\frac{5}{4}\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).$
Proof.
The difference $b-3/2$ of system (4.4) is played by $\lambda$ of Theorem 4.7. The equilibrium $E_{3}$ is a regular singular fold point and there is a Hopf bifurcation point at $b=3/2$. In order to apply Theorem 4.7 we change variables and the parameter by
$\left\{\begin{array}[]{l}\bar{x}=x-1\\ \\ \bar{y}=y+2/3\\ \\ \lambda=b-3/2\\ \end{array}\right.$
to obtain the equivalent system ($\tau=t/\varepsilon$)
$\left\{\begin{array}[]{l}\dot{\bar{x}}=\bar{y}-\frac{1}{3}\bar{x}^{3}-\bar{x}^{2}\\ \\ \dot{\bar{y}}=\varepsilon\left(-\bar{x}-\lambda\bar{y}+\frac{2}{3}\lambda-\frac{3}{2}\bar{y}\right).\end{array}\right.$
In the notation of Theorem 4.7 we have
$l_{1}=-1\qquad l_{2}=-1-\bar{x}/3\qquad l_{3}=0\qquad l_{4}=1\qquad l_{6}=-\frac{3}{2}$
hence $A=-1/2<0$ and $B=-3/4<0$ therefore the Hopf bifurcation is supercritical for $b$ in the original system (4.9). Since for $b<T_{H}^{2,3}(c)$ the eigenvalues of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,$ are negative, then the bifurcating periodic solution is unstable and canards occur when
$b_{c}(\sqrt{\varepsilon})=\frac{3}{2}+\frac{5}{4}\varepsilon+\mathcal{O}(\varepsilon^{3/2}).$
4.3. Digestive remark
We describe the dynamics of (3.1) in the case $a=0$, focusing on the first quadrant of the bifurcation diagram in the $(b,c)$–plane. We suggest the reader follows the description by using Figure 6.
For $b\in(0,1)$ and $c>1$, there exists a unique equilibrium point $E_{1}$, which is unstable. This source is enclosed by a stable periodic orbit denoted by $C^{s}$. At $b=1$, one of the eigenvalues of the source $E_{1}$ becomes zero, leading to the creation of two additional unstable nodes $E_{2}$ and $E_{3}$. $E_{1}$ undergoes a supercritical Pitchfork bifurcation.
The saddles $E_{2}$ and $E_{3}$ subsequently become unstable foci and undergo a supercritical Hopf bifurcation, giving rise to two unstable periodic solutions (symmetric under $\kappa$). Around $b=DH(c)$, the invariant manifolds of $E_{1}$ evolve into a double homoclinic orbit, which is unstable, previously identified in [8].
As $b$ increases smoothly, the double homoclinic loop breaks apart, resulting in an unstable periodic orbit $C^{u}$, which surrounds both the stable and unstable manifolds of $E_{1}$. Near $b=SNL(c)$, the periodic orbits $C^{s}$ and $C^{u}$ collapse, and two stable equilibria remain, dominating the basin of attraction of any compact subset of $\mathbb{R}^{2}$. For $c\gg 1$, there is a small portion of the bifurcation parameter diagram $(b,c)$ of Figure 6, where the periodic orbits emerging from $T_{H}^{2,3}$ correspond to canards. They rapidly (with respect to $b$) collapse into the double homoclinic cycle to $E_{1}$. The values of $b_{c}$ that give rise to canards are at most $\mathcal{O}\left(\varepsilon\right)$ away from the values $b_{H}$ where a Hopf bifurcation occurs (see Figure 7):
$b_{c}-b_{H}=\frac{1}{2}\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).$
Figure 7. Sketch of the singular bifurcations for $b\geq 1$ and $c=c^{\infty}$, in Case A. The smaller $\varepsilon$ is, the narrower the interval of $[b_{c},b_{H}]$. When $\varepsilon=0$, we have $b_{c}\equiv b_{H}$.
4.4. Case B $(b=0)$
Proposition 4.10.
With respect to (4.4), there exists a continuous function $a_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $a$ that gives rise to a canard in the system, asymptotically defined by:
$a_{c}\left(\sqrt{\varepsilon}\right)=-\frac{\varepsilon}{8}+\mathcal{O}\left(\varepsilon^{3/2}\right).$
Proof.
The difference $a-1$ of system (4.4) is played by $\lambda$ in Theorem 4.7. We are interested in only one of the two fold points of (3.1) located at $(1,-2/3)$. After shifting the fold point to the origin, reversing time $t\to-t$ and setting $\lambda=a-1$, we get
$\left\{\begin{array}[]{l}\dot{x}=-y+x^{2}(1+x/3),\\ \\ \dot{y}=\varepsilon(x-\lambda).\end{array}\right.$
In this standard form, the relevant parameters $\ell_{i}$ defined in Theorem 4.7 are
$\ell_{1}=-1,\quad\ell_{2}=1+x/3,\quad\ell_{3}=0,\quad\ell_{4}=1,\quad\ell_{5}=1,\quad\ell_{6}=0,\quad A=1.$
Hence, a maximal canard exists on a curve $\lambda_{c}$ in $(\varepsilon,\lambda)$–space defined by:
$a_{c}(\sqrt{\varepsilon})=-\frac{\varepsilon}{8}+\mathcal{O}(\varepsilon^{3/2}).$
4.5. Digestive remark
The dynamics of system (3.1) for $b=0$ are illustrated in Figures 4 and 9. Our analysis focuses on the first and fourth quadrants of the bifurcation diagram in the $(a,c)$–plane. For $a<-1$, the saddle $E_{1}$ becomes a stable node. Near $a=-1$, a Belyakov transition occurs, followed by a supercritical Hopf bifurcation that gives rise to a stable periodic orbit that involves into a canard in an exponentially small portion of the bifurcation plane. This periodic solution disappears through a subcritical Hopf bifurcation at $a=1$.
5. The phase portraits in the Poincaré disc
In this section, we study the global phase portraits for system (3.1) in the Poincaré disc by using the theory described in Subsections 2.2 and 2.3.
Theorem 5.1.
The global phase portraits of (3.1) in the Poincaré disc are topologically equivalent to one of the phase portraits of Figures 8, 9 and 10 for parameters values in Cases A, B and C, respectively.
Remark 5.2.
The calculations used in the proof of Theorem 5.1 could be streamlined and simplified by using weighted or quasi-homogeneous blow-ups. By a pedagogical point of view, we decided to do three blow-ups, which are illustrated in Figures 8, 9 and 10 for parameters values in Cases A, B and C. Some computations in the following proof were carried out using the software Maple.
Proof.
System (3.1) in the local chart $U_{1}=\{z_{1}>0\}$ may be written as:
$\dot{u}=$ $-cu^{2}v^{2}-\frac{b}{c}uv^{2}-cuv^{2}+\frac{a}{c}v^{3}-\frac{1}{c}v^{2}+\frac{c}{3}u,$ (5.1)
$\dot{v}=$ $-cuv^{3}-cv^{3}+\frac{c}{3}v.$
The origin is the unique equilibrium on $v=0$ (the points of $\mathbb{S}^{1}$ in any chart have $v=0.$) The eigenvalues of Jacobian matrix associated with the vector field (5.1) are $\dfrac{c}{3},\dfrac{c}{3}.$ So, the origin is an unstable node (resp. stable) for $c>0$ (resp. $c<0$.)
System (3.1) in the local chart $U_{2}$ may be written as:
$\dot{u}=$ $\frac{1}{c}u^{2}v^{2}-\frac{a}{c}uv^{3}-\frac{c}{3}u^{3}+\frac{b}{c}uv^{2}+cuv^{2}+cv^{2},$ (5.2)
$\dot{v}=$ $\frac{1}{c}uv^{3}-\frac{a}{c}v^{4}+\frac{b}{c}v^{3}.$
The origin is an identically zero singular point. By doing the blow-up $(u,v)\rightarrow(u_{1},w)$ where $w=\dfrac{v_{1}}{u_{1}},$ we get
$\dot{u}_{1}$ $=-\frac{a}{c}{u_{1}}^{4}{v_{1}}^{3}+\frac{1}{c}{u_{1}}^{4}{v_{1}}^{2}+c{u_{1}}^{3}{v_{1}}^{2}+\frac{b}{c}{u_{1}}^{3}{v_{1}}^{2}+c{u_{1}}^{2}{v_{1}}^{2}-\frac{c}{3}{u_{1}}^{3},$ (5.3)
$\dot{v}_{1}$ $=-c{u_{1}}^{2}{v_{1}}^{3}-cu_{1}{v_{1}}^{3}+\frac{c}{3}{u_{1}}^{2}{v_{1}}.$
By removing the common factor $u_{1},$ we have
$\dot{u}_{1}$ $=-\frac{a}{c}{u_{1}}^{3}{v_{1}}^{3}+\frac{1}{c}{u_{1}}^{3}{v_{1}}^{2}+c{u_{1}}^{2}{v_{1}}^{2}+\frac{b}{c}{u_{1}}^{2}{v_{1}}^{2}+c{u_{1}}{v_{1}}^{2}-\frac{c}{3}{u_{1}}^{2},$ (5.4)
$\dot{v}_{1}$ $=-c{u_{1}}{v_{1}}^{3}-c{v_{1}}^{3}+\frac{c}{3}{u_{1}}{v_{1}}.$
The origin is an identically zero singular point again. For studying the local phase portrait at the origin, we do another blow-up (see §2.2 of the present paper). The characteristic direction at the origin is the real factors of $\frac{2c}{3}{u_{1}}^{2}v_{1}.$ Then the vertical axis $u_{1}=0$ is a characteristic direction. So, we translate the direction $u_{1}=0$ to the direction $u_{1}=v_{1}$ doing the change of variables $(u_{1},v_{1})=(u_{2}-v_{2},v_{2}).$ Then, the new system is given by
$\dot{u_{2}}=$ $\dfrac{1}{3c}\Big{(}3c^{2}u_{2}v_{2}-c^{2}u_{2}^{2}-2c^{2}v_{2}^{2}+3c^{2}u2v_{2}^{2}+3bu_{2}^{2}v_{2}^{2}+3c^{2}u_{2}^{2}v_{2}^{2}+3u_{2}^{3}v_{2}^{2}$ (5.5)
$\qquad-6c^{2}v_{2}^{3}-6bu_{2}v_{2}^{3}-9c^{2}u_{2}v_{2}^{3}-9u_{2}^{2}v_{2}^{3}-3au_{2}^{3}v_{2}^{3}+3bv_{2}^{4}+6c^{2}v_{2}^{4}$
$\qquad+9u_{2}v_{2}^{4}+9au_{2}^{2}v_{2}^{4}-3v_{2}^{5}-9au_{2}v_{2}^{5}+3av_{2}^{6}\Big{)},$
$\dot{v_{2}}=$ $\dfrac{c}{3}v_{2}\left(u_{2}-v_{2}-3v_{2}^{2}-3u_{2}v_{2}^{2}+3v_{2}^{3}\right).$
The origin is the only singular point of system above on the line $u_{2}=0,$ and since the linear part is identically zero, we do the vertical blow-up $(u_{2},v_{2})\rightarrow(u_{3},w)$ where $w=\dfrac{v_{3}}{u_{3}},$ and we have
$\dot{u}_{3}=$ $\dfrac{1}{3c}u_{3}^{2}\Big{(}3c^{2}v_{3}-c^{2}-2c^{2}v_{3}^{2}+3c^{2}u_{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{2}+3c^{2}u_{3}^{2}v_{3}^{2}+3u_{3}^{3}v_{3}^{2}-6c^{2}u_{3}v_{3}^{3}$
$\qquad\,-6bu_{3}^{2}v_{3}^{3}-9c^{2}u_{3}^{2}v_{3}^{3}-9u_{3}^{3}v_{3}^{3}-3au_{3}^{4}v_{3}^{3}+3bu_{3}^{2}v_{3}^{4}+6c^{2}u_{3}^{2}v_{3}^{4}+9u_{3}^{3}v_{3}^{4}$
$\qquad\,+9au_{3}^{4}v_{3}^{4}-3u_{3}^{3}v_{3}^{5}-9au_{3}^{4}v_{3}^{5}+3au_{3}^{4}v_{3}^{6}\Big{)},$ (5.6)
$\dot{v}_{3}=$ $\dfrac{1}{3c}u_{3}v_{3}(1-v_{3})\Big{(}2c^{2}-2c^{2}v_{3}-6c^{2}u_{3}v_{3}^{2}-3bu_{3}^{2}v_{3}^{2}-6c^{2}u_{3}^{2}v_{3}^{2}-3u_{3}^{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{3}$
$\hskip 79.6678pt+6c^{2}u_{3}^{2}v_{3}^{3}+6u_{3}^{3}v_{3}^{3}+3au_{3}^{4}v_{3}^{3}-3u_{3}^{3}v_{3}^{4}-6au_{3}^{4}v_{3}^{4}+3au_{3}^{4}v_{3}^{5}\Big{)}.$
Eliminating the common factor $u_{3}$ leads to
$\dot{u}_{3}=$ $\dfrac{1}{3c}u_{3}\Big{(}3c^{2}v_{3}-c^{2}-2c^{2}v_{3}^{2}+3c^{2}u_{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{2}+3c^{2}u_{3}^{2}v_{3}^{2}+3u_{3}^{3}v_{3}^{2}-6c^{2}u_{3}v_{3}^{3}$
$\qquad\,-6bu_{3}^{2}v_{3}^{3}-9c^{2}u_{3}^{2}v_{3}^{3}-9u_{3}^{3}v_{3}^{3}-3au_{3}^{4}v_{3}^{3}+3bu_{3}^{2}v_{3}^{4}+6c^{2}u_{3}^{2}v_{3}^{4}+9u_{3}^{3}v_{3}^{4}$
$\qquad\,+9au_{3}^{4}v_{3}^{4}-3u_{3}^{3}v_{3}^{5}-9au_{3}^{4}v_{3}^{5}+3au_{3}^{4}v_{3}^{6}\Big{)},$ (5.7)
$\dot{v}_{3}=$ $\dfrac{1}{3c}v_{3}(1-v_{3})\Big{(}2c^{2}-2c^{2}v_{3}-6c^{2}u_{3}v_{3}^{2}-3bu_{3}^{2}v_{3}^{2}-6c^{2}u_{3}^{2}v_{3}^{2}-3u_{3}^{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{3}$
$\hskip 79.6678pt+6c^{2}u_{3}^{2}v_{3}^{3}+6u_{3}^{3}v_{3}^{3}+3au_{3}^{4}v_{3}^{3}-3u_{3}^{3}v_{3}^{4}-6au_{3}^{4}v_{3}^{4}+3au_{3}^{4}v_{3}^{5}\Big{)}.$
This system has two singular points on the line $u_{3}=0,$ namely $\tilde{E}_{1}:(0,0),$ and $\tilde{E}_{2}:(0,1).$ The eigenvalues of the Jacobian matrix associated with the vector field (5.7) at $(0,0)$ are $-\dfrac{c}{3},\dfrac{2c}{3}.$ Hence, $\tilde{E}_{1}$ is a saddle point. The linear part of the system at $\tilde{E}_{2}$ is identically zero. Therefore, in order to know the local phase portrait around this point, we must do blow-up. At first, we translate the singular point $\tilde{E}_{2}$ at the origin by doing the change of variables $(u_{3},v_{3})=(u_{4},1+v_{4}).$ System (5.7) in the new variables becomes
$\dot{u}_{4}=$ $\dfrac{1}{3c}u_{4}\Big{(}-3c^{2}u_{4}-c^{2}v_{4}-12c^{2}u_{4}v_{4}+3c^{2}u_{4}^{2}v_{4}-2c^{2}v_{4}^{2}-15c^{2}u_{4}v_{4}^{2}+3bu_{4}^{2}v_{4}^{2}$
$\qquad\quad+12c^{2}u_{4}^{2}v_{4}^{2}-6c^{2}u_{4}v_{4}^{3}+6bu_{4}^{2}v_{4}^{3}+15c^{2}u_{4}^{2}v_{4}^{3}-3u_{4}3v_{4}^{3}+3au_{4}^{4}v_{4}^{3}$
$\qquad\quad+3bu_{4}^{2}v_{4}^{4}+6c^{2}u_{4}^{2}v_{4}^{4}-6u_{4}^{3}v_{4}^{4}+9au_{4}^{4}v_{4}^{4}-3u_{4}^{3}v_{4}^{5}+9au_{4}^{4}v_{4}^{5}+3au_{4}^{4}v_{4}^{6}\Big{)},$
$\dot{u}_{4}=$ $-\dfrac{1}{3c}v_{4}(1+v_{4})\Big{(}-6c^{2}u_{4}-2c^{2}v_{4}-12c^{2}u_{4}v_{4}+3bu_{4}^{2}v_{4}+6c^{2}u_{4}^{2}v_{4}-6c^{2}u_{4}v_{4}^{2}$
$\qquad\qquad\qquad\quad+6bu_{4}^{2}v_{4}^{2}+12c^{2}u_{4}^{2}v_{4}^{2}-3u_{4}^{3}v_{4}^{2}+3au_{4}^{4}v_{4}^{2}+3bu_{4}^{2}v_{4}^{3}+6c^{2}u_{4}^{2}v_{4}^{3}$
$\hskip 85.35826pt-6u_{4}^{3}v_{4}^{3}+9au_{4}^{4}v_{4}^{3}-3u_{4}^{3}v_{4}^{4}+9au_{4}^{4}v_{4}^{4}+3au_{4}^{4}v_{4}^{5}\Big{)}.$ (5.8)
The origin is the unique singular point of system above on the line $u_{4}=0$ and furthermore, the characteristic direction at the origin of system above is $u_{4}=0.$ Hence, by using the change of variables $(u_{4},v_{4})=(u_{5}-v_{5},v_{5}),$ we translate the direction to $u_{5}=v_{5},$ and we have
$-3c\dot{u_{5}}=$ $3{u_{5}}^{4}{v_{5}}^{5}-15{u_{5}}^{3}{v_{5}}^{6}+27{u_{5}}^{2}{v_{5}}^{7}-21{u_{5}}{v_{5}}^{8}+6{v_{5}}^{9}+6{u_{5}}^{4}{v_{5}}^{4}-33{u_{5}}^{3}{v_{5}}^{5}$
$+63{u_{5}}^{2}{v_{5}}^{6}-51{u_{5}}{v_{5}}^{7}+15{v_{5}}^{8}+3{u_{5}}^{4}{v_{5}}^{3}-21{u_{5}}^{3}{v_{5}}^{4}-6c^{2}{u_{5}}^{3}{v_{5}}^{4}$
$-3b{u_{5}}^{3}{v_{5}}^{4}+12b{u_{5}}^{2}{v_{5}}^{5}+45{u_{5}}^{2}{v_{5}}^{5}+24c^{2}{u_{5}}^{2}{v_{5}}^{5}-15b{u_{5}}{v_{5}}^{6}-30c^{2}{u_{5}}{v_{5}}^{6}$
$-39{u_{5}}{v_{5}}^{6}+12c^{2}{v_{5}}^{7}+6b{v_{5}}^{7}+12{v_{5}}^{7}-3{u_{5}}^{3}{v_{5}}^{3}-15c^{2}{u_{5}}^{3}{v_{5}}^{3}-6b{u_{5}}^{3}{v_{5}}^{3}$
$+9{u_{5}}^{2}{v_{5}}^{4}+27b{u_{5}}^{2}{v_{5}}^{4}+63c^{2}{u_{5}}^{2}{v_{5}}^{4}-36b{u_{5}}{v_{5}}^{5}-9{u_{5}}{v_{5}}^{5}-81c^{2}{u_{5}}{v_{5}}^{5}$
$+33c^{2}{v_{5}}^{6}+15b{v_{5}}^{6}+3{v_{5}}^{6}-3b{u_{5}}^{3}{v_{5}}^{2}-12c^{2}{u_{5}}^{3}{v_{5}}^{2}+18b{u_{5}}^{2}{v_{5}}^{3}-27b{u_{5}}{v_{5}}^{4}$
$+60c^{2}{u_{5}}^{2}{v_{5}}^{3}-90c^{2}{u_{5}}{v_{5}}^{4}+42c^{2}{v_{5}}^{5}+12b{v_{5}}^{5}-3c^{2}{u_{5}}^{3}{v_{5}}+30c^{2}{u_{5}}^{2}{v_{5}}^{2}$
$+3b{u_{5}}^{2}{v_{5}}^{2}-69c^{2}{u_{5}}{v_{5}}^{3}-6b{u_{5}}{v_{5}}^{3}+42c^{2}{v_{5}}^{4}+3b{v_{5}}^{4}+12c^{2}{u_{5}}^{2}{v_{5}}$
$-40c^{2}{u_{5}}{v_{5}}^{2}+26c^{2}{v_{5}}^{3}+3c^{2}{u_{5}}^{2}-11c^{2}{u_{5}}{v_{5}}+6c^{2}{v_{5}}^{2},$ (5.9)
$-3c\dot{v_{5}}=$ ${v_{5}}(1+{v_{5}})(-3{u_{5}}^{3}{v_{5}}^{4}+9{u_{5}}^{2}{v_{5}}^{5}-9{u_{5}}{v_{5}}^{6}+3{v_{5}}^{7}-6{u_{5}}^{3}{v_{5}}^{3}+18{u_{5}}^{2}{v_{5}}^{4}$
$-18{u_{5}}{v_{5}}^{5}+6{v_{5}}^{6}-3{u_{5}}^{3}{v_{5}}^{2}+3b{u_{5}}^{2}{v_{5}}^{3}+6c^{2}{u_{5}}^{2}{v_{5}}^{3}+9{u_{5}}^{2}{v_{5}}^{3}-6b{u_{5}}{v_{5}}^{4}$
$-12c^{2}{u_{5}}{v_{5}}^{4}-9{u_{5}}{v_{5}}^{4}+3b{v_{5}}^{5}+6c^{2}{v_{5}}^{5}+3{v_{5}}^{5}+6b{u_{5}}^{2}{v_{5}}^{2}+12c^{2}{u_{5}}^{2}{v_{5}}^{2}$
$-24c^{2}{u_{5}}{v_{5}}^{3}-12b{u_{5}}{v_{5}}^{3}+12c^{2}{v_{5}}^{4}+6b{v_{5}}^{4}+6c^{2}{u_{5}}^{2}{v_{5}}+3b{u_{5}}^{2}{v_{5}}-18c^{2}{u_{5}}{v_{5}}^{2}$
$-6b{u_{5}}{v_{5}}^{2}+12c^{2}{v_{5}}^{3}+3b{v_{5}}^{3}-12c^{2}{u_{5}}{v_{5}}+12c^{2}{v_{5}}^{2}-6c^{2}{u_{5}}+4c^{2}{v_{5}}).$
Since the linear part is identically zero, we do the vertical blow-up $(u_{5},v_{5})=(u_{6},u_{6}v_{6}),$ for investigating the local phase portrait. The new system writes (applying the blow-up):
$-3c\dot{u}_{6}=$ ${u_{6}}^{2}(3c^{2}-11c^{2}{v_{6}}+12c^{2}{u_{6}}{v_{6}}-3c^{2}{u_{6}}^{2}{v_{6}}+6c^{2}{v_{6}}^{2}-40c^{2}{u_{6}}{v_{6}}^{2}+3b{u_{6}}^{2}{v_{6}}^{2}$
$+30c^{2}{u_{6}}^{2}{v_{6}}^{2}-3b{u_{6}}^{3}{v_{6}}^{2}-12c^{2}{u_{6}}^{3}{v_{6}}^{2}+26c^{2}{u_{6}}{v_{6}}^{3}-6b{u_{6}}^{2}{v_{6}}^{3}-69c^{2}{u_{6}}^{2}{v_{6}}^{3}$
$+18b{u_{6}}^{3}{v_{6}}^{3}+60c^{2}{u_{6}}^{3}{v_{6}}^{3}-3{u_{6}}^{4}{v_{6}}^{3}-6b{u_{6}}^{4}{v_{6}}^{3}-15c^{2}{u_{6}}^{4}{v_{6}}^{3}+3{u_{6}}^{5}{v_{6}}^{3}$
$+3b{u_{6}}^{2}{v_{6}}^{4}+42c^{2}{u_{6}}^{2}{v_{6}}^{4}-27b{u_{6}}^{3}{v_{6}}^{4}-90c^{2}{u_{6}}^{3}{v_{6}}^{4}+9{u_{6}}^{4}{v_{6}}^{4}+27b{u_{6}}^{4}{v_{6}}^{4}$
$+63c^{2}{u_{6}}^{4}{v_{6}}^{4}-21{u_{6}}^{5}{v_{6}}^{4}-3b{u_{6}}^{5}{v_{6}}^{4}-6c^{2}{u_{6}}^{5}{v_{6}}^{4}+6{u_{6}}^{6}{v_{6}}^{4}+12b{u_{6}}^{3}{v_{6}}^{5}$
$+42c^{2}{u_{6}}^{3}{v_{6}}^{5}-9{u_{6}}^{4}{v_{6}}^{5}-36b{u_{6}}^{4}{v_{6}}^{5}-81c^{2}{u_{6}}^{4}{v_{6}}^{5}+45{u_{6}}^{5}{v_{6}}^{5}+12b{u_{6}}^{5}{v_{6}}^{5}$
$+24c^{2}{u_{6}}^{5}{v_{6}}^{5}-33{u_{6}}^{6}{v_{6}}^{5}+3{u_{6}}^{7}{v_{6}}^{5}+3{u_{6}}^{4}{v_{6}}^{6}+15b{u_{6}}^{4}{v_{6}}^{6}+33c^{2}{u_{6}}^{4}{v_{6}}^{6}$
$-39{u_{6}}^{5}{v_{6}}^{6}-15b{u_{6}}^{5}{v_{6}}^{6}-30c^{2}{u_{6}}^{5}{v_{6}}^{6}+63{u_{6}}^{6}{v_{6}}^{6}-15{u_{6}}^{7}{v_{6}}^{6}+12{u_{6}}^{5}{v_{6}}^{7}$
$+6b{u_{6}}^{5}{v_{6}}^{7}+12c^{2}{u_{6}}^{5}{v_{6}}^{7}-51{u_{6}}^{6}{v_{6}}^{7}+27{u_{6}}^{7}{v_{6}}^{7}+15{u_{6}}^{6}{v_{6}}^{8}-21{u_{6}}^{7}{v_{6}}^{8}$
$+6{u_{6}}^{7}{v_{6}}^{9}),$ (5.10)
$3c\dot{v}_{6}=$ ${u_{6}}{v_{6}}({v_{6}}-1)(-9c^{2}+6c^{2}{v_{6}}-30c^{2}{u_{6}}{v_{6}}+3b{u_{6}}^{2}{v_{6}}+9c^{2}{u_{6}}^{2}{v_{6}}+26c^{2}{u_{6}}{v_{6}}^{2}$
$-6b{u_{6}}^{2}{v_{6}}^{2}-51c^{2}{u_{6}}^{2}{v_{6}}^{2}+12b{u_{6}}^{3}{v_{6}}^{2}+30c^{2}{u_{6}}^{3}{v_{6}}^{2}-3{u_{6}}^{4}{v_{6}}^{2}+3b{u_{6}}^{2}{v_{6}}^{3}$
$+42c^{2}{u_{6}}^{2}{v_{6}}^{3}-24b{u_{6}}^{3}{v_{6}}^{3}-72c^{2}{u_{6}}^{3}{v_{6}}^{3}+9{u_{6}}^{4}{v_{6}}^{3}+15b{u_{6}}^{4}{v_{6}}^{3}+33c^{2}{u_{6}}^{4}{v_{6}}^{3}$
$-12{u_{6}}^{5}{v_{6}}^{3}+12b{u_{6}}^{3}{v_{6}}^{4}+42c^{2}{u_{6}}^{3}{v_{6}}^{4}-9{u_{6}}^{4}{v_{6}}^{4}-30b{u_{6}}^{4}{v_{6}}^{4}-66c^{2}{u_{6}}^{4}{v_{6}}^{4}$
$+36{u_{6}}^{5}{v_{6}}^{4}+6b{u_{6}}^{5}{v_{6}}^{4}+12c^{2}{u_{6}}^{5}{v_{6}}^{4}-15{u_{6}}^{6}{v_{6}}^{4}+3{u_{6}}^{4}{v_{6}}^{5}+15b{u_{6}}^{4}{v_{6}}^{5}$
$+33c^{2}{u_{6}}^{4}{v_{6}}^{5}-36{u_{6}}^{5}{v_{6}}^{5}-12b{u_{6}}^{5}{v_{6}}^{5}-24c^{2}{u_{6}}^{5}{v_{6}}^{5}+45{u_{6}}^{6}{v_{6}}^{5}-6{u_{6}}^{7}{v_{6}}^{5}$
$+12{u_{6}}^{5}{v_{6}}^{6}+6b{u_{6}}^{5}{v_{6}}^{6}+12c^{2}{u_{6}}^{5}{v_{6}}^{6}-45{u_{6}}^{6}{v_{6}}^{6}+18{u_{6}}^{7}{v_{6}}^{6}+15{u_{6}}^{6}{v_{6}}^{7}$
$-18{u_{6}}^{7}{v_{6}}^{7}+6{u_{6}}^{7}{v_{6}}^{8}).$
Now, by elimination of the common factor $u_{6},$ we obtain the truncated system
$\dot{u_{6}}=$ $-\dfrac{1}{3c}u_{6}\left(3c^{2}-11c^{2}v_{6}+\mathcal{O}(|u_{6},v_{6}|^{2})\right),$ (5.11)
$\dot{v_{6}}=$ $\dfrac{1}{3c}v_{6}(v_{6}-1)\left(-9c^{2}+6c^{2}v_{6}+\mathcal{O}(|u_{6},v_{6}|^{2})\right).$
The equilibrium points for system (5.11) on the line $u_{6}=0$ are given by
$\tilde{E}_{4}=(0,0),\qquad\tilde{E}_{5}=(0,1),\qquad\tilde{E}_{6}=\left(0,\frac{3}{2}\right).$ (5.12)
The eigenvalues of Jacobian matrix associated with vector field (5.11) at $\tilde{E}_{4}$ and $\tilde{E}_{5}$ are $-c$, $3c$ and $\frac{2c}{3},-c$. So, they are saddle points. On the other hand, the eigenvalues associated with Jacobian matrix at $\tilde{E}_{6}$ are 00 and $\frac{3c}{2}$, and it is a semi-hyperbolic equilibrium. Given the center manifold $\dot{u_{6}}=-\frac{b}{8c}{u_{6}}^{3}$, we conclude that the equilibrium point $\tilde{E}_{6}$ is:
  • 1.
    a saddle point for
    $\{(a,b,c)\big{|}b>0,\,c<0\}\cup\{(a,b,c)\big{|}b>0,\,c>0\}.$ (5.13)
  • 2.
    an unstable node for
    $\{(a,b,c)\big{|}b<0,\,c<0\}.$ (5.14)
  • 3.
    a stable node for
    $\{(a,b,c)\big{|}b<0,\,c>0\}\cup\{(a,b,c)\big{|}b=0\}.$ (5.15)
The final step of the proof follows from the description of Subsection 5.1. ∎
(a) Region a
(b) Region b
(c) Region c
(d) Region d
(e) Region e
(f) Region f
(g) Region g
(h) Region h
(i) Region i
(j) Region j
(k) Region k
(l) Region l
(m) Region m
(n) Region n
(o) Region o
(p) Region p
Figure 8. The global phase portraits of (3.1) for Case A, which is associated with system (3.1) when $a=0,$ associated with regions a–j of Figure (3)(a).
(a) Region I
(b) Region II
(c) Region III
(d) Region IV
Figure 9. The global phase portraits of (3.1) for Case B, which is associated with system (3.1) when $b=0,$ corresponding to regions I–IV of Figure 4(b), in the case where the eigenvalues of (3.5) at $E_{1}$ are complex non-real.
5.1. Description of Figures 11, 12, 13, 14, 15 and 16 in Appendix 2
In order to plot the complete phase diagram in the compactified space for the Cases A, B and C in Figures 8, 9 and 10, we need to carefully go back through each step of the blow-up procedure, explaining the transformations (blow-ups, directional transformations and coordinate maps) used at each stage, and clearly stating the resulting system at every step.
Undoing the rescaling $dt_{3}=u_{6}dt_{2}$ the phase portrait depicted in Figures 11(a), 12(a), 13(a), 14(a), 15(a) and 16(a) yields the local phase portrait at the origin of (5.10), which is topologically equivalent to that of Figures 11(b), 12(b), 13(b), 14(b), 15(b) and 16(b), respectively.
Going back through the change of variables $(u_{5},v_{5})=(u_{6},u_{6}v_{6}),$ the phase portrait depicted in Figures 11(b), 12(b), 13(b), 14(b), 15(b) and 16(b) yields the local phase portrait at the origin of (5.9), which is topologically equivalent to that of Figures 11(c), 12(c), 13(c), (14)(c), 15(c) and 16(c), respectively.
Going back through the change of variables $(u_{4},v_{4})=(u_{5}-v_{5},v_{5}),$ the phase portrait depicted in Figures 11(c), 12(c), 13(c), 14(c), 15(c) and 16(c), yields the local phase portrait at the origin of (5.8), which is topologically equivalent to that of Figures 11(d), 12(d), 13(d), 14(d), 15(d) 16(d), respectively.
Going back through the change of variables $(u_{3},v_{3})=(u_{4},1+v_{4})$ the phase portrait depicted in Figures 11(d), 12(d), 13(d), 14(d), 15(d) and 16(d) yields the local phase portrait at the origin of system (5.7), which is topologically equivalent to that of Figures 11(e), 12(e), 13(e), 14(e), 15(e) and 16(e), respectively.
Undoing the rescaling $dt_{2}=u_{3}dt_{1}$ the phase portrait depicted in Figures 11(e), 12(e), 13(e), 14(e), 15(e) and 16(e) yields the local phase portrait at the origin of (5.6), which is topologically equivalent to that of Figures 11(f) and 12(f), 13(f), 14(f), 15(f) and 16(f), respectively.
Going back through the change of variables $(u_{2},v_{2})\rightarrow(u_{3},v_{3}/u_{3}),$ the phase portrait depicted in Figures 11(f), 12(f), 13(f), 14(f), 15(f) and 16(f), yields the local phase portrait at the origin of system 5.5, which is topologically equivalent to that of Figures 11(g), 12(g), 13(g), 14(g), 15(g) and 16(g), respectively.
Going back through the change of variables $(u_{1},v_{1})=(u_{2}-v_{2},v_{2}),$ the phase portrait depicted in Figures 11(g), 12(g), 13(g), 14(g), 15(g) and 16(g) yields the local phase portrait at the origin of (5.4), which is topologically equivalent to that of Figures 11(h), 12(h), 13(h), 14(h), 15(h) and 16(h), respectively.
Undoing the rescaling $dt_{1}=u_{1}dt$ the phase portrait depicted in Figures 11(h), 12(h), 13(h), 14(h), 15(h) and 16(h) yields the local phase portrait at the origin of (5.3), which is topologically equivalent to that of Figures 11(i), 12(i), 13(i), 14(i), 15(i) and 16(i), respectively.
Going back through the change of variables $(u,v)\rightarrow(u_{1},v_{1}/u_{1}),$ the phase portrait depicted in Figures 11(i), 12(i), 13(i), 14(i), 15(i) and 16(i) yields the local phase portrait at the origin of (5.2), which is topologically equivalent to that of Figures 11(j), 12(j), 13(j), 14(j), 15(j) and 16(j), respectively.
This allows us to plot the complete phase diagram in the compactified space for the Cases A, B and C in Figures 8, 9 and 10. The local phase portraits of the blow-down at the origin of the different local charts have been plotted in Figures 1116.
(a) $c<-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
(b) $-\sqrt{\frac{b}{1-{x^{*}}^{2}}}<c<0$
(c) $0<c<\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
(d) $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
[…]
6. Discussion and concluding remark
Although the FHN equations were created as a simplified model for nerve impulse, they have also been intensively studied for purely mathematical reasons [12, 13, 10, 23] because they provide a very simple example of equations that exhibit rich dynamics.
Results obtained on [21, pp. 180] synthesised the global bifurcation diagram for the FHN model (3.1). It has been obtained by putting together, as in a “huge puzzle”, all local bifurcation diagrams obtained in previous chapters of the latter reference. The authors have concentrated their attention to a particular parameter region relevant to physiology ($|c|>1+\sqrt{3}$).
Trying to complete and understand the bifurcation diagram of (3.1), in this paper we have discussed the finite equilibria of (3.1), as well as their bifurcations, for the following three scenarios identified in Subsection 1.4:
Case A:
$a=0,$
Case B:
$b=0,$
Case C:
$a\neq 0$ and $10<b<$ 1.
The richest scenario is Case A. In Theorem 3.4 we have found a Double-zero Bifurcation with a $\mathbb{Z}_{2}(\kappa)$–symmetry. We have obtained precise expressions of the bifurcation curves passing through the bifurcation points $(b,c)=(1,\pm 1)$, complementing the work started in [8]. We also give an analytical proof of the results stated in Section 2.4 of [20].
We have been able to explain rigorously the dynamics of Regions 8, 11, 17 and 18 of [21] on the line defined by $a=0$ and $b>0$, as we proceed to explain:
  • Point Q: Pitchfork bifurcation of $E_{1}$ (Theorem 3.4, curve $b=T_{P}$);
  • Region 8: existence of three equilibria, one saddle $E_{1}$ and two sources $E_{2},E_{3}$;
  • Point $\textbf{Q}_{0}$: Hopf bifurcations of $E_{2},E_{3}$ (Theorem 3.4, curve $b=T_{H}^{2,3})$;
  • Region 11: three periodic solutions (one stable, two unstable);
  • Point $\textbf{Q}_{6}$: Double Homoclinic (Theorem 3.4, curve $b={DH}$);
  • Region 17: two periodic solutions of different stabilities;
  • Point T: Saddle-node bifurcation of non-hyperbolic solutions, making part of the Bautin bifurcation of $\textbf{Q}_{17}$ and $\textbf{Q}_{18}$ – see Remark 3.5;
  • Region 18: three equilibria $E_{1},E_{2}$ and $E_{3}$.
The precise location the above points/regions depend on $c$ but their relative position does not.
The non-hyperbolic equilibria corresponding to parameter values situated at Q (see [21]) are degenerated saddle-node (cusp) equilibria, which are attracting for $c<1$ and repelling for $c>1$. For $c=1$, following the Remark 3.5, this point might correspond to a degenerated Bogdanov-Takens of order two (with symmetry) – a kind of codimension-three bifurcation. As pointed out in [23], this would be the unique generic codimension-three local bifurcation exhibited by the FHN model (3.1) and its complete understanding is an open problem. Using the same line of argument, in Cases B and C, we have analysed the dynamics of (3.1).
Based on [10, 14], in Section 4, we have studied the asymptotic case $c\to\pm\infty$, where canards are detected. An important consideration is the empirical difficulty in finding canards, since the smaller $\varepsilon$ is, the narrower the interval of $b$ (Case A) or $a$ (Case B) values for which canards appear – see Figure 7. The dynamics produced in the system around these small intervals is called the canard explosion and has been studied in [10, 14, 12].
In Section 5, we have continued the analysis with the compactification of the phase portraits associated with (3.1) on the Poincaré disc. This brings additional information of the trajectories which tend to or come from infinity. We have provided phase and bifurcation diagrams in all the three cases.
Novelty versus limitations
We have considered several cases of the classical FHN system which allow the catalogue of topologically distinct phase portraits. The global dynamics seems to be determined by the local bifurcations found in Theorems 3.4, 3.8 and Proposition 3.10. Although we have not achieved complete phase portraits for all cases, we have described all topological regions worth to be analysed. Numerics suggest that, in the three scenarios under consideration, we have obtained a complete phase portrait (up to conjugacy). However, we are not claiming it analytically.
We have connected the local bifurcation theory (Theorems 3.4, 3.8 and Proposition 3.10) with the asymptotic dynamics $|c|\to+\infty$ (Propositions 4.9 and 4.10). Finally, using a blow-up technique we have depicted the dynamics of (3.1) at infinity (Theorem 5.1). Our contributions do not finish the whole discussion of the bifurcation analysis of (3.1); bifurcations might make part of of high codimension phenomena.
Theoretically, the maximum number of non-trivial periodic solutions of (3.1) is three [21, pp. 218]. There are regions where the dynamics is completely determined and the phase portrait is complete (Regions b, g, o and p of Figure 8 combined with Lemma 3.1) and others in which the complete analysis of the bifurcation diagram of [21, pp. 180] is still ongoing. We defer the complete analysis for a future task.
Acknowledgments
Alexandre Rodrigues has been supported by the Project CEMAPRE/REM (UIDB/05069/ 2020) financed by FCT/MCTES through national funds. Nasrin Sadri is supported by Institute for Research in Fundamental Sciences (IPM) grant 1403370046. This work is based upon research funded by Iran National Science Foundation (INSF) under Project Number 4041099. The authors are also indebted to two reviewers for the suggestions which helped to improve the readability of this manuscript.
Appendix A Blowing-down of the origin of the local chart $U_{2}$:
Figure 11. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b>0$ and $c<0.$
Figure 12. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b<0$ and $c<0.$
Figure 13. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ when $b=0,$ $c<0.$
Figure 14. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ when $b=0,$ $c>0.$
Figure 15. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b>0$ and $c>0.$
Figure 16. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b<0$ and $c>0.$
[Uncaptioned image] Region 1 [Uncaptioned image] Region 2 [Uncaptioned image] Region 3 [Uncaptioned image] Region 4 [Uncaptioned image] Region 5 [Uncaptioned image] Region 6 [Uncaptioned image] Region 7
[Uncaptioned image] Region 8 [Uncaptioned image] Region 9 [Uncaptioned image] Region 10 [Uncaptioned image] Region 11 [Uncaptioned image] Region 12 [Uncaptioned image] Region 13 [Uncaptioned image] Region 14
[Uncaptioned image] Region 15 [Uncaptioned image] Region 16 [Uncaptioned image] Region 17 [Uncaptioned image] Region 18 [Uncaptioned image] Region 19 [Uncaptioned image] Region 20     [Uncaptioned image] Region 21
[Uncaptioned image] Region 22 [Uncaptioned image] Region 23 [Uncaptioned image] Region 24 [Uncaptioned image] Region 25 [Uncaptioned image] Region 26 [Uncaptioned image] Region 27 [Uncaptioned image] Region 28
Table 1. The local phase portraits corresponding to Case A, which is associated with system (3.1) when $a=0,$ for Regions 1–28 of Figure (3)(b).
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Region 1 Region 2 Region 3 Region 4
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Region 5 Region 6 Region 7 Region 8
Table 2. The local phase portraits corresponding to Case B, associated with system (3.1) when $b=0$, for Regions 1–8 of Figure 4(a).

### v3

@@
-FitzHugh-Nagumo equation: from global dynamics to the slow-fast system
+slow-fast system and dynamics near infinity
 Abstract.
-We concentrate our attention on the qualitative study of the phase portraits of the three-parameter FitzHugh-Nagumo family and its compactification.
-We divide the study into three scenarios based on the parameters.
-One of the scenarios is characterised by the existence of a Double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry (singularity of codimension two).
-In this case, we explicitly exhibit the Pitchfork, Hopf, Belyakov, Double Homoclinic bifurcation/transition curves unfolding the singularity of codimension 2 and plot the bifurcation diagrams.
-We bridge this analysis with the theory on the associated slow-fast family and the existence of canards.
-We complete our study with the global compactification of the phase portraits for the family under consideration.
+We focus on the qualitative analysis of the phase portraits arising in the three-parameter FitzHugh-Nagumo system and its compactified form.
+The investigation is split into three parameter-dependent cases.
+In one of these cases, the system displays a double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry, a singularity of codimension two.
+For this case, we provide explicit descriptions of the bifurcation/transition curves unfolding the singularity, including Pitchfork, Hopf, Belyakov, and Double Homoclinic.
+Further, we present the corresponding bifurcation diagrams.
+We bridge this analysis with the theory on the framework of slow-fast family and to the presence of canards.
+We complete our study with the analysis of the dynamics “near infinity” for the family under consideration.
 This study complements the work summarised in Georgescu, Rocşoreanu and Giurgiţeanu, Global Bifurcations in FitzHugh-Nagumo Model, Trends in Mathematics: Bifurcations, Symmetry and Patterns (2003).
-Key words and phrases:
-Keywords: FitzHugh-Nagumo equation, Bifurcations, Global phase portrait, Poincaré compactification, Singular system.
-[…]
+Alexandre A. P. Rodrigues1,† $\dagger$}Corresponding author and Nasrin Sadri2
 1 Lisbon School of Economics and Management, Lisbon University, Centro de Matemática Aplicada à Previsão e Decisão Económica,
@@
 2020 Mathematics Subject Classification:
-Primary: 34C60. Secondary: 34E15, 34E17, 37G15, 37G35
+Primary: 34C60. Secondary: 34E17, 34E15, 37G15, 37G35
 1. Introduction
 Information in nerve fibers is encoded through action potentials (electrical membrane changes).
-Alan Hodgkin and Andrew Huxley, in 1939, managed to make the first intracellular recording of an action potential by inserting microelectrodes into the giant axons of squid.
-In their experiments, they demonstrated that the action potentials are the result of two effects:
-  • a rapid inward current carried by sodium (Na+) ions and
-  • a slow activating outward current carried by potassium (K+) ions.
-Hodgkin and Huxley discovered that the permeability of the membrane for $\text{Na}^{+}$ and $\text{K}^{+}$ was regulated independently, with the conductance depending on both time and the membrane potential.
-They provided a computational model for the action potential in a single cell, known as the Hodgkin-Huxley (HH) model [11].
+A.
+Hodgkin and A.
+Huxley, in 1939, succeeded in recording an action potential from within a neuron for the first time, using microelectrodes inserted into the squid’s giant axon.
+In their experiments, they demonstrated that the action potentials arise from two effects:
+  • a fast inward current mediated by sodium (Na+) ions, and
+  • a slow developing outward current produced by potassium (K+) ions.
+Hodgkin and Huxley showed that the membrane’s permeabilities to $\text{Na}^{+}$ and $\text{K}^{+}$ are controlled separately, with their conductances varying in time and as functions of the membrane voltage.
+They formulated a quantitative description of the action potential in a single neuron, now referred to as the Hodgkin-Huxley (HH) model [11].
 Details and an overview may be found in the review [3].
 1.1. From the Hodgkin-Huxley model to the FitzHugh-Nagumo equations
-Hodgkin and Huxley modelled the measured changes in current by introducing precise probabilistic terms to capture ion channels, which may be either open or closed.
-The HH model elucidates the electrical behaviour of ion channels within the cell membrane, by addressing the passage of $\text{Na}^{+}$ and $\text{K}^{+}$ ions.
-The HH model consists of four ordinary differential equations, each corresponding to one of the state variables [11, Equation 1].
+Hodgkin and Huxley represented the observed current variations using probabilistic variables that describe the closing and opening of ion channels.
+The Hodgkin-Huxley framework characterises the electrical dynamics of these membrane channels by modelling the flow of $\text{Na}^{+}$ and $\text{K}^{+}$ ions across the membrane.
+In the HH framework, the dynamics are governed by four ODEs, each one associated with a specific state variable – [11, Eq.
+1].
 This system is nonlinear, making it analytically difficult to solve explicitly.
-Nonetheless, numerical simulations enable the exploration of certain properties and general behaviours, such as the existence of oscillations and excitability.
-The fact that the HH model correctly captures these excitable dynamics was a major realisation as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (cf.
-[11, Equation 1]).
-While the HH model effectively replicates many neuronal physiological phenomena, it s very complex.
+Nonetheless, numerical simulations allow investigation of specific features and general dynamics, including the presence of excitability and oscillations.
+The ability of the HH model to accurately reproduce these excitable dynamics represented a key insight, offering a complete explanation of the recorded action potential and the mechanisms underlying neural excitability (cf.
+[11]).
+While the HH model successfully reproduces many neuronal physiological phenomena, it is very complex.
 In [7], Richard FitzHugh introduced a simplified model to capture the dynamics of neuronal excitability, which was refined by Jinichi Nagumo two years later [19].
-This model is known as the FitzHugh-Nagumo (FHN) model.
-FitzHugh focused on preserving dynamical features of the HH model, namely the presence of excitability and oscillations.
+This system is known as the FitzHugh-Nagumo (FHN) model.
+FitzHugh concentrated the study on preserving the key dynamical characteristics of the HH model, specifically and oscillatory and excitability behaviour.
 1.2. The FitzHugh-Nagumo equations
-FitzHugh started from the oscillator equation (introduced in 1920) by Balthasar Van der Pol [25], that admits oscillations which are relaxation-like.
-This means there are periods of “low” and “high” states characterised by rapid transitions between them.
-Van der Pol’s equation is built from the simple differential equation for the damped harmonic oscillator.
-He replaced the damping constant by a damping function that depends quadratically on $x,$ thus introducing a nonlinearity ($\dot{x}$ represents the first derivative of the physical position $x$ with respect to $t$ and $c\in\mathbb{R}^{+}_{0}$):
+FitzHugh started with the oscillator equation (introduced in 1920) by Balthasar Van der Pol [25], which supports oscillations resembling relaxation-like dynamics.
+This means there are periods of “low” and “high” states characterised through fast transitions between them.
+The Van der Pol (VdP) equation is derived from the basic differential equation describing a damped harmonic oscillator.
+He replaced the constant damping by a damping function that has a quadratic dependence on $x,$ resulting in the nonlinear differential equation ($\dot{x}$ denotes the first derivative of the physical position $x$ with respect to $t$ and $c\in\mathbb{R}^{+}_{0}$):
 $\ddot{x}+c(x^{2}-1)\dot{x}+x=0.$
-There is only “effective” damping for $|x|<1$ while for $x^{2}>1$ the nonlinear term describes amplification. To interpret the dynamics of the van der Pol equation, one can use the Liénard map [16]
+The system exhibits “effective” damping only when $|x|<1$; for $x^{2}>1$, the nonlinear term leads to amplification. To analyse the dynamics of the VdP equation, the Liénard map can be employed [16]
 $y=\frac{\dot{x}}{c}+\left(\frac{x^{3}}{3}-x\right),$
@@
 $\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}x,\end{array}\right$. (1.4)
-from which one may see the separation in time scales of both equations.
-For $c\gg 1$, while the first one evolves fast (order of $\mathcal{O}(c)$), the second one is much slower of the order of $\mathcal{O}(1/c),$ where $\mathcal{O}$ stands for the usual Landau notation Given a real map $f:\mathbb{R}\to\mathbb{R}$ and $c\in\mathbb{R}$, we say that $f$ is of order $\mathcal{O}(c)$ if there exist $M\in\mathbb{R}^{+}$ and $x_{0}\in\mathbb{R}$ such that $|f(x)|<Mc$, for $x>x_{0}$..
-Building on the van der Pol oscillator, FitzHugh considered the equation:
+from which the time-scale separation of the equations becomes evident.
+For $c\gg 1$, the first variable changes rapidly of the order of $\mathcal{O}(c)$, whereas the second evolves much more slowly, $\mathcal{O}(1/c),$ where $\mathcal{O}$ denotes the standard Landau notation (given a real map $f:\mathbb{R}\to\mathbb{R}$ and $c\in\mathbb{R}$, we say that $f$ is of order $\mathcal{O}(c)$ if there exist $M\in\mathbb{R}^{+}$ and $x_{0}\in\mathbb{R}$ such that $|f(x)|<Mc$, for $x>x_{0}$).
+Based on the VdP oscillator, FitzHugh considered the equation:
 $\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)+z\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x-a+by)\end{array}\right$. (1.5)
@@
   2. (2)
-    the variable $x$ is related to the membrane voltage and the $\text{Na}^{+}$ activation and
+    the variable $x$ corresponds to the membrane voltage and the $\text{Na}^{+}$ activation and
   3. (3)
     the variable $y$ is related to the $\text{Na}^{+}$ inactivation and the $\text{K}^{+}$ activation.
-Equations (1.5) work as an activator-inhibitor model.
-Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with an electrical circuit [19].
+Equations (1.5) act as an activator-inhibitor framework.
+Arimoto, Nagumo, and Yoshizawa proved that it can be realized as an electrical circuit [19].
 Although the FHN model is primarily used to describe neuronal and cardiac systems, it can also be applied in a range of other biological contexts.
-The FHN equations provide a simplified framework for representing interconnected positive and negative feedback loops, enabling the generation of a wide range of responses such as switches, pulses, and stable oscillations [24].
-Among the numerous models used to investigate cardiac cells, the FHN model is notable for being one of the simplest and most extensively researched, effectively representing the general dynamic characteristics of cardiac cells [1, 4].
+The FHN equations provide a simplified framework for representing interconnected negative and positive feedback loops, enabling the generation of a wide range of responses such as switches, pulses, and stable oscillations [24].
+Within the wide range of models used to investigate cardiac cells, the FHN model is notable for being one of the simplest and most extensively researched, effectively representing the general dynamic characteristics of cardiac cells [1, 4].
 Remark 1.1.
 System (1.5) mimics an excitable system and there are limitations regarding the interpretation of the original variables of the HH model.
-Therefore, in the rest of the paper, we will “play safe” and stick to a mathematical motivation in terms of bifurcations and dynamical systems.
+In the rest of the paper, we will take a cautious approach and we stay within a mathematical motivation grounded in bifurcations and dynamical systems.
 1.3. Novelty
-Results in [21, pp.180] synthesised the global bifurcation diagram for the FHN model.
-It has been obtained by putting together as in a “huge puzzle”, all local bifurcation diagrams obtained in Chapters 2–4.
-See also Section 2.4 of [20]
-The main novelty of this paper is the dissection of the global bifurcations diagrams of [21, pp.
-180] and provide a more complete bifurcation analysis of model (1.5), namely the location of Hopf, Pitchfork and Double-zero bifurcations in some cases that will be specified in Subsection 1.4.
-Besides the case-study $c=\mathcal{O}(1)$, we also study the asymptotic case $c\to\pm\infty$ ($\Leftrightarrow\varepsilon=1/c^{2}\to 0$), where canards are observed [12, 14], and we finish the analysis with the compactification of the phase portrait of (1.5) on the Poincaré disc.
-To the best of our knowledge, the application of this procedure to equation (1.5) is new, providing additional information about the trajectories which tend to or come from infinity.
-For some cases, we give the complete description of the phase portraits for (1.5) in the Poincaré disc (i.e.
-in the compactification of $\mathbb{R}^{2}$ adding the circle $\mathbb{S}^{1}$ of the infinity) modulo topological equivalence.
+Results in [21, pp.
+180] synthesised the global bifurcation diagram for the FHN model.
+The result comes from integrating all the local bifurcation diagrams obtained in Chapters 2–4, as if fitting together a “huge puzzle”.
+See Section 2.4 of [20].
+One of the novelties of this paper is the dissection of the bifurcations diagrams of [21, pp.
+180] and provide a more complete bifurcation analysis of model (1.5), namely the location of Hopf, Pitchfork and Double-zero bifurcations, which will be specified in Subsection 1.4.
+Besides the case-study $c=\mathcal{O}(1)$, we also study the asymptotic case
+$c\to\pm\infty\qquad(\Leftrightarrow\quad\varepsilon=1/c^{2}\to 0),$
+where canards are observed [12, 14], and we finish the analysis with the compactification of the phase portrait of (1.5) on the Poincaré disc.
+To the best of our knowledge, this procedure applied to equation (1.5) is new, providing additional information about the trajectories which come from or tend to infinity.
+We classify topologically all phase portraits of system (3.1) “near infinity”, represented on the Poincaré disc (i.e.
+within the compactification of $\mathbb{R}^{2}$ by including the circle at infinity, $\mathbb{S}^{1}$).
+Note that we do not claim to have derived complete phase portraits.
 1.4. Structure
-This paper is structured as follows: Section 2 introduces the definitions and preliminary concepts of blow-up and Poincaré compactification.
+This paper is structured as follows: Section 2 introduces the definitions and preliminary concepts of quasi-homogeneous blow-up and Poincaré compactification.
 In Section 3, we discuss the finite equilibria of (3.1) for the following three cases:
@@
 Case C:
-[…]
+$a\neq 0,\,0<b<1$
 and investigate possible bifurcations, including Pitchfork, Double Homoclinic and Hopf bifurcations.
@@
 In Section 4, we analyse the asymptotic dynamics of (1.5) when $c\to\pm\infty$ and we relate the existence of canards with the periodic solution emerging from the Hopf bifurcation.
-Section 5 is dedicated to the study of the global phase portrait of (1.5) in the Poincaré disc.
-The phase portraits for each connected component in Cases A, B and C are completely illustrated.
-This completes the work started in [8].
+Section 5 is dedicated to studying the dynamics “near infinity” for (1.5) on the Poincaré disc.
+The phase portraits for each connected component in Cases A, B and C are illustrated.
+This complements the work started in [8].
 Section 6 finishes the paper.
-Throughout this paper, we have endeavoured to make a self-contained exposition bringing together all topics related to the proofs.
-We have drawn illustrative figures to make the paper easily readable and all results are illustrated with numerical simulations using the software Matlab R2015b.
+In this paper, we aim to provide a self-contained exposition, integrating all topics pertinent to the proofs.
+Illustrative figures have been included for clarity, and all results are supported by numerical simulations performed using Matlab R2015b and pplane9.
 2. Preparatory section
-In this section, we introduce some terminology for polynomial vector fields on $\mathbb{R}^{2},$ that will be used in the remaining sections.
-Let $f_{(a,b,c)}$ be a smooth vector field on $\mathbb{R}^{2}$ with flow given by the unique solution $x(t)=\varphi(t,x)\in\mathbb{R}^{2}$ of the three-parameter family
+This section provides definitions for polynomial vector fields on $\mathbb{R}^{2}$ that will be used in the sections that follow.
+Let $f_{(a,b,c)}$ be a smooth vector field on $\mathbb{R}^{2}$ whose flow is given by the unique solution $X(t)=\varphi(t,X)\in\mathbb{R}^{2}$ of the three-parameter family
 $\dot{X}=f_{(a,b,c)}(X),\qquad X(0)=X_{0}\in\mathbb{R}^{2},$ (2.1)
-where $\0(a,b,c)\in\mathbb{R}\times\mathbb{R}\times\mathbb{R}\backslash\{0\}$.
+[…]
 2.1. Useful terminology
 The center manifold of a non-hyperbolic equilibrium $p\in\mathbb{R}^{2}$ of (2.1) is the set of solutions whose behaviour around $p$ is controlled neither by the exponential attraction of the stable manifold nor by the exponential repulsion of the unstable manifold.
-If $Df_{(a,b,c)}$ evaluated at a given equilibrium $p$, $Df_{(a,b,c)}(p)$, has an eigenvalue with zero real part, the center manifold plays an important role and this is the set where bifurcations might occur.
-For a fixed triple $\0(a,b,c)\in\mathbb{R}\times\mathbb{R}\times\mathbb{R}\backslash\{0\}$, an equilibrium point $p$ of (2.1) is called nilpotent singularity when all eigenvalues of Jacobian matrix $Df_{(a,b,c)}(p)$ are zero but $Df_{(a,b,c)}(p)\neq 0$, and is called linearly zero when $Df_{(a,b,c)}(p)\equiv 0.$ Further, the singular point $p$ is called semi-hyperbolic if exactly one eigenvalue of $Df_{(a,b,c)}(p)$ is equal to $0$
+If $Df_{(a,b,c)}$ evaluated at an equilibrium $p$, $Df_{(a,b,c)}(p)$, has an eigenvalue with zero real part, the center manifold plays an important role and this is the set where bifurcations might occur.
+[…]
 Throughout this paper, we study the Double-zero (DZ) singularity with symmetry $\mathbb{Z}_{2}$ for a family of differential equations corresponding to the case described in [27, pp.
 400].
-The unfolding of this singularity of codimension two involves lines of Belyakov transitions, Pitchfork, Hopf, Double Homoclinic and saddle-node of limit cycles.
-For the notion of super/subcritical Hopf bifurcation, we use the terminology of Golubitsky and Schaeffer [9, Chapter IV, Section 2].
-We suggest the reading of [15, 18, 27] for a complete understanding of these bifurcations, as well as the sufficient conditions that prompt their existence.
+The unfolding of this singularity of codimension two involves lines of Belyakov transitions, Pitchfork, Hopf, Double Homoclinic and saddle-node bifurcation of limit cycles.
+A Hopf bifurcation is termed supercritical when the first Lyapunov coefficient of its normal form is negative, leading to the emergence of a stable small-amplitude limit cycle, and subcritical when this coefficient is positive, in which an unstable small-amplitude limit cycle is created [26].
+For full details on these bifurcations and the criteria that guarantee their appearance, see [15, 18, 27].
 Remark 2.1.
-In this paper, we refer to an equilibrium point as undergoing a Belyakov transition if at least one pair of eigenvalues of the Jacobian matrix (evaluated at the equilibrium) changes from real to complex conjugate or vice versa, while the sign of their real part remains unchanged.
+In this paper, we refer to an equilibrium point $p$ as undergoing a Belyakov transition if at least one pair of eigenvalues of $Df_{(a,b,c)}(p)$ changes from real to complex conjugate or vice versa, while the sign of their real part remains unchanged.
 Although such a transition is typically considered in higher-dimensional systems, we adopt this terminology here for the sake of brevity and clarity.
-2.2. Blow-up Method
-The blow-up method is one of the important methods to study the topological behaviour of solutions of a dynamical system in the neighbourhood of nilpotent and linearly zero singularities.
-In this method, an equilibrium point is “expanded” into a line or a circle, and new equilibria on the line or circle are examined.
-There are various methods for doing a blow-up.
-In this paper, we concentrate our attention on blow-ups in a given direction, a method which will be used in the Section 5 of the present article.
-In what follows, following [2], we review some concepts related to the topic.
-Consider a planar polynomial differential equation associated with the vector field $\mathbf{X}=(P,Q),$ in $\mathbb{R}^{2}$ of the form:
-$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{2})=P_{m}(x_{1},x_{2})+\cdots\\ \\ \dot{x}_{2}=Q(x_{1},x_{2})=Q_{m}(x_{1},x_{2})+\cdots\end{array}\right.$ (2.2)
-where
-  • $P,Q$ are coprime polynomials;
-  • $P_{m},Q_{m}$ are homogeneous polynomials of degree $m\in\mathbb{N}$ and
-  • the dots $\cdots$ stand for high-order terms in the variables $x_{1},x_{2}.$
-Since $m>0,$ the origin is a singularity. By using the change of coordinates
-$(x_{1},x_{2})\mapsto(r\cos(\theta),r\sin(\theta)),$
-system (2.2) may be written in polar coordinates (if $r\neq 0$) as:
-$\left\{\begin{array}[]{l}\dot{r}=\mathcal{R}(\theta)r+\cdots\\ \\ \dot{\theta}=\mathcal{T}(\theta)+\cdots\end{array}\right.$ (2.3)
-where:
-  • $\mathcal{R}$ and $\mathcal{T}$ are polynomials in $\cos{(\theta)}$ and $\sin{(\theta)}$ and
-  • the dots $\cdots$ stand for high order terms in $r.$
-If $\mathcal{T}\neq 0,$ all the solution curves tending (as $t\to\pm\infty$) to the origin are tangent to the solutions $\theta^{\star}\in[0,2\pi)$ of the equation $\mathcal{T}(\theta)=0.$ The map $\mathcal{T}$ is often called the characteristic polynomial and $\theta^{\star}$ is called characteristic direction associated with (2.2).
-The polynomial $\mathcal{T}$ in the cartesian coordinates $(x_{1},x_{2})$ may be written as:
-$\mathcal{T}(x_{1},x_{2})=x_{1}Q_{m}(x_{1},x_{2})-x_{2}P_{m}(x_{1},x_{2}).$
+2.2. Quasi-homogeneous blow-up method
+Quasi-homogeneous blow-ups are used to analyze and simplify the behaviour of differential equations near non-hyperbolic singularities by rescaling the variables according to their natural weights.
+This technique transforms the singularity into a more regular, analyzable object.
+Although homogeneous blow-ups
+$\phi:\mathbb{S}^{1}\times\mathbb{R}\to\mathbb{R}^{2},\quad(\theta,\rho)\mapsto\left(\rho\cos\theta,\,\rho\sin\theta\right),$
+suffice for the study of isolated singularities of an analytic vector field, the quasi-homogeneous approach (scaling the variables with different degrees) is considerably more efficient.
+Here we review this method briefly – for more details and examples, see [5, Section 3.3].
+Consider the polynomial vector field $f=(P,Q)$ associated with the planar differential equation:
+$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{2})\\ \\ \dot{x}_{2}=Q(x_{1},x_{2}).\end{array}\right$. (2.2)
+In the language of singularity theory, we have:
+$f(x_{1},x_{2})=P(x_{1},x_{2})\frac{\partial}{\partial x_{1}}+Q(x_{1},x_{2})\frac{\partial}{\partial x_{2}},$
+with an isolated singular point at the origin $(0,0)$. Since $P$ and $Q$ are polynomial we may find $p_{ij},q_{ij}\in\mathbb{R}$ such that:
+$P(x_{1},x_{2})=\sum_{i+j\geq 1}p_{ij}x_{1}^{i}x_{2}^{j}\qquad\text{and}\qquad Q(x_{1},x_{2})=\sum_{i+j\geq 1}q_{ij}x_{1}^{i}x_{2}^{j}.$
 Definition 2.2.
-Blow-up in $x_{1}$ direction is a “non-bijective” change of coordinates defined as
-$(x_{1},z)\mapsto(x_{1},zx_{1})=(x_{1},x_{2}),$
-where $z\geq 0$ is a new variable.
-This change of coordinates projects the origin of (2.2) on the line $x_{1}=0.$ The expression of system (2.2) after the blow-up in the $x_{1}$ direction may be written as:
-$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{1}z),\\ \\ \dot{z}=\dfrac{Q(x_{1},x_{1}z)-zP(x_{1},x_{1}z)}{x_{1}}.\end{array}\right.$ (2.4)
-After the blow-up, we may cancel the common factor $x_{1}^{m-1}.$ The $x_{1}$ directional blow-up is equivalent on $x_{1}\neq 0$ and $\theta\neq\pi/2,3\pi/2,$ since there exists an analytic change of coordinates bringing $(r,\theta)$ and $(x_{1},x_{2}).$ To investigate the behavior of the solutions of system (2.2) around equilibria, it is necessary to examine the behaviour of the equilibrium points of (2.4) on the line $x_{1}=0.$ If some of these points are singular, then we may continue the blow-up process.
+The quasi-homogeneous blow-up is defined by
+$\phi:\mathbb{S}^{1}\times\mathbb{R}\to\mathbb{R}^{2},\quad(\theta,\rho)\mapsto\left(\rho^{\alpha}\cos\theta,\,\rho^{\beta}\sin\theta\right)=(x_{1},x_{2}),$
+for some $\alpha,\beta\in\mathbb{N}.$
+If $(\alpha,\beta)=(1,1)$, this corresponds to the standard homogeneous case.
+Doing the quasi-homogeneous blow-up (of the previous definition) on $f=(P,Q)$, it yields the following differential equation:
+$\dot{\rho}=\frac{P\rho^{\beta}\cos\theta+Q\rho^{\alpha}\sin\theta}{\rho^{\alpha+\beta-1}(\alpha\cos^{2}\theta+\beta\sin^{2}\theta)},\qquad\dot{\theta}=\frac{\alpha Q\rho^{\alpha}\cos\theta-\beta P\rho^{\beta}\sin\theta}{\rho^{\alpha+\beta}(\alpha\cos^{2}\theta+\beta\sin^{2}\theta)},$
+in which $P$ and $Q$ are computed at $(\rho^{\alpha}\cos\theta,\rho^{\beta}\sin\theta)$. A common factor $\rho^{d}>0$ can be cancelled for a suitable choice of $d>0$.
+Throughout this article, we will be interested in quasi-homogeneous directional blow-ups in a given direction.
+This technique isolates and desingularizes the dynamics of trajectories approaching the singularity along that direction, creating a chart where the vector field becomes simpler and its behaviour can be analyzed separately from other directions.
+Definition 2.3.
+The quasi-homogeneous directional blow-up along the positive (resp. negative) $x$ axis is given by the map:
+$\overline{f}^{x}_{+}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}^{\alpha},\bar{x}^{\beta}\bar{y}),$
+$\overline{f}^{x}_{-}:\quad(\bar{x},\bar{y})\mapsto(-\bar{x}^{\alpha},\bar{x}^{\beta}\bar{y}),$
+and the quasi-homogeneous directional blow-up along the positive (resp. negative) $y$ axis is described by:
+$\overline{f}^{y}_{+}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}\bar{y}^{\alpha},\bar{y}^{\beta}),$
+$\overline{f}^{y}_{-}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}\bar{y}^{\alpha},-\bar{y}^{\beta}),$
+where $\bar{x}$ and $\bar{y}$ denote new variables and $\alpha,\beta\in\mathbb{N}.$
+Among all possible pairs $(\alpha,\beta)\in\mathbb{N}\times\mathbb{N}$ of the definition of quasi-homogeneous blow-ups, a natural question is which choice is the most “efficient” (in the sense that we desingularize the equilibria in the minimum number of steps) for performing a quasi-homogeneous blow-up to simplify and resolve a singularity.
+Selecting such an efficient pair balances the relative importance of the variables, making the leading-order dynamics near the singularity clear.
+The Newton polygon provides a systematic tool to guide this choice, which we now introduce (adapted from [5]).
+Definition 2.4.
+We define the support of $f$ as the subset
+$S=\{(m-1,n)\mid p_{mn}\neq 0\}\cup\{(m,n-1)\mid q_{mn}\neq 0\}\subset\mathbb{R}^{2},$
+where the point (1,n)(-1,n) corresponds to the monomial $p_{0n}x_{2}^{n}$ and the point (m,1)(m,-1) corresponds to the monomial $q_{m0}x_{1}^{m}.$ Furthermore, $(0,0)$ is associated with the monomials $p_{10}x_{1}$ and $q_{01}x_{2}.$
+Definition 2.5.
+The Newton polygon associated to $f$ is given by the convex hull
+$\bigcup_{(r,s)\in S}\{({r}^{*},{s}^{*})\in\mathbb{R}^{2}\mid r\leq{r}^{*},\,s\leq{s}^{*}\}.$
+The Newton polygon associated to $f$ consists of the union $\gamma$ of the points/segments $\gamma_{k}$, enumerated in order from left to right.
+For each $1\leq k\leq n$, the segment $\gamma_{k}$ lies on the straight line $\alpha_{k}r+\beta_{k}s=d_{k}$ where $\alpha_{k}$ and $\beta_{k}$ are coprime integers and $d_{k}$ is a constant.
+Since $f$ has an isolated singularity at the origin, then at least one of the two points $(0,s)$ or (1,s)(-1,s) belongs to the support $S$ for some $s$ and, similarly, at least one of the two points (r,1)(r,-1) or $(r,0)$ lies in $S$, for some $r$.
+Thus a first segment $\gamma_{1}$ always exists in the Newton diagram ($\Leftrightarrow$ it is non-empty).
+The Newton diagram identifies the dominant monomials near a singularity; the slopes of its compact edges determine the weight ratio $\beta/\alpha$, $\alpha\neq 0$, used to choose the exponents of a quasi-homogeneous blow-up.
+Proposition 2.6.
+[5, Section 3.3, adapted].
+Under the previous conditions, the most efficient pair $(\alpha,\beta)$ for the quasi-homogeneous blow-up corresponds to the edge of the Newton polygon with minimal slope (in absolute value).
 2.3. Poincaré compactification
-In this section, we present a concise overview of the Poincaré compactification following Chapter 5 of [5], a technique used in Section 5 of the present paper.
-Consider the polynomial vector field (2.2) in $\mathbb{R}^{2}$ where $d_{1}$ and $d_{2}$ are the algebraic degrees of $P$ and $Q$ respectively and $d={\max}\{d_{1},d_{2}\}.$
+Here, following Chapter 5 of [5], we present a concise overview of the Poincaré compactification to study how trajectories of a planar differential system behave “near infinity”.
+This technique will be used in Section 5 of the present paper.
+To construct the phase portrait, one would ideally need to consider the complete real plane $\mathbb{R}^{2},$ which is often impractical.
+However, if the vector field is defined by polynomial functions, Poincaré compactification can be applied.
+This method allows the phase portrait to be represented within a finite region and also provides control over orbits that approach or come from infinity.
+Consider the polynomial vector field $f=(P,Q)$, where $d_{1}$ and $d_{2}$ are the algebraic degrees of $P$ and $Q$ respectively and $d={\max}\{d_{1},d_{2}\}.$
 2.3.1. Preliminaries for the construction
-First identify $\mathbb{R}^{2}$ with the plane $\Pi$ in $\mathbb{R}^{3}$ defined by $(y_{1},y_{2},y_{3})=(x_{1},x_{2},1).$ The sphere $\mathbb{S}^{2}=\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}:y_{1}^{2}+y_{2}^{2}+y_{3}^{2}=1\}$ is called the Poincaré sphere and is tangent to the plane $\Pi$ at $(0,0,1).$ The sphere may be written as
+First identify $\mathbb{R}^{2}$ with the plane $\Pi$ in $\mathbb{R}^{3}$ defined as $(y_{1},y_{2},y_{3})=(x_{1},x_{2},1).$ The sphere
+$\mathbb{S}^{2}=\left\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}:\sum_{i=1}^{3}{y_{i}}^{2}=1\right\}$
+is called the Poincaré sphere and is tangent to $\Pi$ at $(0,0,1).$ The sphere may be written as
 $H^{+}\cup H^{-}\cup\mathbb{S}^{1}$
@@
 Definition 2.7.
-The projection of $\mathbf{X}$ from $\mathbb{R}^{2}$ to $\mathbb{S}^{2}$ is given by the central projections
-$f^{+}:\mathbb{R}^{2}\to\mathbb{S}^{2}\quad\text{and}\quad f^{-}:\mathbb{R}^{2}\to\mathbb{S}^{2}$
-where $f^{+}(x_{1},x_{2})$ is the intersection point of the line passing through the origin and $(x_{1},x_{2})\in\mathbb{R}^{2}$, with $H^{+}$. More precisely:
+The projection of $\mathbf{X}:\mathbb{R}^{2}\to\mathbb{S}^{2}$ is defined with the central projections
+$f^{\pm}:\mathbb{R}^{2}\to\mathbb{S}^{2}$
+where $f^{\pm}(x_{1},x_{2})$ is the intersection point of the line that passes through $(0,0)$ and $(x_{1},x_{2})\in\mathbb{R}^{2}$, with $H^{\pm}$. More precisely:
 $f^{+}(x_{1},x_{2})=\left(\frac{x_{1}}{\Delta(x)},\frac{x_{2}}{\Delta(x)},\frac{1}{\Delta(x)}\right).$
@@
 $f^{-}(x_{1},x_{2})=\left(-\frac{x_{1}}{\Delta(x)},-\frac{x_{2}}{\Delta(x)},-\frac{1}{\Delta(x)}\right).$
-We obtain vector fields in each hemisphere (analytically conjugate to the initial vector field $\mathbf{X}$.
-For $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the induced vector field on $H^{+}$ and $H^{-}$ is defined, respectively, by:
-$\overline{\mathbf{X}}(y)$ $=Df^{+}(x)\mathbf{X}(x),\quad\text{where}\quad y=f^{+}(x),$
-$\overline{\mathbf{X}}(y)$ $=Df^{-}(x)\mathbf{X}(x),\quad\text{where}\quad y=f^{-}(x).$
+For $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the induced vector field on $H^{\pm}$ is defined, respectively, by:
+$\overline{\mathbf{X}}(y)$ $=Df^{\pm}(x)\mathbf{X}(x),\quad\text{for}\quad y=f^{\pm}(x).$
 Remark 2.8.
-The vector field $\overline{\mathbf{X}}$ on $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ is everywhere tangent to $\mathbb{S}^{2}.$
-The points at infinity of $\mathbb{R}^{2}$ (each direction is associated with two points) are in bijective correspondence with the points of the equator of $\mathbb{S}^{2}.$ In the next subsection, we are going to extend the induced vector field $\overline{\mathbf{X}}$ from $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ to $\mathbb{S}^{2}$.
+The vector field $\overline{\mathbf{X}}$ is tangent to the sphere $\mathbb{S}^{2}$ at every point $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$.
+The infinite points of $\mathbb{R}^{2}$ (each direction is associated with two points) are in bijective associated to the points of the equator of $\mathbb{S}^{2}.$ In the next subsection, we are going to generalize the induced vector field $\overline{\mathbf{X}}:\mathbb{S}^{2}\backslash\mathbb{S}^{1}\rightarrow\mathbb{S}^{2}.$
 Definition 2.9.
-The extended vector field on $\mathbb{S}^{2}$ is called the Poincaré compactification of the vector field $\mathbf{X}$ on $\mathbb{R}^{2},$ and it is denoted by $p(\mathbf{X}).$
+The extension of $\mathbf{X}$ to $\mathbb{S}^{2}$ is referred to as its Poincaré compactification, denoted by $p(\mathbf{X}).$
 2.3.2. About the construction
-We use smooth charts to make calculations. For $y=(y_{1},y_{2},y_{3})\in\mathbb{S}^{2},$ we use the six local charts given by
-$U_{k}=\{y\in\mathbb{S}^{2}:y_{k}>0\},\quad V_{k}=\{y\in\mathbb{S}^{2}:y_{k}<0\},\quad k=1,2,3.$
-The corresponding local maps $\Phi_{k}:U_{k}\to\mathbb{R}^{2}$ and $\Psi_{k}:V_{k}\to\mathbb{R}^{2}$ are defined as
-$\Phi_{k}(y)=-\Psi_{k}(y)=\left(\frac{y_{m}}{y_{k}},\frac{y_{n}}{y_{k}}\right),\quad\text{for}\quad m<n\quad\text{and}\quad m,n\neq k.$
-We denote by $z=(u,v)$ the value of $\Phi_{k}(y)$ or $\Psi_{k}(y)$ for any $k\in\{1,2,3\}$, such that $(u,v)$ will play different roles depending on the local chart we are considering.
-Observe that points lying in $\mathbb{S}^{1}$, in any chart, have $v=0$.
-We perform a detailed calculation of the expression of $p(\mathbf{X})$ in the local chart $U_{1}$.
-From (2.2), we have $\mathbf{X}(x)=(P(x_{1},x_{2}),Q(x_{1},x_{2}))$.
-Then $\overline{\mathbf{X}}(y)=Df^{+}(x)\mathbf{X}(x)$ with $y=f^{+}(x)$ and, using the Chain rule, one gets:
-$D\Phi_{1}(y)\overline{\mathbf{X}}(y)=D\Phi_{1}(y)\circ Df^{+}(x)X(x)=D(\Phi_{1}\circ f^{+})(x)\mathbf{X}(x).$
-Let $\overline{\mathbf{X}}|_{U_{1}}$ denote the system defined as $D\Phi_{1}(y)\overline{\mathbf{X}}(y)$. Then since
-$(\Phi_{1}\circ f^{+})(x)=\left(\frac{x_{2}}{x_{1}},\frac{1}{x_{1}}\right)=(u,v),$
-we have
-$\overline{\mathbf{X}}|_{U_{1}}=v^{2}\left(-\frac{u}{v}P\left(\frac{1}{v},\frac{u}{v}\right)+\frac{1}{v}Q\left(\frac{1}{v},\frac{u}{v}\right),-P\left(\frac{1}{v},\frac{u}{v}\right)\right).$
-On the other hand, one knows that $\rho(y)=y_{3}^{d-1}=v^{d-1}m(z)$ where
-$m(z)=(1+u^{2}+v^{2})^{\frac{1-d}{2}}.$
-Then we get:
-$\rho\left(\overline{\mathbf{X}}|_{U_{1}}\right)(z)=v^{d+1}m(z)\left(-\frac{u}{v}P\left(\frac{1}{v},\frac{u}{v}\right)+\frac{1}{v}Q\left(\frac{1}{v},\frac{u}{v}\right),-P\left(\frac{1}{v},\frac{u}{v}\right)\right).$
-Remark 2.6.
-In order to prove that the extension of $\rho\overline{\mathbf{X}}$ to $p(\mathbf{X})$ is defined on the whole of $\mathbb{S}^{2}$ we notice that while $\overline{\mathbf{X}}|_{U_{1}}$ is not well defined when $v=0$, and $p(\mathbf{X})|_{U_{1}}=\rho\overline{\mathbf{X}}|_{U_{1}}$ is well defined along $v=0$, since the multiplying factor $v^{d+1}$ cancels any factor of $v$ which may appear in the denominator.
-The same line of arguments may be applied to the other local charts.
-To simplify the extended vector field we also make a change in the time variable and remove the factor $m(z)$.
-We still keep a vector field on $\mathbb{S}^{2}$ which is $C^{\omega}$–equivalent to $\mathbf{X}$ on any of the hemispheres $H^{+}$ and $H^{-}$.
+We use smooth charts to make calculations. For $y=(y_{1},y_{2},y_{3})\in\mathbb{S}^{2},$ we utilize the six local charts:
+$U_{l}=\{y\in\mathbb{S}^{2}:0<y_{l}\},\quad V_{l}=\{y\in\mathbb{S}^{2}:0>y_{l}\},\quad l=1,2,3.$
+The associated local maps $\Phi_{l}:U_{l}\to\mathbb{R}^{2}$ and $\Psi_{l}:V_{l}\to\mathbb{R}^{2}$ are described as
+$\Phi_{l}(y)=-\Psi_{l}(y)=\left(\frac{y_{m}}{y_{l}},\frac{y_{n}}{y_{l}}\right),\quad\text{for}\quad m<n\quad\text{where}\quad m,n\neq l.$
+Consider $z=(u,v)$ as the value of $\Phi_{l}(y)$ or $\Psi_{l}(y)$ for any $l\in\{1,2,3\}$, such that $(u,v)$ plays different roles depending on the local chart we are considering.
+In any chart, points on $\mathbb{S}^{1}$ correspond to those with $v=0$.
 2.3.3. Explicit expressions
-The expression for $p(\mathbf{X})$ in local chart $(U_{1},\Phi_{1})$ is given by:
-$\dot{u}=v^{d}\left[-uP\left(\frac{1}{v},\frac{u}{v}\right)+Q\left(\frac{1}{v},\frac{u}{v}\right)\right],\hskip 14.22636pt\dot{v}=-v^{d+1}P\left(\frac{1}{v},\frac{u}{v}\right),$ (2.5)
+In the local chart $(U_{1},\Phi_{1}),$ $p(\mathbf{X})$ is expressed as
+$\dot{u}=\left[Q\left(\frac{1}{v},\frac{u}{v}\right)-uP\left(\frac{1}{v},\frac{u}{v}\right)\right]v^{d},\hskip 14.22636pt\dot{v}=-P\left(\frac{1}{v},\frac{u}{v}\right)v^{d+1},$ (2.3)
 the expression for $(U_{2},\Phi_{2})$ is:
-$\dot{u}=v^{d}\left[P\left(\frac{u}{v},\frac{1}{v}\right)-uQ\left(\frac{u}{v},\frac{1}{v}\right)\right],\hskip 14.22636pt\dot{v}=-v^{d+1}Q\left(\frac{u}{v},\frac{1}{v}\right),$ (2.6)
+$\dot{u}=\left[P\left(\frac{u}{v},\frac{1}{v}\right)-uQ\left(\frac{u}{v},\frac{1}{v}\right)\right]v^{d},\hskip 14.22636pt\dot{v}=-Q\left(\frac{u}{v},\frac{1}{v}\right)v^{d+1},$ (2.4)
 and the expression for $(U_{3},\Phi_{3})$ is:
@@
 Remark 2.10.
-For $i=1,2,3$, the expressions of the vector field $p(\mathbf{X})$ in the local chart $(V_{i},\Psi_{i})$ are identical to those in the local chart $(U_{i},\Phi_{i}),$ up to a multiplication by $({-1})^{d-1}$.
-To study $\mathbf{X}$ in the complete plane $\mathbb{R}^{2}$, including its behaviour near infinity, it suffices to work on $H^{+}\cup\mathbb{S}^{1}$, which we call the Poincaré disc.
-All calculations can be performed in the three charts $(U_{1},\Phi_{1})$, $(U_{2},\Phi_{2})$, and $(U_{3},\Phi_{3})$ in which case the expressions are given by the formulas (2.5), (2.6) and (2.7).
-To obtain (2.5) we start with (2.2) and introduce coordinates $(u,v)$ through the equality:
-$(x_{1},x_{2})=\left(\frac{1}{v},\frac{u}{v}\right).$
-In each local chart, the local representative of $p(\mathbf{X})$ is a polynomial vector field.
+For $i=1,2,3$, the representations for the vector field $p(\mathbf{X})$ in the local chart $(V_{i},\Psi_{i})$ are identical to those in $(U_{i},\Phi_{i}),$ up to a multiplication by $({-1})^{d-1}$.
+To investigate $\mathbf{X}$ throughout the complete plane $\mathbb{R}^{2}$, as well as its behaviour “near infinity”, it is enough to examine it on $H^{+}\cup\mathbb{S}^{1}$, a region known as the Poincaré disc.
+All calculations can be performed in the three charts $(U_{i},\Phi_{i})$ for $i=1,2,3$ whose expressions are given by the formulas (2.3), (2.4) and (2.5).
 Definition 2.11.
-We call finite (resp. infinite) singular points of $\mathbf{X}$ the singular points of $p(\mathbf{X})$ which lie in $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ (resp. $\mathbb{S}^{1}$).
+We refer to singular points of $\mathbf{X}$ as finite when the corresponding points of $p(\mathbf{X})$ are in $\mathbb{S}^{2}\setminus\mathbb{S}^{1},$ and as infinite when they belong to $\mathbb{S}^{1}$.
 2.3.4. Useful consequences
@@
   1. (1)
-    If $y\in\mathbb{S}^{1}$ is an infinite singular point, then $-y$ is also a singular point.
+    Whenever $y\in\mathbb{S}^{1}$ is an infinite singular point, so is y-y.
   2. (2)
-    Since the local behavior near $-y$ is the local behavior near $y$ multiplied by $(-1)^{d-1}$, then:
-    1. (a)
-      the orientation of the orbits changes when the degree $d$ is even.
-    2. (b)
-      if $d$ is even and $y\in\mathbb{S}^{1}$ is a stable node of $p(\mathbf{X})$, then $-y$ is an unstable node.
+    The local behaviour near y-y is that near $y$ scaled by $(-1)^{d-1}$.
   3. (3)
-    Infinite singular points appear in pairs of diametrically opposite points so that it is enough to study half of them. Using the degree of the vector field one can determine the other half.
+    Because infinite singular points arise in pairs, it is sufficient to examine just half of them. The rest can be inferred via the vector field’s degree.
   4. (4)
-    The integral curves of $\mathbb{S}^{2}$ are symmetric with respect to the origin, such that it is sufficient to represent the flow of $p(\mathbf{X})$ only in the closed northern hemisphere, the so called Poincaré disc.
+    The integral curves on $\mathbb{S}^{2}$ are symmetric aroud the origin, so it suffices to depict the flow of $p(\mathbf{X})$ in the closed northern hemisphere, the so called Poincaré disc.
 2.3.5. Infinite singular points
 The theory described here will be useful in Section 5 of the present paper.
-We aim to study the local phase portrait at infinite singular points.
-For this we choose an infinite singular point $(u,0)$ and start by looking at the expression of the linear part of the vector field $p(\mathbf{X})$.
-Denote by $P_{i}$ and $Q_{i}$ the homogeneous polynomials of degree $i\in\mathbb{N}_{0}$ for $i=0,1,...,d$ such that
+Our goal is to analyse the local phase portrait near infinite singular points.
+To this end, we select a singular point $(u,0)$ and begin with examining the linear part of the vector field $p(\mathbf{X})$.
+Denote by $P_{i}$ and $Q_{i}$ the homogeneous polynomials of degree $i\in\mathbb{N}_{0}$ for i=0,1,,di=0,1,...,d such that
 $\left\{\begin{array}[]{l}P=P_{0}+P_{1}+...+P_{d},\\ \\ Q=Q_{0}+Q_{1}+...+Q_{d}.\end{array}\right$.
-Then $(u,0)\in\mathbb{S}^{1}\cap(U_{1}\cup V_{1})$ is an infinite singular point of $p(\mathbf{X})$ if and only if
-$F(u)\equiv Q_{d}(1,u)-uP_{d}(1,u)=0.$
-Similarly $(u,0)\in\mathbb{S}^{1}\cap(U_{2}\cup V_{2})$ is an infinite singular point of $p(\mathbf{X})$ if and only if
-$G(u)\equiv P_{d}(u,1)-uQ_{d}(u,1)=0.$
+Then $(u,0)\in(U_{1}\cup V_{1})\cap\mathbb{S}^{1}$ is an infinite singular point of $p(\mathbf{X})$ if and only if
+$F(u)\equiv-uP_{d}(1,u)+Q_{d}(1,u)=0.$
+In the same way $(u,0)\in(U_{2}\cup V_{2})\cap\mathbb{S}^{1}$ is an infinite singular point for $p(\mathbf{X})$ if and only if
+$G(u)\equiv-uQ_{d}(u,1)+P_{d}(u,1)=0.$
 The Jacobian of the vector field $p(\mathbf{X})$ at the point $(u,0)$ is
 $\left(\begin{array}[]{cc}F^{\prime}(u)&Q_{d-1}(1,u)-uP_{d-1}(1,u)\\ \\ 0&-P_{d}(1,u)\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}G^{\prime}(u)&P_{d-1}(u,1)-uQ_{d-1}(u,1)\\ \\ 0&-Q_{d}(u,1)\end{array}\right)$
-if $(u,0)$ belongs to $U_{1}\cup V_{1}$ or $U_{2}\cup V_{2}$, respectively.
+if $(u,0)$ lies in $U_{1}\cup V_{1}$ or $U_{2}\cup V_{2}$, respectively.
 Our discussion is concentrated on isolated singularities in the equator.
-Following [5], among the hyperbolic singular points at infinity only nodes and saddles may appear.
-All the semi-hyperbolic singular points can appear at infinity.
-If one of these hyperbolic or semi-hyperbolic singularities at infinity is a (topological) saddle, then the straight line defined by $v=0$, representing the equator of $\mathbb{S}^{2}$, is a stable or unstable manifold, or a center manifold.
-The same property also holds for semi-hyperbolic singularities of saddle-node type.
-They can have their hyperbolic sectors split in two different ways depending on the Jacobian matrix of the system in the charts $U_{1}$ or $U_{2}$.
-These matrices may be either
-$\left(\begin{array}[]{cc}a&\star\\ \\ 0&0\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}0&\star\\ \\ 0&a\end{array}\right)$
-with $a\neq 0$ and $\star\in\mathbb{R}$.
-The sense of the orbits can also be the opposite.
-The nilpotent points, as well as the singularities with zero linear part, have a behavior at infinity that is quite a bit more complicated than the hyperbolic and elementary singular points.
-Blow-up is needed to study them.
+According to [5], only nodes and saddles can arise as hyperbolic singularities at infinity.
+All semi-hyperbolic singular points may also emerge at infinity.
+In the case that one of these singular points at infinity is a (topological) saddle, the line $v=0,$ corresponding to the equator of $\mathbb{S}^{2}$, serves as a stable, unstable, or center manifold.
+This property also applies to semi-hyperbolic points of saddle-node type, whose hyperbolic sectors may be divided into two different forms depending on the Jacobian matrix in charts $U_{1}$ or $U_{2}.$ These matrices may be either
+$\left(\begin{array}[]{cc}0&\star\\ \\ 0&a\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}a&\star\\ \\ 0&0\end{array}\right)$
+where $a\neq 0$ and $\star\in\mathbb{R}$.
+Nilpotent equilibria exhibit behaviour at infinity that is considerably more complex than those of hyperbolic equilibria.
+Blow-up techniques of Subsection 2.2 are needed to study them.
 3. Finite equilibria and bifurcation analysis
 Our object of study is the analysis of the family of differential equations (1.5), which can be written as:
-$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x-a+by)\end{array}\right.$ (3.1)
-where $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}.$ We can absorb the parameter $z$ of (1.5) into the variable through the change of variable $y\to y+z$ and redefinition of $a\to a-bz.$ The model (3.1) is invariant under the transformation $(t,c)\to(-t,-c)$ so we can restrict to the case $c>0$, which means that the stability of the equilibria for $c>0$ is the reverse of that of $c<0$.
-Let us denote by $f_{(a,b,c)}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ the vector field associated with (3.1).
-For $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}$, the Jacobian matrix of $f_{(a,b,c)}$ at a general point $(x,y)\in\mathbb{R}^{2}$ is given by:
+$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x+by-a)\end{array}\right$. (3.1)
+[…]
+[…]
 $Df_{(a,b,c)}(x,y)=\left(\begin{array}[]{cc}c-cx^{2}&c\\ \\ -1/c&-b/c\end{array}\right).$ (3.5)
@@
 The vector field $f_{(0,b,c)}$ is $\mathbb{Z}_{2}(\kappa)$–equivariant.
-For all $a=0$, $b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\},$ we know that $E_{1}=(0,0)$ is an equilibrium of (3.1).
-If $b\in(-\infty,0)\cup[1,+\infty),$ then system (3.1) has two extra equilibria given explicitly by
+[…]
 $E_{2}=\kappa(E_{3})=\left(-\sqrt{\dfrac{3(b-1)}{b}},\dfrac{1}{b}\sqrt{\dfrac{3(b-1)}{b}}\right)$ (3.6)
@@
 $\mbox{Tr}(E_{2,3})$ $=-\dfrac{2c^{2}b-3c^{2}+b^{2}}{cb},\qquad\mbox{Det}(E_{2,3})=2(b-1).$
-It is easy to see that $E_{1}$ is a hyperbolic saddle for $b>1$ (since $\mbox{Det}(E_{1})<0$).
+It can be easily shown that $E_{1}$ is a hyperbolic saddle for $b>1$ (since $\mbox{Det}(E_{1})<0$).
 Using (3.5), the jacobian matrix of the vector field $f_{(0,1,\pm 1)}$ at $E_{1}\equiv E_{2}\equiv E_{3}$ is given by:
-(±$\left(\begin{array}[]{cc}\pm 1&\pm 1\\ \\ \mp 1&\mp 1\end{array}\right).$ (3.10)
-It is easy to verify that the matrix (3.10) is non-hyperbolic and has a double zero eigenvalue.
+$\left(\begin{array}[]{cc}\pm 1&\pm 1\\ \\ \mp 1&\mp 1\end{array}\right).$ (3.10)
+It is straightforward to see that the matrix (3.10) is non-hyperbolic and has a double zero eigenvalue.
 Our main result relies on the existence of a Double-zero bifurcation (DZ) with $\mathbb{Z}_{2}(\kappa)$–symmetry of $f_{(0,1,\pm 1)}$ at the singularity $E_{1}$.
@@
 $=$ $\left(y,-\frac{1}{3}(x\pm y)^{3}\right),$
-so that the equations in Jordan form are given by
+so that the equations in Jordan form are represented by
 $\left\{\begin{array}[]{l}\dot{x}=y\\ \\ \dot{y}=-\frac{1}{3}(x\pm y)^{3},\end{array}\right$. (3.11)
-which contains part of the truncated form of degree 3 of the versal deformation of a Double-zero (DZ) bifurcation with $\mathbb{Z}_{2}$–symmetry [27, pp.
-400, Case $c=-1$].
+which contains part of the truncated form of degree 3 of the versal deformation of a Double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry [27, pp.
+400, Case c=1c=-1].
 The next result deals with the description of all bifurcation curves passing through the bifurcation point $(b,c)=(1,\pm 1)$ and that characterises a DZ bifurcation of $f_{(0,b,c)}$ with $\mathbb{Z}_{2}(\kappa)$–symmetry.
@@
 Theorem 3.4.
-With respect to the vector field $f_{(0,b,c)},$ the equilibrium $E_{1}\equiv E_{2}\equiv E_{3}$ undergoes a DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1).$ The global representations of the transition/bifurcation curves in the space of parameters $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$ are as follows (schematic lines for $a=0$ have been plotted in Figures 1 and 3):
+Regarding to the vector field $f_{(0,b,c)},$ the equilibrium $E_{1}\equiv E_{2}\equiv E_{3}$ undergoes a DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1).$ The global representations of the transition/bifurcation curves in the parameters space $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$ are as follows (schematic lines for $a=0$ have been plotted in Figures 1 and 3):
   1. (1)
@@
   3. (3)
-    The equilibrium $E_{1}$ undergoes a subcritical (with respect to $b$) Hopf bifurcation along part of the parabola:
+    The equilibrium $E_{1}$ undergoes a supercritical (with respect to $b$) Hopf bifurcation along part of the parabola:
     $T_{H}^{1}=\left\{(b,c):b=c^{2},\,|c|<1\right\}.$
@@
   5. (5)
-    The equilibria $E_{2}$ and $E_{3}$ undergo a supercritical (with respect to $b$) Hopf bifurcation along the algebraic curves:
+    The equilibria $E_{2}$ and $E_{3}$ undergo a subcritical (with respect to $b$) Hopf bifurcation along the algebraic curves:
     $T^{2,3}_{H}=\left\{(b,c):b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>1\right\}.$
@@
     $DH:\quad 0=\frac{7b^{2}+10bc^{2}-17c^{2}}{15c^{3}}\sqrt{\frac{-7b^{2}+5bc^{2}+2c^{2}}{5b}},\quad b\in(0,c).$
+Remark 3.5.
+For every pair $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$, the origin is always an equilibrium point.
+When $b<0$, two additional nontrivial equilibria appear; they move away from the origin and diverge to infinity as $b\to 0^{-}$.
+For $0<b<1$, the origin is the only equilibrium.
+Thus, although the number of equilibria changes as the line $b=0$ is crossed, this change results from equilibria escaping to (or emerging from) infinity rather than from a local bifurcation at the origin.
 Remark 3.6.
@@
 The notation $T_{P}$, $T_{H}^{1}$, $T_{F}^{1}$, $T_{P}$ $T_{H}^{2,3}$, $T_{F}^{2,3}$ follows from Theorem 3.4 and sd, uf, un, sf, and sn refer to saddle, unstable focus, unstable node, stable focus and stable node.
-The curve $b=0$ corresponds to a subcritical pitchfork bifurcation from where the equilibria $E_{2}$ and $E_{3}$ (3.6) collapse into $E_{1}$.
 Remark 3.7.
-Theorem 3.4 generalizes the analysis of [8] and [21] by locating a codimension 2 bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$ for the vector field $f_{(0,b,c)}.$ For the sake of completeness, we plot a computer-assisted global phase portraits on Table 1 of Appendix.
-Figure 8 sketches the main features of the phase portraits in the different regions – within each region of the figure, the dynamics is $C^{1}$–conjugated to the respective figure.
+Theorem 3.4 generalizes the analysis of [8] and [21] by locating a codimension 2 bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$ for the vector field $f_{(0,b,c)}.$ For the sake of completeness, we present local numerical phase portraits on Table 1 of Appendix.
+All of them agree well with the theory.
 Proof of Theorem 3.4.
@@
 Let $b=1+\mu$, where $\mu\in\mathbb{R}$ is an additional bifurcation parameter.
-For $\mu=0$ the Jacobian matrix (3.5) associated with $E_{1}$ admits eigenvalues: 00 and $\frac{c^{2}-1}{c}$, $c\neq 0,1$.
+For $\mu=0$, the Jacobian matrix (3.5) associated with $E_{1}$ admits eigenvalues: 0 and $\frac{c^{2}-1}{c}$, $c\neq 0,1$.
 Define the diffeomorphism $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ as
-    $\mathbf{x}\equiv(x,y)\mapsto\left(\frac{c^{2}y-x}{c^{2}-1},\frac{x-y}{c^{2}-1}\right).$
-[…]
+    $\mathbf{x}\equiv(x,y)\mapsto\left(\frac{c^{2}y-x}{c^{2}-1},-\frac{y-x}{c^{2}-1}\right).$
+    In the new coordinates $\mathbf{x}=(x,y),$ where $\mathbf{0}=(0,0)$, let
     $\bar{f}_{(0,b,c)}(\mathbf{x})\equiv f_{(0,b,c)}(\mathbf{x})-Df_{(0,b,c)}(\mathbf{0})\mathbf{x}.$
@@
     $\left\{\begin{array}[]{l}\dot{x}=-\frac{c}{3(c^{2}-1)^{3}}\Big(3c^{2}x^{2}y-x^{3}-3c^{4}xy^{2}+c^{6}y^{3}+3(c^{2}-1)^{2}\mu x-3(c^{2}-1)^{2}\mu y\Big),\\ \\ \dot{y}=-\frac{1}{c(c^{2}-1)^{3}}\Big(3c^{4}x^{2}y-c^{2}x^{3}-3c^{6}xy^{2}+c^{8}y^{3}+3(c^{2}-1)^{2}\mu x-3(c^{2}-1)^{2}\mu y-3(c^{2}-1)^{4}y\Big).\end{array}\right$.
-    We aim to find a center manifold in the form of $h(x,\mu):={\alpha_{0}}^{2}\mu+\alpha_{1}x^{2}+\alpha_{2}x\mu.$ Since the center manifold must satisfy the equation ([27, Equation 3.2.7])
+    We aim to find a center manifold in the form of
+    $h(x,\mu):={\alpha_{0}}\mu^{2}+\alpha_{1}x^{2}+\alpha_{2}x\mu+\dots.$
+    Since the center manifold must satisfy the equation (cf. [27, Equation 3.2.7])
     $\dot{x}\frac{\partial h}{\partial x}-\dot{y}\big|_{(x,h(x,\mu))}=0,$
@@
 If $c>1$, then the bifurcation is supercritical with respect to $\mu$ and also with respect to $b$ in the original equation, as depicted in Figure 2.
-The analysis for the case $c>1$ is different from the case $10<c<$ 1, although the conclusion is the same.
+The analysis for the case $c>1$ is different from the case $0<c<1$, although the argument runs along the same lines.
     Figure 2. Illustration of the Pitchfork bifurcation of (3.14) for $c>1$ and close to $\mu=0(\Leftrightarrow b=1)$. (a) $\mu<0$, (b) $\mu=0$ and (c) $\mu>0$.
@@
     Let $b=c^{2}$ and $|c|<1$.
-The Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ has a pair of purely imaginary eigenvalues $i\sqrt{1-c^{2}}$ with $-1<c<1.$ Let $b=c^{2}+\mu$ where $\mu\in\mathbb{R}$ is a bifurcation parameter.
+The Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ has a pair of purely imaginary eigenvalues $i\sqrt{1-c^{2}}$ with 1<c<1.-1<c<1.
+Let $b=c^{2}+\mu$ where $\mu\in\mathbb{R}$ is a bifurcation parameter.
 The eigenvalues associated with the Jacobian matrix at $E_{1}$ for $\mu\neq 0$
@@
     this means that there is a non-degenerate Hopf bifurcation at $\mu=0$ ([17, pp. 299–300]). By using the linear transformation
-    $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}&\dfrac{1}{2}\dfrac{c}{\sqrt{1-c^{2}}}\\ 0&-\dfrac{1}{2}\dfrac{1}{c\sqrt{1-c^{2}}}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix},$
+    $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}&\dfrac{c}{2\sqrt{1-c^{2}}}\\ 0&-\dfrac{1}{2c\sqrt{1-c^{2}}}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}$,
     the system (3.1) turns into its Jordan form
     $\left\{\begin{array}[]{l}f^{1}:\,\dot{u}=-\frac{1}{12}c{u}^{3}-\frac{1}{4}\frac{{c}^{2}}{\sqrt{1-{c}^{2}}}v{u}^{2}-\frac{1}{4}{\frac{{c}^{3}}{1-{c}^{2}}}{v}^{2}u-\frac{1}{12}\frac{{c}^{4}}{\left(1-{c}^{2}\right)^{3/2}}{v}^{3}-\sqrt{1-{c}^{2}}v,\\ \\ f^{2}:\,\dot{v}=\sqrt{1-{c}^{2}}u.\end{array}\right$.
-    By calculating the first Lyapunov coefficient ([26, Equations (2) and (3)]), we have:
-    $l:=\dfrac{1}{16\omega}\left(R_{1}+\omega R_{2}\right),$
+    The first Lyapunov coefficient depends on $c$ and is given by:
+    $\ell(c)=\dfrac{1}{16\omega}\left(R_{1}+\omega R_{2}\right),$
     where
-    $\omega:=$ $\sqrt{1-{c}^{2}}$
+    $\omega:=$ $\sqrt{1-{c}^{2}}\neq 0$
     $R_{1}:=$ $f^{1}_{uv}\left(f^{1}_{uu}+f^{1}_{vv}\right)-f^{2}_{uv}\left(f^{2}_{uu}+f^{2}_{vv}\right)-f^{1}_{uu}f^{2}_{uu}+f^{1}_{vv}f^{2}_{vv}\quad\text{and}$
     $R_{2}:=$ $f^{1}_{uuu}+f^{1}_{uvv}+f^{2}_{uuv}+f^{2}_{vvv}.$
-    Therefore, using Maple, we get
-    $l=\frac{c}{32(c^{2}-1)}<0$
-    and we may conclude that the limit cycle is stable for $10<c<$ 1 and the bifurcation is subcritical – [27, Remarks 1 and 2, pp.
+    Using Maple, we get
+    $\ell(c)=\frac{c}{32(c^{2}-1)}<0$
+    and we may conclude that the limit cycle is stable for $0<c<1$ and the bifurcation is supercritical – [27, Remark 1, pp.
 383].
-In this case $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable for $\mu>0$ with an asymptotically stable periodic solution for $\mu<0$.
-The same conclusion holds for $b$.
+In this case, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ ($\Leftrightarrow b>c^{2}$) and unstable for $\mu<0$ ($\Leftrightarrow b<c^{2}$), with an asymptotically stable periodic solution for $\mu<0$ ($\Leftrightarrow b<c^{2}$).
   4. (4)
@@
   5. (5)
-    Using linear analysis, one knows that Hopf bifurcation exists only if $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0$ and $\mbox{Det}(E_{2}),\mbox{Det}(E_{3})>0.$ Indeed,
-    $1\displaystyle 3{c}^{2}-{b}^{2}-2{c}^{2}b=0\Leftrightarrow b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>$ 1 .
-    Figure 3.
-The bifurcation curves associated to (3.1) for Case A.
-Within each depicted region, phase portraits are $C^{1}$–equivalent.
-(a) The global phase portraits on the Poincaré disc corresponding to Regions a–p have been sketched in Figures 8.
-(b) The local portraits corresponding to Regions 1–28 have been plotted in Table 1 of Appendix.
+    Using linear analysis, one knows that Hopf bifurcation exists only if $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0$ and $\mbox{Det}(E_{2}),\mbox{Det}(E_{3})>0.$ Indeed, we have:
+    $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0\mbox{ and }b>1$
+    $\Leftrightarrow$ $3{c}^{2}-{b}^{2}-2{c}^{2}b=0\mbox{ and }b>1$
+    $\Leftrightarrow$ $b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>1.$
+    Figure 3. The bifurcation curves associated to (3.1) for Case A. The local portraits corresponding to Regions 1–28 have been plotted in Table 1 of Appendix.
     The inequality $b>1$ serves to ensure that the equilibria $E_{2}$ and $E_{3}$ exist.
-    Let $b=c\left(-c\pm\sqrt{c^{2}+3}\right).$ The Jacobian matrix at $E_{2,3}$ has a pair of purely imaginary eigenvalues ±$\pm i\sqrt{2c\sqrt{c^{2}+3}-2-2c^{2}}$ with $c>1.$ Let $b=-c^{2}\pm c\sqrt{c^{2}+3}+\mu$ where $\mu\in\mathbb{R}$ denotes the bifurcation parameter.
+    Let $b=c\left(-c\pm\sqrt{c^{2}+3}\right).$ The Jacobian matrix at $E_{2,3}$ has a pair of purely imaginary eigenvalues $\pm i\sqrt{2c\sqrt{c^{2}+3}-2-2c^{2}}$ with $c>1.$ Let $b=-c^{2}\pm c\sqrt{c^{2}+3}+\mu$ where $\mu\in\mathbb{R}$ denotes the bifurcation parameter.
 The eigenvalues, $\lambda(\mu),$ associated with the Jacobian matrix at $E_{2,3}$
@@
     $\lambda(\mu)=\dfrac{-2c\sqrt{c^{2}+3}\mu-\mu^{2}\pm\sqrt{\sigma}}{2c(\mu-c^{2}+c\sqrt{c^{2}+3})}$
-    where :
+    where
     $\sigma=\mu^{4}+4c(\sqrt{c^{2}+3}-2c)\mu^{3}-4c^{2}(6c\sqrt{c^{2}+3}-7c^{2}-5)\mu^{2}$
     $+8c^{3}(2\sqrt{c^{2}+3}(3c^{2}+1)-6c^{3}-11c)\mu-8c^{4}(c\sqrt{c^{2}+3}\left(4c^{2}+5\right)-11c^{2}-4c^{4}-3).$
-    There is a pair of complex conjugate eigenvalues when $\sigma<0$ and
+    There is a pair of complex conjugate eigenvalues when $\sigma<0$. In addition, we have:
     $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}=-\frac{\sqrt{c^{2}+3}}{c(\sqrt{c^{2}+3}\mp c)}.$
-    Since the first Lyapunov coefficient is
-    $l={\frac{c\left(\left(4c^{2}+3\right)\left(\sqrt{c^{2}+3}-c\right)-3c\right)\left(\sqrt{c^{2}+3}c\left(21+44c^{2}+16c^{4}\right)-c^{2}\left(69+68c^{2}+16c^{4}\right)-9\right)}{32\left(\sqrt{c^{2}+3}c-c^{2}-1\right)^{2}\left(\sqrt{c^{2}+3}-c\right)^{7}}},$
-    we may deduce, using Maple, that these limit cycles are unstable for $c>0$.
+    The first Lyapunov coefficient $\ell$ associated to $f_{(0,b,c)}$ at $E_{2,3}$ depends on $c$ and is given explicitly by (computations in Maple):
+    $\ell(c)={\frac{c\left(\left(4c^{2}+3\right)\left(\sqrt{c^{2}+3}-c\right)-3c\right)\left(\sqrt{c^{2}+3}c\left(21+44c^{2}+16c^{4}\right)-c^{2}\left(69+68c^{2}+16c^{4}\right)-9\right)}{32\left(\sqrt{c^{2}+3}c-c^{2}-1\right)^{2}\left(\sqrt{c^{2}+3}-c\right)^{7}}},$
+    Since $\lim_{c\to 1^{+}}\ell(c)=\lim_{c\to+\infty}\ell(c)=+\infty$ and $\frac{\partial\ell(c)}{\partial c}=0\quad\text{at}\quad c^{\star}\approx 1.315$, then $\ell$ remains positive in the neighbourhood of $c^{\star}$.
+Therefore, the corresponding limit cycles for $c>1$ are unstable and the bifurcation is subcritical.
   6. (6)
     This item follows from [8, pp. 201] and [21, Formula (3.2.29)].
-Figure 4.
-The classification of the equilibrium point $E_{1}$ and the corresponding bifurcation curves for (3.1) in Case B.
-In this figure, un, sn, uf and sf refer to unstable (stable) node and unstable (stable) focus, respectively.
-(a) The local phase portraits associated with Regions 1–8 have been plotted in Figure 2.
-(b) The global phase portraits on the Poincaré disc associated with Regions I–IV have been plotted in Figure 9.
 Remark 3.8.
-Numerically, we may observe that the non-degenerate Hopf bifurcation of items (3) and (5) of Theorem 3.4 are, respectively, subcritical and supercritical.
+Numerically, we may observe that the non-degenerate Hopf bifurcation of items (3) and (5) of Theorem 3.4 are, respectively, supercritical and subcritical.
 The bifurcation $T_{H}^{1}$ yields a stable periodic solution, say $C^{s}$, and $T_{H}^{2,3}$ generates two unstable and $\kappa$–symmetric periodic solutions – confirm on Table 1 of Appendix.
@@
 Theorem 3.9.
-With respect to the vector field $f_{(a,0,c)}$ of (3.1), the global representations of the transition/bifurcation curves in the space of parameters $\0(a,c)\in\mathbb{R}\times\mathbb{R}\backslash\{0\}$ are as follows:
+[…]
   • I.
@@
     $T^{b0}_{F}=\left\{(a,c):c=\pm\frac{2}{a^{2}-1}\right\}.$
+    Figure 4.
+The classification of the equilibrium point $E_{1}$ and the corresponding bifurcation curves for (3.1) in Case B.
+In this figure, un, sn, uf and sf refer to unstable (stable) node and unstable (stable) focus, respectively.
+The local phase portraits associated with Regions 1–8 have been plotted in Table 2.
   • II.
@@
     $T^{b0}_{H}=\left\{(a,c):a=\pm 1\right\}.$
-    The Hopf bifurcation is supercritical at $a=-1$ and subcritical at $a=1$.
-Schematic bifurcation lines for $b=0$ have been plotted in Figure 4.
-Illustrative pictures in the phase space have been depicted in Figure 9 in the case where the eigenvalues of (3.5) at $E_{1}$ are complex non-real.
+    The Hopf bifurcation is supercritical(resp. subcritical) if $c>0$ (resp. $c<0$).
+Schematic bifurcation lines for $b=0$ have been plotted in Figure 4. The local phase portraits associated with Regions 1–8 have been plotted in Table 2
 Proof.
@@
   • II.
-    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues ±$\pm i.$ Let $a=\pm 1+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter.
+    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues $\pm i.$ Let $a=\pm 1+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter.
 The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix (3.5) at $E_{1}$ are (computations performed with Maple):
     $\lambda(\mu)=\mp c\mu-\frac{1}{2}c{\mu}^{2}\pm\frac{1}{2}\sqrt{4{c}^{2}{\mu}^{2}\pm 4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4}.$
-    For $\mu$ close to zero, we have $04{c}^{2}{\mu}^{2}+4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4<$ 0, which means that $\lambda(\mu)$ is complex non-real.
-Furthermore $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}\big{|}_{\mu=0}=\mp c\neq 0.$ Following [26, Equations (2) and (3)], the first Lyapunov coeficient is $l=-\frac{1}{32}c$ which is negative for $c>0.$ Hence the bifurcation is subcritical at $a=1$ and supercritical for $a=-1$.
-    For the case $a=-1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable focus for $\mu>0$ with an asymptotically stable periodic solution emerging for $\mu>0$.
+    For $\mu$ close to zero, we have $4{c}^{2}{\mu}^{2}+4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4<0$, which means that $\lambda(\mu)$ is complex non-real. Furthermore
+    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}\big|_{\mu=0}=\mp c\neq 0.$
+    The first Lyapunov coefficient depends on $c$ and is given by
+    $\ell(c)=-\frac{1}{32}c$
+    which has opposite sign to that of $c\neq 0$. Hence the bifurcation is subcritical for $c<0$ and supercritical for $c>0$.
+    For the case a=1a=-1, $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable focus for $\mu>0$ with an asymptotically stable periodic solution emerging for $\mu>0$.
 For the case $a=1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ and unstable focus for $\mu<0$ with an asymptotically stable periodic solution emerging $\mu>0$.
 This finishes the proof.
-Remark 3.9.
-We may observe in Figure 9 that, for $c>0$ (resp.
-$c<0$), the Hopf bifurcation (with respect to $a$) is supercritical for $a=-1$ giving rise to a stable (resp.
-unstable) periodic solution which disappears at a subcritical Hopf bifurcation at $a=1,$ see regions 5-8 of Table 2.
-3.3. Case C ($a>0,$ $10<b<$ 1)
+3.3. Case C ($a>0,$ $0<b<1$)
 If $(x^{\star},y^{\star})$ is an equilibrium of (3.1) under the conditions $a>0,$ $0<b<1,$ then:
@@
 $\Delta=\frac{4}{3}c^{4}\left(1-\frac{1}{b}\right)^{3}-3\frac{c^{4}a^{2}}{b^{2}}.$
-For $10<b<$ 1, the discriminant satisfies $\Delta<0$.
-According to Cardano’s formula, this implies that the polynomial $H(x)$ has exactly one real root, denoted by $x^{\star}.$ Then the equilibrium of (3.1) is explicitly given by
+For $0<b<1$, the discriminant satisfies $\Delta<0$.
+According to Cardano’s formula, this implies that the polynomial $H(x)$ has exactly one real root, denoted by $x^{\star}.$ Then the equilibrium for system (3.1) is explicitly given by
 $\left(x^{\star},\frac{a-x^{\star}}{b}\right)=(x^{\star},y^{\star})=:E_{1}.$
@@
 Proposition 3.10.
-The following assertions hold for the vector field $f_{(a,b,c)}$ of (3.1), under the conditions $a>0,$ and $10<b<$ 1:
+The following assertions hold for the vector field $f_{(a,b,c)}$ of (3.1), under the conditions $a>0,$ and $0<b<1$:
   1. (1)
@@
   2. (2)
-    A non-degenerate Hopf bifurcation occurs at $c^{2}=\frac{b}{1-{x^{*}}^{2}}.$ The resulting limit cycle is attracting when $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and repelling for $c<\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
-Note that $x^{\star}\neq\pm 1$ (if $x^{\star}=\pm 1$, then $b=0$, which is a contradiction).
-Illustrative pictures in the phase space have been depicted in Figure 10.
-As before, we use the notation $\mbox{Tr}(E_{1})$ and $\mbox{Det}(E_{1})$ the trace and determinant operators of $Df_{(a,b,c)},$ evaluated at $E_{1}$.
+    A non-degenerate Hopf bifurcation occurs at $c^{2}=\frac{b}{1-{x^{*}}^{2}}.$ The numerical simulations indicate the resulting limit cycle is attracting when $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and repelling for $c<-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
+Note that $x^{\star}\neq\pm 1$ (if $x^{\star}=\pm 1$ then $b=0$, which is a contradiction).
+As before, we use the notation $\mbox{Det}(E_{1})$ and $\mbox{Tr}(E_{1})$ for the determinant and trace operators of $Df_{(a,b,c)},$ evaluated at $E_{1}$.
 Proof.
@@
   2. (2)
-    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues ±$\pm i\sqrt{b{x^{*}}^{2}-b+1}.$ Let $b=c^{2}\left(1-{x^{*}}^{2}\right)+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter.
+    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues $\pm i\sqrt{b{x^{*}}^{2}-b+1}.$ Let $b=c^{2}\left(1-{x^{*}}^{2}\right)+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter.
 The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix at $E_{1},$
@@
 The statement of Proposition 3.10 says nothing about the criticality of the Hopf Bifurcation.
-Analytically, it is very difficult to compute the first Lyapunov exponent of (3.1) at an equilibrium whose explicit expression we do not know.
-However, numerics of Figure 10 suggest that the Hopf bifurcation is supercritical for for $c=\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and subcritical for $c=-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
-4. Canards for the singular case $c\to+\infty$ in Cases A and B
-The goal of this section is the study of the dynamics of (3.1) when $|c|\to+\infty$.
+Analytically, it is very difficult to compute the first Lyapunov coefficient of (3.1) at an equilibrium whose explicit expression we do not know.
+However, numerics suggest that the Hopf bifurcation is supercritical for $c=\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and subcritical for $c=-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
+4. Analysis of the singular case $c\to+\infty$ in Cases A and B
+The aim of this section is the study of the dynamics of (3.1) when $|c|\to+\infty$.
 Many results concerning the asymptotic behaviour of the solution of the FN model as $|c|\to+\infty$ are available in the literature – see [21, 22, 23].
 Our concern is to match these results with the local bifurcation study performed in Section 3.
-We are particularly interested in Cases A and B whose results have been described in Theorems 3.4 and 3.8.
-By changing the time scale of (3.1), considering $\tau\to t/c$ and $\varepsilon\to 1/c^{2},$ we obtain an equivalent version of (3.1) as follows (written with respect to the slow time $\tau$):
-$\left\{\begin{array}[]{l}\varepsilon\dot{x}=\displaystyle\left[y-\left(\frac{x^{3}}{3}-x\right)\right]=:f(x,y,\varepsilon)\\ \\ \dot{y}=\displaystyle-(x-a+by)=:g(x,y,\varepsilon)\end{array}\right.$ (4.1)
-where $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}.$ For $0<\varepsilon\ll 1,$ or equivalently $|c|\gg 1,$ system (4.1) is what we call a fast-slow system (see [10, 14] and references therein).
-A fast-slow system is a system of differential equations in which some variables have their derivatives with larger magnitude than others.
+We are particularly interested in Cases A and B whose results have been described in Theorems 3.4 and 3.9.
+By changing the time scale of (3.1), taking into account $\tau=t/c$ and $\varepsilon=1/c^{2},$ we get an equivalent form for (3.1) as follows (written with respect to the slow time $\tau$):
+$\left\{\begin{array}[]{l}\varepsilon\dot{x}=\displaystyle\left[y-\left(\frac{x^{3}}{3}-x\right)\right]=:f(x,y,\varepsilon)\\ \\ \dot{y}=\displaystyle-(x+by-a)=:g(x,y,\varepsilon)\end{array}\right$.
+(4.1)
+[…]
+A fast-slow system consists of differential equations where the rates of change of some variables are much greater than those of others.
 This leads to a two time-scale system.
 The general approach to this type of system starts by grouping the variables in two disjoint sets: fast variable $x$ and slow variable $y$.
-By allowing the dynamics to be separated into two distinct phases, slow and fast, we can study each one independently and understand its underlying dynamics.
-This separation is introduced in system (4.1) via the parameter $\varepsilon.$
+By decomposing the dynamics into slow and fast phases, each can be studied independently, allowing a clearer understanding of its underlying behaviour.
+The parameter $\varepsilon$ introduces this separation in system (4.1).
 4.1. The critical manifold and bifurcations
 We first analyse the singular case $\varepsilon=0.$ We follow the analysis performed in [10, 12, 13]. The critical manifold associated with (4.1) is the set
-$\mathcal{C}_{0}=\{(x,y)\in\mathbb{R}^{2}:f(x,y,0)=0\}=\{(x,y)\in\mathbb{R}^{2}:y=x^{3}/3-x\}$ (4.2)
+$\mathcal{C}_{0}=\{(x,y)\in\mathbb{R}^{2}:f(x,y,0)=0\}=\left\{(x,y)\in\mathbb{R}^{2}:y=\frac{x^{3}}{3}-x\right\}.$ (4.2)
 Definition 4.1.
@@
   1. (1)
-    The subset $S\subset\mathcal{C}_{0}$ is said to be normally hyperbolic if for all $p=(x,y)\in{S}$, we have:
-[…]
+    The subset $S\subset\mathcal{C}_{0}$ is called normally hyperbolic provided that for all $e=(x_{0},y_{0})\in{S}$, we get:
+    $\frac{\partial f}{\partial x}\left(e,0\right)\neq 0$
   2. (2)
-    A normally hyperbolic set $S$ is said to be attracting (resp.
-repelling) if, for all $p\in{S}$, all eigenvalues of $\frac{\partial f}{\partial x}\left(p,0\right)$ have negative (resp.
+    A normally hyperbolic set $S$ is called attracting (resp.
+repelling) when, for every $e\in{S}$, all eigenvalues of $\frac{\partial f}{\partial x}\left(e,0\right)$ possess negative (resp.
 positive) real part, respectively.
 Definition 4.2.
-With respect to (4.1) with critical manifold $\mathcal{C}_{0}$, the point $p\in\mathcal{C}_{0}$ is said to be a fold point if:
-$\frac{\partial f}{\partial x}\left(p,0\right)=0\,,\quad\frac{\partial^{2}f}{\partial x^{2}}\left(p,0\right)\neq 0\quad\text{ and }\quad\frac{\partial f}{\partial y}\left(p,0\right)\neq 0.$
-If $g\left(p,0\right)\neq 0$, the fold point is called regular.
-Coming back to system (4.1), we have $f\left(x,y,0\right)=y-x^{3}/3+x$ and, for $p=(x,y)\in\mathcal{C}_{0}$, one has:
-$\frac{\partial f}{\partial x}\left(p,0\right)=x^{2}-1\neq 0\qquad\text{ if }\,\,x\neq\pm 1.$
-Therefore, if $b=3/2$, the equilibria $E_{2}=(-1,2/3)$ and $E_{3}=(1,-2/3)$ are fold points (according to the Definition 4.2) since:
+With respect to (4.1) with critical manifold $\mathcal{C}_{0}$ (see (4.2)), the point $e\in\mathcal{C}_{0}$ is called a fold point if:
+$\frac{\partial f}{\partial x}\left(e,0\right)=0\,,\quad\frac{\partial^{2}f}{\partial x^{2}}\left(e,0\right)\neq 0\quad\text{ and }\quad\frac{\partial f}{\partial y}\left(e,0\right)\neq 0.$
+If $f\left(e,0\right)\neq 0$, the fold point is termed regular.
+Coming back to system (4.1), we get $f\left(x,y,0\right)=x-\frac{x^{3}}{3}+y$ and, for $e=(x_{0},y_{0})\in\mathcal{C}_{0}$, one has:
+$\frac{\partial f}{\partial x}\left(e,0\right)=1-x_{0}^{2}\neq 0\qquad\text{ if }\,\,x_{0}\neq\pm 1.$
+Therefore, if $b=3/2$, the equilibria $E_{2}=\left(-1,2/3\right)$ and $E_{3}=(1,-2/3)$ are fold points (according to the Definition 4.2) since:
 $\frac{\partial f}{\partial x}(E_{2,3},0)=0,\quad\frac{\partial^{2}f}{\partial^{2}x}(E_{2,3},0)\neq 0\quad\text{and}\quad\frac{\partial f}{\partial y}(E_{2,3},0)\neq 0$
-and $S_{0}=\mathcal{C}_{0}\setminus\left\{E_{2},E_{3}\right\}$ is normally hyperbolic. We may split $S_{0}$ into three disjoint open subsets:
+and $S_{0}=\mathcal{C}_{0}\setminus\left\{E_{2},E_{3}\right\}$ is normally hyperbolic according to Definition 4.1. We may split $S_{0}$ into three disjoint open subsets:
 $\mathcal{C}_{0\text{L}}$ $=\mathcal{C}_{0}\cap{\textstyle\left\{\left(x,y\right)\in\mathbb{R}^{2}:x<-1\right\}},$
@@
 $S_{0}=\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{M}}\cup\mathcal{C}_{0\text{R}},$ (4.3)
-where $\mathcal{C}_{0\text{L}}$ and $\mathcal{C}_{0\text{R}}$ are repelling subsets and $\mathcal{C}_{0\text{M}}$ is attracting.
-This means that in the fast flow, trajectories of (4.1) move backwards either to $\mathcal{C}_{0\text{L}}$ or $\mathcal{C}_{0\text{R}}$ and towards to $\mathcal{C}_{0\text{M}}$ – see Figure 5.
+where $\mathcal{C}_{0\text{L}}$ and $\mathcal{C}_{0\text{R}}$ are attracting and $\mathcal{C}_{0\text{M}}$ is repelling.
+As a result, trajectories of (4.1) in the fast flow move towards either to $\mathcal{C}_{0\text{L}}$ or $\mathcal{C}_{0\text{R}}$ and backwards to $\mathcal{C}_{0\text{M}}$ – see Figure 5.
 Figure 5. (a) Illustration of the critical manifold associated to system (4.1). (b) Illustration of the dynamics of system (4.1) for the case $a=0,$ when there is a unique equilibrium.
-Yet in the singular case ($\varepsilon=0$), depending on the values of $a,b\in\mathbb{R}$, there exists at least one and at most three equilibria, as result of the intersection of the cubic curve, $f(x,y,0)=0,$ with the line defined by $x-a+by=0$ ($\Leftrightarrow\dot{y}=0).$ Moreover, when the system has three equilibria only one of the following scenarios may occur (cf.
+Yet in the singular case ($\varepsilon=0$), depending on the values of $a,b\in\mathbb{R}$ in system (4.1), one, two, or three equilibria might arise as a consequence of the cubic curve’s intersection, $f(x,y,0)=0,$ and the line defined by xa+by=0x-a+by=0 ($\Leftrightarrow\dot{y}=0).$ If the system possesses three equilibria, only one of the following cases arise (cf.
 [10]):
 Case 1:
-either all three equilibria belong to $\mathcal{C}_{0\text{M}}$ or else
+each of the three equilibria lies in $\mathcal{C}_{0\text{M}}$ or else
 Case 2:
-they are each one in a distinct region, $\mathcal{C}_{0\text{L}},$ $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$.
-The independence of the time scales when $\varepsilon=0$ makes the system simpler to understand.
-Fenichel’s Theorem [6] says that, near the normally hyperbolic part of $\mathcal{C}_{0}$, when $0<\varepsilon\ll 1,$ the dynamics of the perturbed system is similar to that of the singular system, with a deviation of $\mathcal{O}\left(\varepsilon\right).$ The smaller the $\varepsilon,$ the more similar the system trajectories are to those described in the singular system ($\varepsilon=0$).
-Fenichel’s Theorem guarantees that for $0<\varepsilon\ll 1,$ there exists a set $S_{\varepsilon}$, $C^{1}$–close to $S_{0}$ (in the Haussdorf topology), that exhibits the same behaviour as $S_{0}$.
+each of the equilibria belongs to a different region $\mathcal{C}_{0\text{L}}$, $\mathcal{C}_{0\text{M}}$, and $\mathcal{C}_{0\text{R}}$.
+When $\varepsilon=0,$ the independence of the time scales simplifies the system’s analysis.
+Fenichel’s Theorem [6] says that, near the normally hyperbolic part of $\mathcal{C}_{0}$, for $0<\varepsilon\ll 1,$ the perturbed system exhibits dynamics “analogous” to the singular system (in the sense of differentiable conjugacy), with a deviation of the order $\mathcal{O}\left(\varepsilon\right).$ Decreasing $\varepsilon$ makes the system trajectories increasingly similar to those of the singular system ($\varepsilon=0$).
+According to Fenichel’s Theorem, it is ensured that when $0<\varepsilon\ll 1,$ there is a set $S_{\varepsilon}$, $\mathcal{O}\left(\varepsilon\right)$–close to $S_{0}$ (in the $\mathbb{C}^{1}$–topology), which is locally hyperbolic and attracting (if $S_{0}$ is normally hyperbolic and attracting).
 Definition 4.3.
-The set $S_{\varepsilon}$ is referred to as the slow manifold of system (4.1).
-When the system’s equilibrium is also a fold point, a supercritical (with respect to $b$) Hopf bifurcation occurs, giving rise to a periodic solution.
-In this section, the canard phenomenon described in [12, 13] bridges the local periodic solution from the Hopf bifurcation to the global limit cycle, as we proceed to explain.
+The set $S_{\varepsilon}$ is known as the slow manifold for (4.1).
+When the equilibrium of the system is a fold point, a subcritical (with respect to $b$) Hopf bifurcation arises, leading to the emergence of a periodic solution.
+In this section, the canard phenomenon described in [12, 13] links the local periodic solution generated by the Hopf bifurcation to the global limit cycle, like we proceed to explain (the canard point is correlated with the passage of a reduced problem equilibrium over the fold point under parameter variation – see Definition 4.6).
 Definition 4.4.
-A trajectory segment of a fast-slow system (4.1) is a canard if it stays within $\mathcal{O}\left(\varepsilon\right)$ distance to a repelling branch of a slow manifold for a time that is $\mathcal{O}\left(1\right)$ on the slow time scale $\tau=t\varepsilon$.
-Canards are related to equilibria of (4.1) that occur at fold points of the slow manifold. The next definitions deal with the general equation parametrised by $\lambda\in\mathbb{R}$:
+A trajectory segment of a fast-slow system (4.1) is a canard if it stays within a $\mathcal{O}\left(\varepsilon\right)$–distance to a repelling branch of a slow manifold for a time that is $\mathcal{O}\left(1\right)$ on the slow time scale $\tau=t\varepsilon$.
+Canards are related to equilibria of (4.1) that arise at fold points of the slow manifold. The next definitions deal with the general equation parametrised by $\lambda\in\mathbb{R}$:
 $\left\{\begin{array}[]{l}\varepsilon\dot{x}=f(x,y,\lambda,\varepsilon)\\ \\ \,\dot{y}=g(x,y,\lambda,\varepsilon)\end{array}\right$. (4.4)
-We now introduce the definition of singular fold point, which differs from that of a fold point – it is required to coincide with an equilibrium point of (4.1):
+We now define a singular fold point, a different concept from that of a fold point; it is necessary to coincide with an equilibrium point of (4.1):
 Definition 4.5.
-With respect to (4.4), let $p=\left(x,y\right)$ be a fold point. The point $p$ is called a singular fold if:
-$f\left(p,\lambda,0\right)=0,$ $\frac{\partial f}{\partial x}\left(p,\lambda,0\right)=0,$ (4.5)
-$\frac{\partial^{2}f}{\partial x^{2}}\left(p,\lambda,0\right)\neq 0,$ $\frac{\partial f}{\partial y}\left(p,\lambda,0\right)\neq 0\quad\text{and}\quad g\left(p,\lambda,0\right)=0.$
+With respect to (4.4), let $e=\left(x,y\right)$ denote a fold point. We refer to $e$ as a singular fold provided:
+$f\left(e,\lambda,0\right)=0,$ $\frac{\partial f}{\partial x}\left(e,\lambda,0\right)=0,$ (4.5)
+$\frac{\partial^{2}f}{\partial x^{2}}\left(e,\lambda,0\right)\neq 0,$ $\frac{\partial f}{\partial y}\left(e,\lambda,0\right)\neq 0\quad\text{and}\quad g\left(e,\lambda,0\right)=0.$
 Definition 4.6.
-A singular fold point $p$ is said to be regular if:
-$\frac{\partial g}{\partial x}\left(p,\lambda,0\right)\neq 0\quad\text{ and }\quad\frac{\partial g}{\partial\lambda}\left(p,\lambda,0\right)\neq 0.$ (4.6)
-The behaviour around a regular singular fold point is described in the next theorem.
-The existence of canards is coupled with the existence of a Hopf bifurcation, which in such systems is referred to as a singular Hopf bifurcation.
+A singular fold point $e$ (also known as a canard point) is called regular if:
+$\frac{\partial g}{\partial x}\left(e,\lambda,0\right)\neq 0\quad\text{ and }\quad\frac{\partial g}{\partial\lambda}\left(e,\lambda,0\right)\neq 0.$ (4.6)
+The dynamics near a regular singular fold point is explored in Theorem 4.7. The occurrence of canards is linked to a Hopf bifurcation, known in these systems as a singular Hopf bifurcation.
 Theorem 4.7 ([12, 13], adapted).
-Consider a fast-slow system of the form (4.4) where $\left(x,y\right)=\left(0,0\right)$ is a regular singular fold point for $\lambda=0$, and that may be written in the following form:
+Consider a fast-slow system of the form (4.4) where $\left(x,y\right)=\left(0,0\right)$ is canard point for $\lambda=0$, and
 $\dot{x}=-y\,l_{1}\left(x,y,\lambda,\varepsilon\right)+x^{2}\,l_{2}\left(x,y,\lambda,\varepsilon\right)+\varepsilon\,l_{3}\left(x,y,\lambda,\varepsilon\right)$ (4.7)
@@
 for $j\in\{1,2,4,5,6\}$.
-Assume that, for $\varepsilon=0$, there is a slow trajectory connecting the repelling and attracting regions of the critical manifold $\mathcal{C}_{0}$.
-Then, there exist $\varepsilon_{0}>0$ and $\lambda_{0}>0$ such that for $0<\varepsilon<\varepsilon_{0}$ and $|\lambda|<\lambda_{0}$, the system has an equilibrium point $p\in\mathbb{R}^{2}$ near the origin where $p\to\left(0,0\right)$ as $\left(\lambda,\varepsilon\right)\to\left(0,0\right)$.
-Then, there exists a smooth function $\lambda_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $\lambda$ that gives rise to a family of canards, asymptotically defined by:
-$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\left(B+A\right)\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
-and there exists a continuous function $\lambda_{H}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $\lambda$ of Hopf bifurcations in the system, asymptotically defined by:
-$\lambda_{H}\left(\sqrt{\varepsilon}\right)=-B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
+Suppose that, when $\varepsilon=0$, a slow trajectory connects attracting and repelling regions of the critical manifold $\mathcal{C}_{0}$.
+Therefore, there is $\varepsilon_{0}>0$ and $\lambda_{0}>0$ so that for $0<\varepsilon<\varepsilon_{0}$ and $|\lambda|<\lambda_{0}$, the system possesses an equilibrium point $p\in\mathbb{R}^{2}$ around $(0,0)$ where $p$ approaches $(0,0)$ as $\left(\lambda,\varepsilon\right)$ tends to $(0,0)$.
+Moreover, there exists a smooth map $\lambda_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that assigns every $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a real value $\lambda$, leading to a family of canards, defined asymptotically as
+$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\varepsilon\left(A+B\right)+\mathcal{O}\left(\varepsilon^{3/2}\right).$
+and In addition, there is a continuous map $\lambda_{H}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $\lambda$, associated to a Hopf bifurcations defined asymptotically as
+$\lambda_{H}\left(\sqrt{\varepsilon}\right)=-\varepsilon B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
 where
-$A=\dfrac{-\dfrac{\partial l_{1}}{\partial x}+3\dfrac{\partial l_{2}}{\partial x}-2\dfrac{\partial l_{4}}{\partial x}+2l_{6}}{8}\qquad B=\dfrac{\dfrac{\partial l_{3}}{\partial x}+l_{6}}{2},$
-where the functions $l_{i},\,i=1,\ldots,6$ and their partial derivatives are evaluated at the point $\left(x,y,\lambda,\varepsilon\right)=\left(0,0,0,0\right)$.
-The Hopf bifurcation is nondegenerate when $A\neq 0$, it is supercritical if $A<0$ and subcritical if $A>0$.
+$A=-\dfrac{\dfrac{\partial l_{1}}{\partial x}-3\dfrac{\partial l_{2}}{\partial x}+2\dfrac{\partial l_{4}}{\partial x}-2l_{6}}{8},\qquad B=\dfrac{\dfrac{\partial l_{3}}{\partial x}+l_{6}}{2},$
+and the functions $l_{i},$ for $1\leq i\leq 6$, and their partial derivatives are calculated at $\left(x,y,\lambda,\varepsilon\right)=\left(0,0,0,0\right)$.
+The latter bifurcation is non-degenerate when $A\neq 0$, it occurs for positive values of $\lambda-\lambda_{H}$ if $A<0$, and for negative values of $\lambda-\lambda_{H}$ if $A>0$.
 4.2. Case A $(a=0)$
-The curves $T_{F}^{2,3}$, $T_{H}^{2,3}$, $DH$ and SNL, introduced in Theorem 3.4 and Remark 3.5, can be parameterised by $c\in[1,+\infty)$.
-From now on, we restrict our attention to the portions of these curves lying within the region defined by $b,c\geq 1$ – see Figure 6.
-We are going to denote by $T_{F,+}^{2,3}\backslash\{(1,1)\}$ the connected component of $T_{F}^{2,3}\backslash\{(1,1)\}$ with highest values of $b$ for the same $c$, and $T_{F,-}^{2,3}\backslash\{(1,1)\}$ the other.
-It is straightforward to verify that, for sufficiently large values of $c$ in $[1,+\infty)$, the following holds:
-$T_{F,-}^{2,3}(c)\leq T_{H}^{2,3}(c)\leq DH(c)\leq SNL(c)\leq T_{F,+}^{2,3}.$ (4.8)
-All the curves admit $(b,c)=(1,1)$ as accumulation point.
-Denote by $A_{1}\leq A_{2}<A_{3}<A_{4}$ the limits when $c\to+\infty$ of the previous curves (when it exists), respectively.
-More specifically, we have:
+[…]
+$T_{F,-}^{2,3}(c)\leq T_{H}^{2,3}(c)\leq DH(c)\leq SNL(c).$ (4.8)
+All the curves pass through $(b,c)=(1,1)$. Denote by $A_{1}\leq A_{2}<A_{3}<A_{4}$ the limits when $c\to+\infty$ of the previous curves (when it exists), respectively. More specifically, we have:
 Figure 6.
 Partial scheme of the codimension 1 bifurcation curves emerging from the point $(b,c)=(1,1)$ associated with $f_{(0,b,c)}$ of (3.1), for $b,c>0$.
-The symbol $c^{\infty}$ represents the limit of the associated curves when $c\to+\infty$ (implicit compactification in $c$).
+The symbol $c^{\infty}$ represents “very large” values of $c$.
 The notation of the curves correspond to those of Theorem 3.4, for Case A.
@@
 Proposition 4.9.
-With respect to (4.4), there exists a smooth function $b_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $b$ that gives rise to a family of canards in the system, asymptotically defined by:
+With respect to (4.4), a smooth function $b_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ exists that assigns every $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $b$ leading to a family of canards in the system, asymptotically defined by:
 $b_{c}\left(\sqrt{\varepsilon}\right)=\frac{3}{2}+\frac{5}{4}\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).$
 Proof.
-The difference $b-3/2$ of system (4.4) is played by $\lambda$ of Theorem 4.7.
-The equilibrium $E_{3}$ is a regular singular fold point and there is a Hopf bifurcation point at $b=3/2$.
-In order to apply Theorem 4.7 we change variables and the parameter by
-$\left\{\begin{array}[]{l}\bar{x}=x-1\\ \\ \bar{y}=y+2/3\\ \\ \lambda=b-3/2\\ \end{array}\right.$
+The difference b3/2b-3/2 of system (4.4) is played by $\lambda$ of Theorem 4.7.
+The equilibrium $E_{3}$ is a singular fold point and a Hopf bifurcation occurs at $b=3/2$.
+Using Theorem 4.7, we perform a change of variables and parameter as
+[…]
 to obtain the equivalent system ($\tau=t/\varepsilon$)
-$\left\{\begin{array}[]{l}\dot{\bar{x}}=\bar{y}-\frac{1}{3}\bar{x}^{3}-\bar{x}^{2}\\ \\ \dot{\bar{y}}=\varepsilon\left(-\bar{x}-\lambda\bar{y}+\frac{2}{3}\lambda-\frac{3}{2}\bar{y}\right).\end{array}\right.$
-In the notation of Theorem 4.7 we have
+$\left\{\begin{array}[]{l}\dot{\bar{x}}=\bar{y}-\bar{x}^{2}-\frac{1}{3}\bar{x}^{3}\\ \\ \dot{\bar{y}}=\varepsilon\left(-\bar{x}+\frac{2}{3}\lambda-\lambda\bar{y}-\frac{3}{2}\bar{y}\right).\end{array}\right$.
+Using the notation introduced in Theorem 4.7, we obtain
 $l_{1}=-1\qquad l_{2}=-1-\bar{x}/3\qquad l_{3}=0\qquad l_{4}=1\qquad l_{6}=-\frac{3}{2}$
-hence $A=-1/2<0$ and $B=-3/4<0$ therefore the Hopf bifurcation is supercritical for $b$ in the original system (4.9).
+hence A=1/2<0A=-1/2<0 and B=3/4<0B=-3/4<0 therefore the Hopf bifurcation occurs for values $\lambda>\lambda_{H}(\Leftrightarrow b>\frac{3}{2}+\lambda_{H})$ .
 Since for $b<T_{H}^{2,3}(c)$ the eigenvalues of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,$ are negative, then the bifurcating periodic solution is unstable and canards occur when
@@
 4.3. Digestive remark
-We describe the dynamics of (3.1) in the case $a=0$, focusing on the first quadrant of the bifurcation diagram in the $(b,c)$–plane. We suggest the reader follows the description by using Figure 6.
+We describe the dynamics of (3.1) in the case $a=0$, focusing on the first quadrant of the bifurcation diagram in the (b,c)(b,c)–plane. We suggest the reader follows the description by using Figure 6.
 For $b\in(0,1)$ and $c>1$, there exists a unique equilibrium point $E_{1}$, which is unstable.
-This source is enclosed by a stable periodic orbit denoted by $C^{s}$.
-At $b=1$, one of the eigenvalues of the source $E_{1}$ becomes zero, leading to the creation of two additional unstable nodes $E_{2}$ and $E_{3}$.
-$E_{1}$ undergoes a supercritical Pitchfork bifurcation.
-The saddles $E_{2}$ and $E_{3}$ subsequently become unstable foci and undergo a supercritical Hopf bifurcation, giving rise to two unstable periodic solutions (symmetric under $\kappa$).
+This source is enclosed by a stable periodic solution denoted by $C^{s}$.
+At $b=1$, one of the eigenvalues of $Df_{(0,b,c)}$ at $E_{1}$ becomes zero, leading to the creation of two additional unstable nodes $E_{2}$ and $E_{3}$.
+In other words, $E_{1}$ undergoes a supercritical (with respect to $b$) Pitchfork bifurcation.
+The saddles $E_{2}$ and $E_{3}$ subsequently become unstable foci and undergo a supercritical Hopf bifurcation, leading to two unstable periodic solutions (symmetric under $\kappa$).
 Around $b=DH(c)$, the invariant manifolds of $E_{1}$ evolve into a double homoclinic orbit, which is unstable, previously identified in [8].
@@
 For $c\gg 1$, there is a small portion of the bifurcation parameter diagram $(b,c)$ of Figure 6, where the periodic orbits emerging from $T_{H}^{2,3}$ correspond to canards.
-They rapidly (with respect to $b$) collapse into the double homoclinic cycle to $E_{1}$.
-The values of $b_{c}$ that give rise to canards are at most $\mathcal{O}\left(\varepsilon\right)$ away from the values $b_{H}$ where a Hopf bifurcation occurs (see Figure 7):
+They rapidly collapse into the double homoclinic cycle to $E_{1}$.
+Parameters $b_{c}$ that yield canards are only $\mathcal{O}\left(\varepsilon\right)$ away from the corresponding Hopf bifurcation values $b_{H}$ (see Figure 7):
 $b_{c}-b_{H}=\frac{1}{2}\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).$
 Figure 7.
-Sketch of the singular bifurcations for $b\geq 1$ and $c=c^{\infty}$, in Case A.
-The smaller $\varepsilon$ is, the narrower the interval of $[b_{c},b_{H}]$.
-When $\varepsilon=0$, we have $b_{c}\equiv b_{H}$.
+Sketch of the singular bifurcations for $b\geq 1$ and $c<\infty$, in Case A.
+As $\varepsilon$ decreases, the interval $b_{c},b_{H}]$ becomes narrower.
+When $\varepsilon=0\,(\Leftrightarrow c=+\infty$), we have $b_{c}\equiv b_{H}$.
 4.4. Case B $(b=0)$
 Proposition 4.10.
-With respect to (4.4), there exists a continuous function $a_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $a$ that gives rise to a canard in the system, asymptotically defined by:
+With respect to (4.4), a continuous mapping $a_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ exists that assigns to each $\varepsilon\in\left[0,\varepsilon_{0}\right]$ a corresponding value $a$ giving rise to a canard, defined asymptotically by
 $a_{c}\left(\sqrt{\varepsilon}\right)=-\frac{\varepsilon}{8}+\mathcal{O}\left(\varepsilon^{3/2}\right).$
 Proof.
-The difference $a-1$ of system (4.4) is played by $\lambda$ in Theorem 4.7.
-We are interested in only one of the two fold points of (3.1) located at $(1,-2/3)$.
+The difference a1a-1 of system (4.4) is played by $\lambda$ in Theorem 4.7.
+We are interested in only one of the two fold points of (3.1) located at (1,2/3)(1,-2/3).
 After shifting the fold point to the origin, reversing time $t\to-t$ and setting $\lambda=a-1$, we get
@@
 4.5. Digestive remark
-The dynamics of system (3.1) for $b=0$ are illustrated in Figures 4 and 9.
-Our analysis focuses on the first and fourth quadrants of the bifurcation diagram in the $(a,c)$–plane.
-For $a<-1$, the saddle $E_{1}$ becomes a stable node.
-Near $a=-1$, a Belyakov transition occurs, followed by a supercritical Hopf bifurcation that gives rise to a stable periodic orbit that involves into a canard in an exponentially small portion of the bifurcation plane.
-This periodic solution disappears through a subcritical Hopf bifurcation at $a=1$.
-5. The phase portraits in the Poincaré disc
-In this section, we study the global phase portraits for system (3.1) in the Poincaré disc by using the theory described in Subsections 2.2 and 2.3.
+The dynamics of system (3.1) for $b=0$ is illustrated in Figures 4 and Table 2.
+Our analysis focuses on the first and fourth quadrants of the bifurcation diagram in the (a,c)(a,c)–plane and $c>0$.
+For a<1a<-1, the saddle $E_{1}$ becomes a stable node.
+Near a=1a=-1, a Belyakov transition occurs, followed by a supercritical Hopf bifurcation that gives rise to a stable periodic orbit that involves into a canard in an exponentially small portion of the bifurcation plane.
+This periodic solution disappears through another Hopf bifurcation at $a=1$.
+5. The phase portraits “near infinity”
+Studying the dynamics “near infinity” provides insight into the system’s asymptotic behaviour and allows the identification of trajectories that leave bounded sets.
+Using the techniques of Subsections 2.2 and 2.3, we compactify system (3.1) and describe the dynamics of (3.1) at infinity – this is what we call the dynamics “near infinity”.
+We distinguish between four cases
+$b\geq 0,\,c>0,\qquad b\geq 0,\,c<0,\qquad b<0,\,c>0,\qquad b<0,\,c>0.$ (5.1)
+The main theorem of this section is the following:
 Theorem 5.1.
-The global phase portraits of (3.1) in the Poincaré disc are topologically equivalent to one of the phase portraits of Figures 8, 9 and 10 for parameters values in Cases A, B and C, respectively.
-Remark 5.2.
-The calculations used in the proof of Theorem 5.1 could be streamlined and simplified by using weighted or quasi-homogeneous blow-ups.
-By a pedagogical point of view, we decided to do three blow-ups, which are illustrated in Figures 8, 9 and 10 for parameters values in Cases A, B and C.
-Some computations in the following proof were carried out using the software Maple.
+The dynamics of system (3.1) “near infinity” is topologically equivalent to one of the phase portraits depicted in Figure 8.
+Insight into the reasoning
+Before proceeding with the formal proof of the result, we give a geometrical interpretation of what is going to be done: we investigate the behaviour of trajectories of the planar vector field (3.1) as their position tends toward infinity.
+To do this, we apply the Poincaré compactification of Subsection 2.3, which transforms points at infinity into a bounded region represented by a disk containing a finite number of equilibria.
+We then study the qualitative dynamics of the “compactified” vector field paying particular attention to the equilibria located there, the directions along which trajectories diverge, and the conditions under which the flow at infinity becomes attracting or repelling.
+Some equilibria are singularities.
+In order to study their dynamics we need to desingularize them using the techniques of Subsection 2.2.
+(a) $b\geq 0$ and $c>0$
+(b) $b\geq 0$ and $c<0$
+(c) $b<0$ and $c<0$
+(d) $b<0$ and $c>0$
+Figure 8. Dynamics near infinity on the Poincaré disc of system (3.1).
 Proof.
-System (3.1) in the local chart $U_{1}=\{z_{1}>0\}$ may be written as:
+Using (2.3) with $d=3$, system (3.1) in the local chart
+$U_{1}=\{(y_{1},y_{2},y_{3})=(x_{1},x_{2},1)\in\mathbb{S}^{2}:y_{1}>0\}$
+may be written as:
 $\dot{u}=$ $-cu^{2}v^{2}-\frac{b}{c}uv^{2}-cuv^{2}+\frac{a}{c}v^{3}-\frac{1}{c}v^{2}+\frac{c}{3}u,$ (5.2)
-$\dot{v}=$ $-cuv^{3}-cv^{3}+\frac{c}{3}v.$
-The origin is the unique equilibrium on $v=0$ (the points of $\mathbb{S}^{1}$ in any chart have $v=0.$) The eigenvalues of Jacobian matrix associated with the vector field (5.1) are $\dfrac{c}{3},\dfrac{c}{3}.$ So, the origin is an unstable node (resp.
-stable) for $c>0$ (resp.
-$c<0$.)
-System (3.1) in the local chart $U_{2}$ may be written as:
+$\dot{v}=$ $-cuv^{3}-cv^{3}+\frac{c}{3}v,$
+where
+$u=\frac{x_{2}}{x_{1}}\quad\text{and}\quad v=\frac{1}{x_{1}}.$
+The system (5.2) has a unique equilibrium at the origin on $v=0$ (in any chart, points of $\mathbb{S}^{1}$ are characterized by $v=0$).
+The eigenvalues of the Jacobian matrix associated to (5.2) computed at the origin, are $\dfrac{c}{3},\dfrac{c}{3}.$ So, the origin is an unstable (resp.
+stable) node for $c>0$ (resp.
+$c<0$).
+Analogously, using (2.4) and $d=3$, the system (3.1) in chart $U_{2}$ takes the form:
 $\dot{u}=$ $\frac{1}{c}u^{2}v^{2}-\frac{a}{c}uv^{3}-\frac{c}{3}u^{3}+\frac{b}{c}uv^{2}+cuv^{2}+cv^{2},$ (5.3)
 $\dot{v}=$ $\frac{1}{c}uv^{3}-\frac{a}{c}v^{4}+\frac{b}{c}v^{3}.$
-The origin is an identically zero singular point. By doing the blow-up $(u,v)\rightarrow(u_{1},w)$ where $w=\dfrac{v_{1}}{u_{1}},$ we get
-$\dot{u}_{1}$ $=-\frac{a}{c}{u_{1}}^{4}{v_{1}}^{3}+\frac{1}{c}{u_{1}}^{4}{v_{1}}^{2}+c{u_{1}}^{3}{v_{1}}^{2}+\frac{b}{c}{u_{1}}^{3}{v_{1}}^{2}+c{u_{1}}^{2}{v_{1}}^{2}-\frac{c}{3}{u_{1}}^{3},$ (5.3)
-$\dot{v}_{1}$ $=-c{u_{1}}^{2}{v_{1}}^{3}-cu_{1}{v_{1}}^{3}+\frac{c}{3}{u_{1}}^{2}{v_{1}}.$
-By removing the common factor $u_{1},$ we have
-$\dot{u}_{1}$ $=-\frac{a}{c}{u_{1}}^{3}{v_{1}}^{3}+\frac{1}{c}{u_{1}}^{3}{v_{1}}^{2}+c{u_{1}}^{2}{v_{1}}^{2}+\frac{b}{c}{u_{1}}^{2}{v_{1}}^{2}+c{u_{1}}{v_{1}}^{2}-\frac{c}{3}{u_{1}}^{2},$ (5.4)
-$\dot{v}_{1}$ $=-c{u_{1}}{v_{1}}^{3}-c{v_{1}}^{3}+\frac{c}{3}{u_{1}}{v_{1}}.$
-The origin is an identically zero singular point again.
-For studying the local phase portrait at the origin, we do another blow-up (see §2.2 of the present paper).
-The characteristic direction at the origin is the real factors of $\frac{2c}{3}{u_{1}}^{2}v_{1}.$ Then the vertical axis $u_{1}=0$ is a characteristic direction.
-So, we translate the direction $u_{1}=0$ to the direction $u_{1}=v_{1}$ doing the change of variables $(u_{1},v_{1})=(u_{2}-v_{2},v_{2}).$ Then, the new system is given by
-$\dot{u_{2}}=$ $\dfrac{1}{3c}\Big{(}3c^{2}u_{2}v_{2}-c^{2}u_{2}^{2}-2c^{2}v_{2}^{2}+3c^{2}u2v_{2}^{2}+3bu_{2}^{2}v_{2}^{2}+3c^{2}u_{2}^{2}v_{2}^{2}+3u_{2}^{3}v_{2}^{2}$ (5.5)
-$\qquad-6c^{2}v_{2}^{3}-6bu_{2}v_{2}^{3}-9c^{2}u_{2}v_{2}^{3}-9u_{2}^{2}v_{2}^{3}-3au_{2}^{3}v_{2}^{3}+3bv_{2}^{4}+6c^{2}v_{2}^{4}$
-$\qquad+9u_{2}v_{2}^{4}+9au_{2}^{2}v_{2}^{4}-3v_{2}^{5}-9au_{2}v_{2}^{5}+3av_{2}^{6}\Big{)},$
-$\dot{v_{2}}=$ $\dfrac{c}{3}v_{2}\left(u_{2}-v_{2}-3v_{2}^{2}-3u_{2}v_{2}^{2}+3v_{2}^{3}\right).$
-The origin is the only singular point of system above on the line $u_{2}=0,$ and since the linear part is identically zero, we do the vertical blow-up $(u_{2},v_{2})\rightarrow(u_{3},w)$ where $w=\dfrac{v_{3}}{u_{3}},$ and we have
-$\dot{u}_{3}=$ $\dfrac{1}{3c}u_{3}^{2}\Big{(}3c^{2}v_{3}-c^{2}-2c^{2}v_{3}^{2}+3c^{2}u_{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{2}+3c^{2}u_{3}^{2}v_{3}^{2}+3u_{3}^{3}v_{3}^{2}-6c^{2}u_{3}v_{3}^{3}$
-$\qquad\,-6bu_{3}^{2}v_{3}^{3}-9c^{2}u_{3}^{2}v_{3}^{3}-9u_{3}^{3}v_{3}^{3}-3au_{3}^{4}v_{3}^{3}+3bu_{3}^{2}v_{3}^{4}+6c^{2}u_{3}^{2}v_{3}^{4}+9u_{3}^{3}v_{3}^{4}$
-$\qquad\,+9au_{3}^{4}v_{3}^{4}-3u_{3}^{3}v_{3}^{5}-9au_{3}^{4}v_{3}^{5}+3au_{3}^{4}v_{3}^{6}\Big{)},$ (5.6)
-$\dot{v}_{3}=$ $\dfrac{1}{3c}u_{3}v_{3}(1-v_{3})\Big{(}2c^{2}-2c^{2}v_{3}-6c^{2}u_{3}v_{3}^{2}-3bu_{3}^{2}v_{3}^{2}-6c^{2}u_{3}^{2}v_{3}^{2}-3u_{3}^{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{3}$
-$\hskip 79.6678pt+6c^{2}u_{3}^{2}v_{3}^{3}+6u_{3}^{3}v_{3}^{3}+3au_{3}^{4}v_{3}^{3}-3u_{3}^{3}v_{3}^{4}-6au_{3}^{4}v_{3}^{4}+3au_{3}^{4}v_{3}^{5}\Big{)}.$
-Eliminating the common factor $u_{3}$ leads to
-$\dot{u}_{3}=$ $\dfrac{1}{3c}u_{3}\Big{(}3c^{2}v_{3}-c^{2}-2c^{2}v_{3}^{2}+3c^{2}u_{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{2}+3c^{2}u_{3}^{2}v_{3}^{2}+3u_{3}^{3}v_{3}^{2}-6c^{2}u_{3}v_{3}^{3}$
-$\qquad\,-6bu_{3}^{2}v_{3}^{3}-9c^{2}u_{3}^{2}v_{3}^{3}-9u_{3}^{3}v_{3}^{3}-3au_{3}^{4}v_{3}^{3}+3bu_{3}^{2}v_{3}^{4}+6c^{2}u_{3}^{2}v_{3}^{4}+9u_{3}^{3}v_{3}^{4}$
-$\qquad\,+9au_{3}^{4}v_{3}^{4}-3u_{3}^{3}v_{3}^{5}-9au_{3}^{4}v_{3}^{5}+3au_{3}^{4}v_{3}^{6}\Big{)},$ (5.7)
-$\dot{v}_{3}=$ $\dfrac{1}{3c}v_{3}(1-v_{3})\Big{(}2c^{2}-2c^{2}v_{3}-6c^{2}u_{3}v_{3}^{2}-3bu_{3}^{2}v_{3}^{2}-6c^{2}u_{3}^{2}v_{3}^{2}-3u_{3}^{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{3}$
-$\hskip 79.6678pt+6c^{2}u_{3}^{2}v_{3}^{3}+6u_{3}^{3}v_{3}^{3}+3au_{3}^{4}v_{3}^{3}-3u_{3}^{3}v_{3}^{4}-6au_{3}^{4}v_{3}^{4}+3au_{3}^{4}v_{3}^{5}\Big{)}.$
-This system has two singular points on the line $u_{3}=0,$ namely $\tilde{E}_{1}:(0,0),$ and $\tilde{E}_{2}:(0,1).$ The eigenvalues of the Jacobian matrix associated with the vector field (5.7) at $(0,0)$ are $-\dfrac{c}{3},\dfrac{2c}{3}.$ Hence, $\tilde{E}_{1}$ is a saddle point.
-The linear part of the system at $\tilde{E}_{2}$ is identically zero.
-Therefore, in order to know the local phase portrait around this point, we must do blow-up.
-At first, we translate the singular point $\tilde{E}_{2}$ at the origin by doing the change of variables $(u_{3},v_{3})=(u_{4},1+v_{4}).$ System (5.7) in the new variables becomes
-$\dot{u}_{4}=$ $\dfrac{1}{3c}u_{4}\Big{(}-3c^{2}u_{4}-c^{2}v_{4}-12c^{2}u_{4}v_{4}+3c^{2}u_{4}^{2}v_{4}-2c^{2}v_{4}^{2}-15c^{2}u_{4}v_{4}^{2}+3bu_{4}^{2}v_{4}^{2}$
-$\qquad\quad+12c^{2}u_{4}^{2}v_{4}^{2}-6c^{2}u_{4}v_{4}^{3}+6bu_{4}^{2}v_{4}^{3}+15c^{2}u_{4}^{2}v_{4}^{3}-3u_{4}3v_{4}^{3}+3au_{4}^{4}v_{4}^{3}$
-$\qquad\quad+3bu_{4}^{2}v_{4}^{4}+6c^{2}u_{4}^{2}v_{4}^{4}-6u_{4}^{3}v_{4}^{4}+9au_{4}^{4}v_{4}^{4}-3u_{4}^{3}v_{4}^{5}+9au_{4}^{4}v_{4}^{5}+3au_{4}^{4}v_{4}^{6}\Big{)},$
-$\dot{u}_{4}=$ $-\dfrac{1}{3c}v_{4}(1+v_{4})\Big{(}-6c^{2}u_{4}-2c^{2}v_{4}-12c^{2}u_{4}v_{4}+3bu_{4}^{2}v_{4}+6c^{2}u_{4}^{2}v_{4}-6c^{2}u_{4}v_{4}^{2}$
-$\qquad\qquad\qquad\quad+6bu_{4}^{2}v_{4}^{2}+12c^{2}u_{4}^{2}v_{4}^{2}-3u_{4}^{3}v_{4}^{2}+3au_{4}^{4}v_{4}^{2}+3bu_{4}^{2}v_{4}^{3}+6c^{2}u_{4}^{2}v_{4}^{3}$
-$\hskip 85.35826pt-6u_{4}^{3}v_{4}^{3}+9au_{4}^{4}v_{4}^{3}-3u_{4}^{3}v_{4}^{4}+9au_{4}^{4}v_{4}^{4}+3au_{4}^{4}v_{4}^{5}\Big{)}.$ (5.8)
-The origin is the unique singular point of system above on the line $u_{4}=0$ and furthermore, the characteristic direction at the origin of system above is $u_{4}=0.$ Hence, by using the change of variables $(u_{4},v_{4})=(u_{5}-v_{5},v_{5}),$ we translate the direction to $u_{5}=v_{5},$ and we have
-$-3c\dot{u_{5}}=$ $3{u_{5}}^{4}{v_{5}}^{5}-15{u_{5}}^{3}{v_{5}}^{6}+27{u_{5}}^{2}{v_{5}}^{7}-21{u_{5}}{v_{5}}^{8}+6{v_{5}}^{9}+6{u_{5}}^{4}{v_{5}}^{4}-33{u_{5}}^{3}{v_{5}}^{5}$
-$+63{u_{5}}^{2}{v_{5}}^{6}-51{u_{5}}{v_{5}}^{7}+15{v_{5}}^{8}+3{u_{5}}^{4}{v_{5}}^{3}-21{u_{5}}^{3}{v_{5}}^{4}-6c^{2}{u_{5}}^{3}{v_{5}}^{4}$
-$-3b{u_{5}}^{3}{v_{5}}^{4}+12b{u_{5}}^{2}{v_{5}}^{5}+45{u_{5}}^{2}{v_{5}}^{5}+24c^{2}{u_{5}}^{2}{v_{5}}^{5}-15b{u_{5}}{v_{5}}^{6}-30c^{2}{u_{5}}{v_{5}}^{6}$
-$-39{u_{5}}{v_{5}}^{6}+12c^{2}{v_{5}}^{7}+6b{v_{5}}^{7}+12{v_{5}}^{7}-3{u_{5}}^{3}{v_{5}}^{3}-15c^{2}{u_{5}}^{3}{v_{5}}^{3}-6b{u_{5}}^{3}{v_{5}}^{3}$
-$+9{u_{5}}^{2}{v_{5}}^{4}+27b{u_{5}}^{2}{v_{5}}^{4}+63c^{2}{u_{5}}^{2}{v_{5}}^{4}-36b{u_{5}}{v_{5}}^{5}-9{u_{5}}{v_{5}}^{5}-81c^{2}{u_{5}}{v_{5}}^{5}$
-$+33c^{2}{v_{5}}^{6}+15b{v_{5}}^{6}+3{v_{5}}^{6}-3b{u_{5}}^{3}{v_{5}}^{2}-12c^{2}{u_{5}}^{3}{v_{5}}^{2}+18b{u_{5}}^{2}{v_{5}}^{3}-27b{u_{5}}{v_{5}}^{4}$
-$+60c^{2}{u_{5}}^{2}{v_{5}}^{3}-90c^{2}{u_{5}}{v_{5}}^{4}+42c^{2}{v_{5}}^{5}+12b{v_{5}}^{5}-3c^{2}{u_{5}}^{3}{v_{5}}+30c^{2}{u_{5}}^{2}{v_{5}}^{2}$
-$+3b{u_{5}}^{2}{v_{5}}^{2}-69c^{2}{u_{5}}{v_{5}}^{3}-6b{u_{5}}{v_{5}}^{3}+42c^{2}{v_{5}}^{4}+3b{v_{5}}^{4}+12c^{2}{u_{5}}^{2}{v_{5}}$
-$-40c^{2}{u_{5}}{v_{5}}^{2}+26c^{2}{v_{5}}^{3}+3c^{2}{u_{5}}^{2}-11c^{2}{u_{5}}{v_{5}}+6c^{2}{v_{5}}^{2},$ (5.9)
-$-3c\dot{v_{5}}=$ ${v_{5}}(1+{v_{5}})(-3{u_{5}}^{3}{v_{5}}^{4}+9{u_{5}}^{2}{v_{5}}^{5}-9{u_{5}}{v_{5}}^{6}+3{v_{5}}^{7}-6{u_{5}}^{3}{v_{5}}^{3}+18{u_{5}}^{2}{v_{5}}^{4}$
-$-18{u_{5}}{v_{5}}^{5}+6{v_{5}}^{6}-3{u_{5}}^{3}{v_{5}}^{2}+3b{u_{5}}^{2}{v_{5}}^{3}+6c^{2}{u_{5}}^{2}{v_{5}}^{3}+9{u_{5}}^{2}{v_{5}}^{3}-6b{u_{5}}{v_{5}}^{4}$
-$-12c^{2}{u_{5}}{v_{5}}^{4}-9{u_{5}}{v_{5}}^{4}+3b{v_{5}}^{5}+6c^{2}{v_{5}}^{5}+3{v_{5}}^{5}+6b{u_{5}}^{2}{v_{5}}^{2}+12c^{2}{u_{5}}^{2}{v_{5}}^{2}$
-$-24c^{2}{u_{5}}{v_{5}}^{3}-12b{u_{5}}{v_{5}}^{3}+12c^{2}{v_{5}}^{4}+6b{v_{5}}^{4}+6c^{2}{u_{5}}^{2}{v_{5}}+3b{u_{5}}^{2}{v_{5}}-18c^{2}{u_{5}}{v_{5}}^{2}$
-$-6b{u_{5}}{v_{5}}^{2}+12c^{2}{v_{5}}^{3}+3b{v_{5}}^{3}-12c^{2}{u_{5}}{v_{5}}+12c^{2}{v_{5}}^{2}-6c^{2}{u_{5}}+4c^{2}{v_{5}}).$
-Since the linear part is identically zero, we do the vertical blow-up $(u_{5},v_{5})=(u_{6},u_{6}v_{6}),$ for investigating the local phase portrait. The new system writes (applying the blow-up):
-$-3c\dot{u}_{6}=$ ${u_{6}}^{2}(3c^{2}-11c^{2}{v_{6}}+12c^{2}{u_{6}}{v_{6}}-3c^{2}{u_{6}}^{2}{v_{6}}+6c^{2}{v_{6}}^{2}-40c^{2}{u_{6}}{v_{6}}^{2}+3b{u_{6}}^{2}{v_{6}}^{2}$
-$+30c^{2}{u_{6}}^{2}{v_{6}}^{2}-3b{u_{6}}^{3}{v_{6}}^{2}-12c^{2}{u_{6}}^{3}{v_{6}}^{2}+26c^{2}{u_{6}}{v_{6}}^{3}-6b{u_{6}}^{2}{v_{6}}^{3}-69c^{2}{u_{6}}^{2}{v_{6}}^{3}$
-$+18b{u_{6}}^{3}{v_{6}}^{3}+60c^{2}{u_{6}}^{3}{v_{6}}^{3}-3{u_{6}}^{4}{v_{6}}^{3}-6b{u_{6}}^{4}{v_{6}}^{3}-15c^{2}{u_{6}}^{4}{v_{6}}^{3}+3{u_{6}}^{5}{v_{6}}^{3}$
-$+3b{u_{6}}^{2}{v_{6}}^{4}+42c^{2}{u_{6}}^{2}{v_{6}}^{4}-27b{u_{6}}^{3}{v_{6}}^{4}-90c^{2}{u_{6}}^{3}{v_{6}}^{4}+9{u_{6}}^{4}{v_{6}}^{4}+27b{u_{6}}^{4}{v_{6}}^{4}$
-$+63c^{2}{u_{6}}^{4}{v_{6}}^{4}-21{u_{6}}^{5}{v_{6}}^{4}-3b{u_{6}}^{5}{v_{6}}^{4}-6c^{2}{u_{6}}^{5}{v_{6}}^{4}+6{u_{6}}^{6}{v_{6}}^{4}+12b{u_{6}}^{3}{v_{6}}^{5}$
-$+42c^{2}{u_{6}}^{3}{v_{6}}^{5}-9{u_{6}}^{4}{v_{6}}^{5}-36b{u_{6}}^{4}{v_{6}}^{5}-81c^{2}{u_{6}}^{4}{v_{6}}^{5}+45{u_{6}}^{5}{v_{6}}^{5}+12b{u_{6}}^{5}{v_{6}}^{5}$
-$+24c^{2}{u_{6}}^{5}{v_{6}}^{5}-33{u_{6}}^{6}{v_{6}}^{5}+3{u_{6}}^{7}{v_{6}}^{5}+3{u_{6}}^{4}{v_{6}}^{6}+15b{u_{6}}^{4}{v_{6}}^{6}+33c^{2}{u_{6}}^{4}{v_{6}}^{6}$
-$-39{u_{6}}^{5}{v_{6}}^{6}-15b{u_{6}}^{5}{v_{6}}^{6}-30c^{2}{u_{6}}^{5}{v_{6}}^{6}+63{u_{6}}^{6}{v_{6}}^{6}-15{u_{6}}^{7}{v_{6}}^{6}+12{u_{6}}^{5}{v_{6}}^{7}$
-$+6b{u_{6}}^{5}{v_{6}}^{7}+12c^{2}{u_{6}}^{5}{v_{6}}^{7}-51{u_{6}}^{6}{v_{6}}^{7}+27{u_{6}}^{7}{v_{6}}^{7}+15{u_{6}}^{6}{v_{6}}^{8}-21{u_{6}}^{7}{v_{6}}^{8}$
-$+6{u_{6}}^{7}{v_{6}}^{9}),$ (5.10)
-$3c\dot{v}_{6}=$ ${u_{6}}{v_{6}}({v_{6}}-1)(-9c^{2}+6c^{2}{v_{6}}-30c^{2}{u_{6}}{v_{6}}+3b{u_{6}}^{2}{v_{6}}+9c^{2}{u_{6}}^{2}{v_{6}}+26c^{2}{u_{6}}{v_{6}}^{2}$
-$-6b{u_{6}}^{2}{v_{6}}^{2}-51c^{2}{u_{6}}^{2}{v_{6}}^{2}+12b{u_{6}}^{3}{v_{6}}^{2}+30c^{2}{u_{6}}^{3}{v_{6}}^{2}-3{u_{6}}^{4}{v_{6}}^{2}+3b{u_{6}}^{2}{v_{6}}^{3}$
-$+42c^{2}{u_{6}}^{2}{v_{6}}^{3}-24b{u_{6}}^{3}{v_{6}}^{3}-72c^{2}{u_{6}}^{3}{v_{6}}^{3}+9{u_{6}}^{4}{v_{6}}^{3}+15b{u_{6}}^{4}{v_{6}}^{3}+33c^{2}{u_{6}}^{4}{v_{6}}^{3}$
-$-12{u_{6}}^{5}{v_{6}}^{3}+12b{u_{6}}^{3}{v_{6}}^{4}+42c^{2}{u_{6}}^{3}{v_{6}}^{4}-9{u_{6}}^{4}{v_{6}}^{4}-30b{u_{6}}^{4}{v_{6}}^{4}-66c^{2}{u_{6}}^{4}{v_{6}}^{4}$
-$+36{u_{6}}^{5}{v_{6}}^{4}+6b{u_{6}}^{5}{v_{6}}^{4}+12c^{2}{u_{6}}^{5}{v_{6}}^{4}-15{u_{6}}^{6}{v_{6}}^{4}+3{u_{6}}^{4}{v_{6}}^{5}+15b{u_{6}}^{4}{v_{6}}^{5}$
-$+33c^{2}{u_{6}}^{4}{v_{6}}^{5}-36{u_{6}}^{5}{v_{6}}^{5}-12b{u_{6}}^{5}{v_{6}}^{5}-24c^{2}{u_{6}}^{5}{v_{6}}^{5}+45{u_{6}}^{6}{v_{6}}^{5}-6{u_{6}}^{7}{v_{6}}^{5}$
-$+12{u_{6}}^{5}{v_{6}}^{6}+6b{u_{6}}^{5}{v_{6}}^{6}+12c^{2}{u_{6}}^{5}{v_{6}}^{6}-45{u_{6}}^{6}{v_{6}}^{6}+18{u_{6}}^{7}{v_{6}}^{6}+15{u_{6}}^{6}{v_{6}}^{7}$
-$-18{u_{6}}^{7}{v_{6}}^{7}+6{u_{6}}^{7}{v_{6}}^{8}).$
-Now, by elimination of the common factor $u_{6},$ we obtain the truncated system
-$\dot{u_{6}}=$ $-\dfrac{1}{3c}u_{6}\left(3c^{2}-11c^{2}v_{6}+\mathcal{O}(|u_{6},v_{6}|^{2})\right),$ (5.11)
-$\dot{v_{6}}=$ $\dfrac{1}{3c}v_{6}(v_{6}-1)\left(-9c^{2}+6c^{2}v_{6}+\mathcal{O}(|u_{6},v_{6}|^{2})\right).$
-The equilibrium points for system (5.11) on the line $u_{6}=0$ are given by
-$\tilde{E}_{4}=(0,0),\qquad\tilde{E}_{5}=(0,1),\qquad\tilde{E}_{6}=\left(0,\frac{3}{2}\right).$ (5.12)
-The eigenvalues of Jacobian matrix associated with vector field (5.11) at $\tilde{E}_{4}$ and $\tilde{E}_{5}$ are $-c$, $3c$ and $\frac{2c}{3},-c$.
-So, they are saddle points.
-On the other hand, the eigenvalues associated with Jacobian matrix at $\tilde{E}_{6}$ are 00 and $\frac{3c}{2}$, and it is a semi-hyperbolic equilibrium.
-Given the center manifold $\dot{u_{6}}=-\frac{b}{8c}{u_{6}}^{3}$, we conclude that the equilibrium point $\tilde{E}_{6}$ is:
+The origin is an identically zero singular point for system (5.3). In order to desingularize this equilibrium, we are going to employ the quasi-homogeneous blow-up
+$(u,v)=(\bar{u}\rho^{2},\bar{v}\rho^{3}),$ (5.4)
+where $(\alpha,\beta)=(2,3)$ has been obtain by using the Newton diagram discussed in Subsection 2.2. The Newton polygon associated to (5.3) is illustrated in Figure 9.
+Figure 9.
+The points/segments associated to (5.3) are $(2,2)$, $(1,3)$, $(3,0)$, $(1,2)$, $(0,2)$, $(0,4)$ and $(0,3)$.
+The Newton polygon is defined by the line $3y+2x\geq 6$.
+Therefore $(\alpha,\beta)=(2,3)$.
+Applying the transformation (5.4) where $\bar{u}=1$ and $\bar{u}=-1,$ the $x$ directional blow-up, converts the system (5.3) into the vector fields defined by:
+$X_{1}^{+}:\begin{cases}\frac{d{\rho}}{d\eta}=\frac{\rho\left(-3a\rho^{5}{\bar{v}}^{3}+3{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+3b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{6c},\\ \\ \frac{d\bar{v}}{d\eta}=-\frac{\bar{v}\left(-a\rho^{5}{\bar{v}}^{3}+{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{2c},\end{cases}$
+and
+$X_{1}^{-}:\begin{cases}\frac{d{\rho}}{d\eta}=\frac{\rho\left(-3a\rho^{5}{\bar{v}}^{3}-3{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+3b{\rho}^{2}{\bar{v}}^{2}-3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{6c},\\ \\ \frac{d\bar{v}}{d\eta}=-\frac{\bar{v}\left(-a\rho^{5}{\bar{v}}^{3}+5{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{2c},\end{cases}$
+respectively, where $d\eta=\rho^{4}d\tau$ (the old time variable is $\tau$ and the new one is $\eta$; when we calculate $X^{1}$, we can cancel $\rho^{4}$ by using this time rescaling).
+We now need to analyze the equilibria at $\rho=0$ of $X_{1}^{\pm}.$ Regarding the vector field $X_{1}^{+},$ and $X_{1}^{-},$ one concludes that:
+  1. Claim 1:
+    The equilibria of system $X_{1}^{+}$ on $\rho=0$ are directly determined by the roots of
+    $\bar{v}\left(3\bar{v}^{2}-1\right)=0.$
+    There are three equilibria at $\rho=0$ denoted by
+    $D_{1}:(0,0)\qquad\text{and}\qquad f^{\pm}:\left(0,\pm\dfrac{\sqrt{3}}{3}\right).$
+    The following statements are valid:
     • 1.
-    a saddle point for
-    $\{(a,b,c)\big{|}b>0,\,c<0\}\cup\{(a,b,c)\big{|}b>0,\,c>0\}.$ (5.13)
+      The equilibrium $D_{1}$ always is a saddle point for $X_{1}^{+}$.
     • 2.
-    an unstable node for
-    $\{(a,b,c)\big{|}b<0,\,c<0\}.$ (5.14)
+      The points $f^{\pm}$ are semi-hyperbolic and the dynamics of $X_{1}^{\pm}$ restricted to the center manifold $\bar{v}=\mp\frac{\sqrt{3}(b+3c^{2})\rho^{2}}{18c^{2}}$ is given by $\dot{\rho}=\frac{b}{9c}{\rho^{3}}.$
+      Proof.
+      A center manifold in the form of $h(\rho):=\alpha_{1}\rho^{2}+h.o.t$ must satisfy the equation
+      $\dot{\rho}\frac{\partial h}{\partial\rho}-\dot{\bar{v}}\big|_{(\rho,h(\rho))}=0.$
+      After shifting $f^{\pm}$ to the origin, we obtain $\alpha_{1}=\mp\frac{\sqrt{3}(b+3c^{2})}{18c^{2}}.$ Then, the dynamics of $X_{1}^{\pm}$ restricted to the center manifold $\bar{v}=\mp\frac{\sqrt{3}(b+3c^{2})\rho^{2}}{18c^{2}}$ is given by $\dot{\rho}=\frac{b}{9c}{\rho^{3}}$ (more details in [5, Theorem 2.19]).
+∎
     • 3.
-    a stable node for
-    $\{(a,b,c)\big{|}b<0,\,c>0\}\cup\{(a,b,c)\big{|}b=0\}.$ (5.15)
-The final step of the proof follows from the description of Subsection 5.1. ∎
-(a) Region a
-(b) Region b
-(c) Region c
-(d) Region d
-(e) Region e
-(f) Region f
-(g) Region g
-(h) Region h
-(i) Region i
-(j) Region j
-(k) Region k
-(l) Region l
-(m) Region m
-(n) Region n
-(o) Region o
-(p) Region p
-Figure 8. The global phase portraits of (3.1) for Case A, which is associated with system (3.1) when $a=0,$ associated with regions a–j of Figure (3)(a).
-(a) Region I
-(b) Region II
-(c) Region III
-(d) Region IV
+      The equilibria $f^{\pm}$ are:
+      • 1a.
+[…]
+      • 2a.
+        an unstable node for $\{(a,b,c)\big|b<0,\,c<0\}$ – see Figure 10(c).
+      • 3a.
+        a stable node for $\{(a,b,c)\big|b<0,\,c>0\}.$ – see Figure 10(d).
+  2. Claim 2:
+    The equilibria of system $X_{1}^{-}$ on $\rho=0$ are determined by the roots of
+    $\bar{v}\left(3\bar{v}^{2}-1\right)=0.$
+    There are three equilibria at $\rho=0$ denoted by $D_{2}:(0,0)$ and $g^{\pm}:\left(0,\pm\frac{\sqrt{3}}{3}\right),$ and the results of Claim 1 are valid in this case (changing $D_{1}$ and $f^{\pm}$ by $D_{2}$ and $g_{\pm}$, respectively).
+    Using the transformation (5.4) with $\bar{v}=1$ and $\bar{v}=-1,$ the $y$ directional blow-up transforms system (5.3) into
+    $Y_{1}^{+}:\begin{cases}\frac{d{\rho}}{d\eta}=\mathcal{O}(|\rho,\bar{u}|),\\ \\ \frac{d\bar{u}}{d\eta}=c,\end{cases}\quad\quad Y_{1}^{-}:\begin{cases}\frac{d{\rho}}{d\eta}=\mathcal{O}(|\rho,\bar{u}|),\\ \\ \frac{d\bar{u}}{d\eta}=c,\end{cases}$
+    respectively, where $d\eta=\rho^{4}d\tau.$ We now need to analyze the point (0, 0) of $Y_{1}^{\pm}.$ Indeed,
+  3. Claim 3:
+    For the vector fields $Y_{1}^{\pm},$ point (0, 0) is a regular point (not an equilibrium).
+From the Claims 1–3, we know the dynamics of the vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}$.
+For the vector fields $Y_{1}^{\pm},$ $(0,0)$ is a regular point.
+Thus, Figure 10 illustrates the qualitative properties of the vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}.$ By shrinking the circles in these figures to the origin (as the radii of the circles decrease to zero and the circles collapse at $(0,0)$ ), we achieve to the phase portraits shown in Figure 11.
+So, as shown in Figure 8, the equilibrium of chart $U_{2}$ is a saddle with four hyperbolic sectors for $b\geq 0.$ Moreover, it becomes an unstable (resp.
+stable) node when $b,c<0$ and $b<0,c>0,$ respectively.
+(a) $b\geq 0$ and $c>0$
+(b) $b\geq 0$ and $c<0$
+(c) $b<0$ and $c<0$
+(d) $b<0$ and $c>0$
 Figure 10.
-The global phase portraits of (3.1) for Case B, which is associated with system (3.1) when $b=0,$ corresponding to regions I–IV of Figure 4(b), in the case where the eigenvalues of (3.5) at $E_{1}$ are complex non-real.
-5.1. Description of Figures 11, 12, 13, 14, 15 and 16 in Appendix 2
-In order to plot the complete phase diagram in the compactified space for the Cases A, B and C in Figures 8, 9 and 10, we need to carefully go back through each step of the blow-up procedure, explaining the transformations (blow-ups, directional transformations and coordinate maps) used at each stage, and clearly stating the resulting system at every step.
-Undoing the rescaling $dt_{3}=u_{6}dt_{2}$ the phase portrait depicted in Figures 11(a), 12(a), 13(a), 14(a), 15(a) and 16(a) yields the local phase portrait at the origin of (5.10), which is topologically equivalent to that of Figures 11(b), 12(b), 13(b), 14(b), 15(b) and 16(b), respectively.
-Going back through the change of variables $(u_{5},v_{5})=(u_{6},u_{6}v_{6}),$ the phase portrait depicted in Figures 11(b), 12(b), 13(b), 14(b), 15(b) and 16(b) yields the local phase portrait at the origin of (5.9), which is topologically equivalent to that of Figures 11(c), 12(c), 13(c), (14)(c), 15(c) and 16(c), respectively.
-Going back through the change of variables $(u_{4},v_{4})=(u_{5}-v_{5},v_{5}),$ the phase portrait depicted in Figures 11(c), 12(c), 13(c), 14(c), 15(c) and 16(c), yields the local phase portrait at the origin of (5.8), which is topologically equivalent to that of Figures 11(d), 12(d), 13(d), 14(d), 15(d) 16(d), respectively.
-Going back through the change of variables $(u_{3},v_{3})=(u_{4},1+v_{4})$ the phase portrait depicted in Figures 11(d), 12(d), 13(d), 14(d), 15(d) and 16(d) yields the local phase portrait at the origin of system (5.7), which is topologically equivalent to that of Figures 11(e), 12(e), 13(e), 14(e), 15(e) and 16(e), respectively.
-Undoing the rescaling $dt_{2}=u_{3}dt_{1}$ the phase portrait depicted in Figures 11(e), 12(e), 13(e), 14(e), 15(e) and 16(e) yields the local phase portrait at the origin of (5.6), which is topologically equivalent to that of Figures 11(f) and 12(f), 13(f), 14(f), 15(f) and 16(f), respectively.
-Going back through the change of variables $(u_{2},v_{2})\rightarrow(u_{3},v_{3}/u_{3}),$ the phase portrait depicted in Figures 11(f), 12(f), 13(f), 14(f), 15(f) and 16(f), yields the local phase portrait at the origin of system 5.5, which is topologically equivalent to that of Figures 11(g), 12(g), 13(g), 14(g), 15(g) and 16(g), respectively.
-Going back through the change of variables $(u_{1},v_{1})=(u_{2}-v_{2},v_{2}),$ the phase portrait depicted in Figures 11(g), 12(g), 13(g), 14(g), 15(g) and 16(g) yields the local phase portrait at the origin of (5.4), which is topologically equivalent to that of Figures 11(h), 12(h), 13(h), 14(h), 15(h) and 16(h), respectively.
-Undoing the rescaling $dt_{1}=u_{1}dt$ the phase portrait depicted in Figures 11(h), 12(h), 13(h), 14(h), 15(h) and 16(h) yields the local phase portrait at the origin of (5.3), which is topologically equivalent to that of Figures 11(i), 12(i), 13(i), 14(i), 15(i) and 16(i), respectively.
-Going back through the change of variables $(u,v)\rightarrow(u_{1},v_{1}/u_{1}),$ the phase portrait depicted in Figures 11(i), 12(i), 13(i), 14(i), 15(i) and 16(i) yields the local phase portrait at the origin of (5.2), which is topologically equivalent to that of Figures 11(j), 12(j), 13(j), 14(j), 15(j) and 16(j), respectively.
-This allows us to plot the complete phase diagram in the compactified space for the Cases A, B and C in Figures 8, 9 and 10.
-The local phase portraits of the blow-down at the origin of the different local charts have been plotted in Figures 1116.
-(a) $c<-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
-(b) $-\sqrt{\frac{b}{1-{x^{*}}^{2}}}<c<0$
-(c) $0<c<\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
-(d) $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
-[…]
+Qualitative properties of vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}.$ The axes corresponding to systems $X_{1}^{\pm}$ and $Y_{1}^{\pm}$ are $(\rho,\bar{v})$ and $(\rho,\bar{u}),$ respectively.
+(a) $b\geq 0$ and $c>0$
+(b) $b\geq 0$ and $c<0$
+(c) $b<0$ and $c<0$
+(d) $b<0$ and $c>0$
+Figure 11. Local phase portraits of system (5.3) (the representation of system (3.1) in the chart $U_{2}$) at $(0,0)$.
 6. Discussion and concluding remark
-Although the FHN equations were created as a simplified model for nerve impulse, they have also been intensively studied for purely mathematical reasons [12, 13, 10, 23] because they provide a very simple example of equations that exhibit rich dynamics.
+Although the FHN equations were originally developed as a simplified representation of nerve impulses, they have also been extensively investigated in purely mathematical purposes [12, 13, 10, 23] – they consist of a very simple model of equations with rich dynamics.
 Results obtained on [21, pp.
-180] synthesised the global bifurcation diagram for the FHN model (3.1).
-It has been obtained by putting together, as in a “huge puzzle”, all local bifurcation diagrams obtained in previous chapters of the latter reference.
+180] synthesised the global bifurcation diagram for the FHN system (3.1).
+It has been obtained by putting together, as in a “huge puzzle”, all local bifurcation diagrams obtained.
 The authors have concentrated their attention to a particular parameter region relevant to physiology ($|c|>1+\sqrt{3}$).
-Trying to complete and understand the bifurcation diagram of (3.1), in this paper we have discussed the finite equilibria of (3.1), as well as their bifurcations, for the following three scenarios identified in Subsection 1.4:
+Trying to complete and understand the bifurcation diagram of (3.1), in this paper we have discussed the finite equilibria of (3.1), as well as their bifurcations, for the three scenarios identified in Subsection 1.4:
 Case A:
@@
 Case C:
-$a\neq 0$ and $10<b<$ 1.
+$a\neq 0$ and $0<b<1$.
 The richest scenario is Case A.
@@
   • Region 11: three periodic solutions (one stable, two unstable);
-  • Point $\textbf{Q}_{6}$: Double Homoclinic (Theorem 3.4, curve $b={DH}$);
+  • Point $\textbf{Q}_{6}$: Double Homoclinic (Theorem 3.4, curve b=DHb={DH});
   • Region 17: two periodic solutions of different stabilities;
@@
 The precise location the above points/regions depend on $c$ but their relative position does not.
-The non-hyperbolic equilibria corresponding to parameter values situated at Q (see [21]) are degenerated saddle-node (cusp) equilibria, which are attracting for $c<1$ and repelling for $c>1$.
+The non-hyperbolic equilibria associated with parameter values located at Q (see [21]) are degenerated saddle-node (cusp) equilibria, which are attracting for $c<1$ and repelling for $c>1$.
 For $c=1$, following the Remark 3.6, this point might correspond to a degenerated Bogdanov-Takens of order two (with symmetry) – a kind of codimension-three bifurcation.
@@
 Using the same line of argument, in Cases B and C, we have analysed the dynamics of (3.1).
-Based on [10, 14], in Section 4, we have studied the asymptotic case $c\to\pm\infty$, where canards are detected.
-An important consideration is the empirical difficulty in finding canards, since the smaller $\varepsilon$ is, the narrower the interval of $b$ (Case A) or $a$ (Case B) values for which canards appear – see Figure 7.
-The dynamics produced in the system around these small intervals is called the canard explosion and has been studied in [10, 14, 12].
-In Section 5, we have continued the analysis with the compactification of the phase portraits associated with (3.1) on the Poincaré disc.
-This brings additional information of the trajectories which tend to or come from infinity.
+In Section 4, we have studied the asymptotic case $c\to\pm\infty$, where canards are detected.
+A key point in this research is the empirical difficulty to detect canards since the smaller value of $\varepsilon$ is, the narrower the interval of $b$ (Case A) or $a$ (Case B) values for which canards appear – see Figure 7.
+The dynamics that emerges in the system within these narrow parameter ranges is known as the canard explosion and has been studied in [10, 14, 12].
+In Section 5, we have continued the analysis with the compactification of the phase portraits of (3.1) on the Poincaré disc.
+This brings additional information of the trajectories which come from or tend to infinity.
 We have provided phase and bifurcation diagrams in all the three cases.
-Novelty versus limitations
-We have considered several cases of the classical FHN system which allow the catalogue of topologically distinct phase portraits.
-The global dynamics seems to be determined by the local bifurcations found in Theorems 3.4, 3.8 and Proposition 3.10.
-Although we have not achieved complete phase portraits for all cases, we have described all topological regions worth to be analysed.
-Numerics suggest that, in the three scenarios under consideration, we have obtained a complete phase portrait (up to conjugacy).
-However, we are not claiming it analytically.
-We have connected the local bifurcation theory (Theorems 3.4, 3.8 and Proposition 3.10) with the asymptotic dynamics $|c|\to+\infty$ (Propositions 4.9 and 4.10).
-Finally, using a blow-up technique we have depicted the dynamics of (3.1) at infinity (Theorem 5.1).
-Our contributions do not finish the whole discussion of the bifurcation analysis of (3.1); bifurcations might make part of of high codimension phenomena.
+Throughout this article, we have connected the local bifurcation theory (Theorems 3.4, 3.9 and Proposition 3.10) with dynamics “near infinity” (Theorem 5.1).
+Our contributions do not finish the whole discussion of the bifurcation analysis of (3.1); bifurcations might make part of high codimension phenomena.
 Theoretically, the maximum number of non-trivial periodic solutions of (3.1) is three [21, pp.
 218].
-There are regions where the dynamics is completely determined and the phase portrait is complete (Regions b, g, o and p of Figure 8 combined with Lemma 3.1) and others in which the complete analysis of the bifurcation diagram of [21, pp.
+There are regions where the dynamics is completely determined and the phase portrait is complete and others in which the complete analysis of the bifurcation diagram of [21, pp.
 180] is still ongoing.
-We defer the complete analysis for a future task.
-Acknowledgments
-Alexandre Rodrigues has been supported by the Project CEMAPRE/REM (UIDB/05069/ 2020) financed by FCT/MCTES through national funds.
-Nasrin Sadri is supported by Institute for Research in Fundamental Sciences (IPM) grant 1403370046.
-This work is based upon research funded by Iran National Science Foundation (INSF) under Project Number 4041099.
-The authors are also indebted to two reviewers for the suggestions which helped to improve the readability of this manuscript.
-Appendix A Blowing-down of the origin of the local chart $U_{2}$:
-Figure 11. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b>0$ and $c<0.$
-Figure 12. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b<0$ and $c<0.$
-Figure 13. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ when $b=0,$ $c<0.$
-Figure 14. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ when $b=0,$ $c>0.$
-Figure 15. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b>0$ and $c>0.$
-Figure 16. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b<0$ and $c>0.$
+7. Appendix
+All phase portrait are illustrated with numerical simulations in Tables 1 and 2 using the software pplane9.
+Initial conditions were rigorously chosen to illustrate the characteristic behaviour in each region.
+Furthermore, limit cycles are indicated by the red trajectories in the phase plot.
+The phase portraits of Table 1 and 2 have been plotted for the parameter values,
+$(c,b)=$ $(2.5,1.1),\;(2,1.2),\;(2,1.3),\;(2,1.34),\;(1.3,1.3),\;(0.5,1.1),\;(-0.5,1.1),\;(-2,1.5),$
+$(-2,1.34),\;(-2,1.33),\;(-2,1.2),\;(-2.5,1.1),\;(-3,0.5),\;(-1,0.5),\;(-0.5,0.4)\;$
+$(-0.1,0.9),\;(0.1,0.9),\;(0.4,0.4),\;(1.5,0.5),\;(3,0.5),\;(2.5,-0.5),\;(2,-0.5),\;$
+$(1,-1),\;(0.2,-1),\;(-0.2,-1),\;(-1,-1),\;(-1.5,-0.5),\;(-2.5,-0.5),$
+and
+$(c,a)=$ $(1,0.2),\;(-1,0.2),\;(5,2),\;(-5,2),\;(-1,0.5)\;(1,0.5)\;(-5,0.5)\;(5,0.5),$
+respectively.
 [Uncaptioned image] Region 1 [Uncaptioned image] Region 2 [Uncaptioned image] Region 3 [Uncaptioned image] Region 4 [Uncaptioned image] Region 5 [Uncaptioned image] Region 6 [Uncaptioned image] Region 7
@@
 [Uncaptioned image] Region 22 [Uncaptioned image] Region 23 [Uncaptioned image] Region 24 [Uncaptioned image] Region 25 [Uncaptioned image] Region 26 [Uncaptioned image] Region 27 [Uncaptioned image] Region 28
-Table 1. The local phase portraits corresponding to Case A, which is associated with system (3.1) when $a=0,$ for Regions 1–28 of Figure (3)(b).
+Table 1. The local phase portraits corresponding to Case A, which is associated with system (3.1) when $a=0,$ for Regions 1–28 of Figure 3.
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
 Region 5 Region 6 Region 7 Region 8
-Table 2. The local phase portraits corresponding to Case B, associated with system (3.1) when $b=0$, for Regions 1–8 of Figure 4(a).
+Table 2. The local phase portraits corresponding to Case B, which is associated with system (3.1) when $b=0,$ for Regions 1–8 of Figure 4.